### v1

# Edge-Cloud Collaborative Computing on Distributed Intelligence and Model Optimization: A Survey

Abstract
Edge-cloud collaborative computing (ECCC) has emerged as a pivotal paradigm for addressing the computational demands of modern intelligent applications, integrating cloud resources with edge devices to enable efficient, low-latency processing. Recent advancements in AI, particularly deep learning and large language models (LLMs), have dramatically enhanced the capabilities of these distributed systems, yet introduce significant challenges in model deployment and resource management. In this survey, we comprehensive examine the intersection of distributed intelligence and model optimization within edge-cloud environments, providing a structured tutorial on fundamental architectures, enabling technologies, and emerging applications. Additionally, we systematically analyze model optimization approaches, including compression, adaptation, and neural architecture search, alongside AI-driven resource management strategies that balance performance, energy efficiency, and latency requirements. We further explore critical aspects of privacy protection and security enhancement within ECCC systems and examines practical deployments through diverse applications, spanning autonomous driving, healthcare, and industrial automation. Performance analysis and benchmarking techniques are also thoroughly explored to establish evaluation standards for these complex systems. Furthermore, the review identifies critical research directions including LLMs deployment, 6G integration, neuromorphic computing, and quantum computing, offering a roadmap for addressing persistent challenges in heterogeneity management, real-time processing, and scalability. By bridging theoretical advancements and practical deployments, this survey offers researchers and practitioners a holistic perspective on leveraging AI to optimize distributed computing environments, fostering innovation in next-generation intelligent systems.
Index Terms:
Edge-Cloud Collaborative Computing, Distributed Intelligence, Model Optimization, AI, Survey
I Introduction
I-A Background
The increasing prevalence of Internet of things (IoT) devices and the demand for real-time applications have propelled a shift from centralized cloud computing to a distributed edge-cloud paradigm [1, 2, 3]. While cloud computing initially offers scalability and cost-effectiveness, limitations such as latency and bandwidth constraints spur the emergence of edge computing, bringing computation closer to data sources [4, 5, 6]. Consequently, edge-cloud collaborative computing (ECCC) arises, integrating the strengths of both paradigms to facilitate efficient resource utilization and enable complex AI tasks across the edge-cloud continuum [7, 8]. AI integration within edge-cloud systems addresses the inherent complexity of distributed environments, enabling optimized resource allocation and strategic model deployment for enhanced system performance [9]. In addition, AI empowers intelligent decision-making at the edge, enabling real-time responses. However, realizing distributed intelligence in edge-cloud systems introduces challenges, including efficient model partitioning and resource-aware model optimization, along with ensuring seamless edge-cloud collaboration and addressing security and privacy concerns [10, 11].
Edge-cloud collaborative computing has emerged as a crucial paradigm to support the growing demands of real-time, intelligent applications, particularly with the proliferation of artificial intelligence of things (AIoT) devices [12]. However, inherent limitations in edge devices, such as constrained computational power and limited battery life, hinder their ability to independently handle computationally intensive AI workloads [13]. Additionally, the dynamic nature of workloads in these environments requires sophisticated resource allocation and task offloading strategies to maintain performance and prevent bottlenecks [14]. Applications like autonomous driving [15] and industrial automation [16] further necessitate low latency and high reliability, demanding intelligent resource management and optimization [17]. Consequently, AI offers a powerful set of tools, including deep reinforcement learning (DRL) for dynamic resource allocation [18] and multi-objective optimization for task scheduling [19], to address these challenges and improve the efficiency and performance of edge-cloud systems. For instance, federated learning enables collaborative model training across distributed devices while preserving privacy [13], ultimately contributing to enhanced quality of service (QoS) by optimizing resource utilization and minimizing latency.
I-B Related Works
Over the past few years, several surveys have explored various aspects of edge computing and cloud collaboration, yet with varying scopes and technical emphases. Early works like Abbas et al. [20] and Ferrer et al. [21] establish foundational understanding of edge computing and decentralized cloud architectures respectively, but lack emphasis on AI-driven model optimization. Similarly, Liu et al. [22] provide valuable insights into edge computing systems and tools, while Wang et al. [23] contribute significantly to offloading algorithms, yet neither fully address the integration of AI techniques with edge-cloud architectures. More recent surveys have begun exploring particular dimensions of this space. Cao et al. [24] examine edge-cloud assisted cyber-physical systems with strong coverage of resource management but limited attention to model optimization techniques. Domain-specific surveys like Arthurs et al. [25] focus on intelligent transportation systems without broader applicability across diverse sectors. Despite promising advancements, emerging paradigms such as large language models (LLMs) at the network edge remain underexplored, with Qu et al. [26] providing a dedicated investigation into mobile edge intelligence for LLMs, yet primarily focusing on resource-efficient techniques rather than holistic system integration across the computing continuum. Additionally, specialized applications of edge intelligence for digital twin-enabled metaverse environments, as examined by Van Huynh et al. [3], offer valuable insights into ultra-reliable and low-latency communications but lack comprehensive treatment of model optimization frameworks. Moreover, while Yang et al. [4] extensively analyze reinforcement learning methodologies for mobile edge computing through an advanced exploration of optimization techniques, their work predominantly emphasizes network dynamics rather than systematically addressing privacy and cross-domain applications. Even the most recent comprehensive works by Yao et al. [11], Gu et al. [27], and Wang et al. [2] exhibit limitations—addressing either model optimization or edge-cloud collaboration but not systematically integrating with comprehensive privacy frameworks and cross-domain applications. As summarized in Table I, these surveys collectively fail to address the full spectrum of ECCC, particularly in integrating model optimization, privacy, security, and diverse applications. Consequently, our survey stands out by providing a holistic analysis that encompasses the entire computing continuum, advanced technical focuses including LLMs, and practical applications across manufacturing, autonomous driving, and healthcare, thereby filling a critical gap in the literature.
TABLE I: Comparative analysis of existing surveys related to AI-driven edge-cloud collaborative computing. Our work uniquely provides comprehensive coverage across the distributed computing continuum, technical focuses, and applications.
Ref. Year Topic Computing Continuum Technical Focuses Applications
Edge Cloud EC Comp. ECC Model Opt. Task Off. Res. Mgmt. Privacy Security LLMs Manuf. Auto. Drive Health
[20] 2018 Mobile Edge Comp., Network Architecture
[21] 2019 Decentralized Cloud, Ad Hoc Computing
[22] 2019 Edge Computing, Systems, Tools
[23] 2019 Edge Cloud, Offloading Algorithms
[24] 2021 Edge Computing, ECC, CPS
[25] 2022 EC Computing, ITS, Connected Vehicles
[3] 2022 Edge Intel., URLLC, Digital Twin
[28] 2023 Offloading, Federated Cloud-Edge-Fog
[29] 2023 Edge Computing, Sensor-Cloud
[11] 2023 EC Polarization, Collab. AI, Edge Comp.
[30] 2024 AIGC, Mobile Networks, EC Generative AI
[27] 2024 AI, CET Network, Collaborative Computing
[2] 2024 DL, Distributed Computing, Model Opt.
[4] 2025 RL, Mobile Edge Comp., Res. Mgmt.
[26] 2025 Mobile Edge Intel., LLMs, Res. Mgmt.
Ours 2025 ECCC, AI, Distributed Intel., Model Opt.
[…]
I-C Scope and Contributions
Figure 1: Overall structure of this survey. The framework is organized into five layers: (1) Foundations & architecture covering introduction and edge-cloud fundamentals, (2) technical implementation focusing on model optimization and resource management techniques, (3) protection framework addressing privacy and security enhancement, (4) practical deployment discussing applications and performance analysis, and (5) research vision outlining challenges and future directions.
As illustrated in Fig. 1, we provide a systematic and multi-layered examination of ECCC, addressing the critical intersection of distributed intelligence and model optimization. Edge-cloud systems face unprecedented challenges due to heterogeneous infrastructures, resource limitations, and growing computational demands from AI applications, making a comprehensive review of this rapidly evolving field both timely and necessary. By systematically analyzing edge-cloud collaborative computing architectures, resource management strategies, and performance metrics, the survey bridges existing gaps in understanding how AI can enhance decentralized computing environments. Moreover, it investigates critical enabling technologies such as model compression, adaptation, resource-aware optimization, and security mechanisms, which are pivotal for overcoming current limitations. The scope spans from architectural foundations and optimization techniques to practical applications and future research directions, offering researchers and practitioners a holistic perspective of this transformative paradigm. The contributions of this work are:
  • First Comprehensive Survey. To the best of our knowledge, this is the first comprehensive survey that specifically targets the intersection of distributed intelligence and model optimization within broader field of AI-driven edge-cloud collaborative computing, reviewing recent advances across architectures, techniques, and security.
  • Unified Architectural Framework. Our survey establishes a cohesive framework for understanding edge-cloud collaborative computing architectures, resource constraints, and distributed intelligence models. Through systematic categorization, researchers gain a structured understanding of how computational resources can be effectively distributed across the edge-cloud continuum.
  • Multi-domain Application Analysis. Through detailed applications spanning autonomous driving, smart healthcare, industrial automation, and smart cities, our work demonstrates the practical impact of AI-driven edge-cloud computing across diverse domains and identifies domain-specific requirements and challenges.
  • Future Research Roadmap. By identifying technical challenges and promising research directions, including advanced AI techniques, system improvements, and emerging applications, we provide a comprehensive roadmap for researchers and practitioners to advance the field of AI-driven edge-cloud collaborative computing.
I-D Research Methodology and Organization
Figure 2: Publication trends and citation analysis of research in distributed computing paradigms from 2021 to 2025.
Recent analysis reveals significant research momentum in ECCC, as illustrated in Fig. 2, where publication trends and citation patterns demonstrate the evolving landscape across four interrelated domains (The statistics data is collected as of 2025/03/05). While traditional edge computing maintains consistent publication volumes with steadily increasing citation impact, cloud computing exhibits characteristics of a mature research area with decreasing publication numbers yet sustained high citation rates. Most notably, edge-cloud computing shows clear upward trajectories in both publications and citations, signaling growing research interest, whereas the specialized subfield of edge-cloud collaboration displays exponential citation growth peaking in 2024, indicating its emerging importance. Our systematic review methodology employed comprehensive literature searches across prominent scholarly databases including IEEE Xplore, ACM Digital Library, ScienceDirect, and Web of Science, utilizing carefully selected keywords such as ”edge-cloud computing”, ”distributed intelligence”, ”model optimization”, ”AI”, and ”machine learning”. Following retrieval, selected papers pertinent to AI-driven ECCC were carefully reviewed and categorized based on their primary contributions aligned with the scope of this work.
The remainder of this survey is structured to provide a clear and comprehensive overview of AI-driven ECCC. We begin by establishing the fundamentals of edge-cloud computing architecture, including paradigms, distributed intelligence concepts, resources, and metrics in Sec. II. Sec. III examines model optimization techniques including compression, adaptation, resource-aware optimization, and neural architecture search. Subsequently, Sec. IV explores AI-driven resource management covering task offloading, resource allocation, and energy efficiency optimization, while Sec. V addresses privacy and security enhancement strategies. Practical applications appear in Sec. VI through applications spanning autonomous driving, smart healthcare, industrial automation, and smart cities. Sec. VII delivers comprehensive performance analysis through evaluation metrics, benchmarking methods, and system design evaluation, followed by identification of current challenges and future research directions in Sec. VIII, with concluding remarks in Sec. IX.
II Architecture and Fundamentals
Addressing the demands of modern intelligent applications requires architectures that efficiently process vast data volumes with low latency, thus motivating the edge-cloud computing paradigm. Fig. 3 provides a conceptual illustration of such AI-driven edge-cloud collaborative systems, depicting the core infrastructure interplay between cloud servers and edge nodes, alongside essential AI-enabled techniques (e.g., model adaptation, resource-aware optimization) that facilitate distributed intelligence across various application scenarios. Building upon this visual overview, the subsequent subsections delve into the foundational aspects of this architecture. We begin by outlining the core computing paradigm in Sec. II-A, subsequently examining the distributed intelligence framework enabling collaboration in Sec. II-B. Following this, an analysis of crucial resource types and their inherent constraints is presented in Sec. II-C.
II-A Edge-Cloud Computing Paradigm
Edge-cloud computing distributes computation and storage closer to data sources, leveraging a three-tier architecture of cloud, edge, and end devices [31]. The cloud tier provides extensive computational and storage resources for large-scale data processing and model training, while the edge tier, comprising edge servers and gateways, offers localized processing to reduce latency [32]. End devices, such as smartphones and IoT sensors, generate and consume data at the network edge. Effective resource management and orchestration are crucial for efficient resource allocation across these tiers. Distributing intelligence across the network, a key characteristic of this paradigm, can be represented by a function $D(R_{i},L_{j})$, quantifying the proportion of resource $R_{i}$ allocated to layer $L_{j}$ [33]. Edge-cloud systems aim to optimize resource allocation, considering factors like application requirements and network conditions, to maximize a utility function $U(P,C,L)$ encompassing performance $P$, cost $C$, and latency $L$, subject to resource constraints [34]. Consequently, processing data closer to its source reduces latency and bandwidth consumption. Additionally, distributing workloads enhances scalability and fault tolerance, improving the system’s ability to handle increasing data volumes and withstand node failures [35]. Edge federation models further enable collaborative resource sharing, optimizing service delivery across a federated network [36], especially for mission-critical applications [37].
II-B Distributed Intelligence Framework
Collaborative intelligence in edge-cloud systems leverages both edge and cloud resources to achieve optimal performance by distributing intelligence across the architecture. The distributed intelligence framework adheres to three key principles: task parallelism, data parallelism, and model parallelism. Task parallelism executes different parts of complex AI tasks concurrently on separate devices [38], represented as $T=\{t_{1},t_{2},...,t_{n}\}$, where $T$ represents the overall task and $t_{i}$ represents individual sub-tasks. Data parallelism trains identical models on different data subsets distributed across multiple devices [11], expressed as $D=\{d_{1},d_{2},...,d_{m}\}$, where $D$ represents the complete dataset and $d_{i}$ represents the data subset on device $i$. Model parallelism deploys different parts of large models on separate devices [39], formulated as $M=\{m_{1},m_{2},...,m_{p}\}$, where $M$ represents the complete model and $m_{i}$ represents the model partition on device $i$. Edge intelligence focuses on real-time processing with low latency [40], while cloud intelligence handles complex computations and large-scale training [41]. Model partitioning strategies balance workloads by considering model complexity, data size, and resource constraints [42]. Resource allocation algorithms dynamically assign resources based on real-time demands [43], while performance optimization techniques like model compression enhance efficiency [44].
II-C Resource Types and Constraints
Figure 3: Concepts of architecture. The diagram provides a holistic view of the technical scope covered in this survey, where the architecture integrates centralized cloud resources, decentralized edge nodes, and diverse mobile/edge devices. AI-enabled technologies include model compression, adaptation, and resource-aware optimization, which components jointly support distributed intelligence and efficient model execution across multiple application scenarios.
Computation, communication bandwidth, and storage capacity are fundamental resources in edge-cloud systems, each with inherent constraints impacting system design and performance [32]. Specifically, computation resources, denoted as $C_{k}$ for node $k$, represent processing power; communication bandwidth, $B_{ij}$, limits data transfer rates between nodes $i$ and $j$, directly affecting latency; and storage capacity, $S_{k}$, dictates the data volume node $k$ can store, influencing data-intensive applications [45]. Managing these constrained and heterogeneous resources across diverse devices, under dynamic workloads and with limited edge resources, especially energy for battery-powered devices [33], presents significant challenges.
Dynamic resource allocation addresses these challenges by adapting to workload fluctuations, aiming to find an optimal mapping of tasks $\mathcal{T}$ to nodes $\mathcal{N}$ that minimizes a cost function (e.g., latency) subject to resource constraints [46]. Mathematically, this involves minimizing $\sum_{t\in\mathcal{T}}Cost(M(t),t)$ subject to $\forall n\in\mathcal{N}:\sum_{t:M(t)=n}R(t)\leq C_{n}$, where $Cost(M(t),t)$ is the cost of executing task $t$ on node $M(t)$, and $R(t)$ is the resource requirement of task $t$. In addition, energy-efficient offloading minimizes energy consumption, $E_{ij}(t)=E_{comm}(t,B_{ij})+E_{comp}(t,C_{j})$, by strategically offloading tasks, where $E_{comm}$ and $E_{comp}$ represent communication and computation energy, respectively [23]. Efficient edge storage management, leveraging techniques like data caching and prefetching [47], is also crucial, especially for data-intensive applications, alongside integrating QoS considerations [48].
III Model Optimization for Edge-Cloud Systems
III-A Model Compression
Deploying sophisticated AI models on resource-constrained edge devices presents substantial challenges due to limited computational power, memory, and energy resources. Various model compression techniques [2, 49, 50] have emerged as critical solutions to reduce model size and computational demands while maintaining acceptable performance levels. Table II provides a comprehensive comparison of recent approaches across three primary compression categories: pruning, quantization, and knowledge distillation. Notably, pruning techniques demonstrate significant improvements in inference latency and energy efficiency by strategically eliminating less important parameters, as further detailed in Sec. III-A1. Meanwhile, quantization methods, explored in Sec. III-A2, reduce parameter precision requirements, with implementations achieving remarkable energy efficiency in healthcare applications. Knowledge distillation approaches, discussed in Sec. III-A3, enable smaller models to effectively mimic larger counterparts, with recent innovations specifically optimized for efficient LLM inference across edge-cloud boundaries. Across these diverse techniques, a common pattern emerges wherein most successful implementations leverage both edge and cloud resources collaboratively, highlighting the synergistic nature of modern compression strategies in distributed computing environments.
TABLE II: Summary of model compression techniques in AI-driven ECCC AIoT systems.
Method Year Venue Type Domain Performance Metrics Edge Cloud
TSCF [51] 2024 IEEE IoTJ P Edge-Cloud Systems Inference Latency $\checkmark$ $\checkmark$
Hybrid SD [52] 2024 ArXiv P Image Synthesis Parameter Efficiency, Image Quality $\checkmark$ $\checkmark$
PSFed [53] 2024 IEEE JSAC P Satellite Edge Cloud Latency, Energy, Privacy $\checkmark$ $\checkmark$
3CE2C [54] 2024 ASC Q Healthcare/ECG Accuracy, Energy Efficiency $\checkmark$ $\checkmark$
CDC [55] 2020 IJCAI P Edge-Cloud Systems Bandwidth Efficiency $\checkmark$ $\checkmark$
QuAsyncFL [56] 2024 IEEE IoTJ Q AIoT Communication Efficiency $\checkmark$ $\checkmark$
Zhang et al. [57] 2025 IEEE JSAC Q Aerial Edge Computing Mean Average Precision (mAP) $\checkmark$ $\checkmark$
DTS-KD [49] 2024 IEEE IoTJ K Industrial IoT F1 Score, Accuracy, Precision, Recall $\checkmark$ $\checkmark$
Wang et al. [58] 2024 CVPR K Multimodal LLMs Generalization Performance $\checkmark$ $\checkmark$
Wu et al. [59] 2024 IEEE TII P Aviation Manufacturing Edge Detection Accuracy
FedAgg [60] 2024 INFOCOM K End-Edge-Cloud Computing Model Accuracy, Convergence Rate $\checkmark$ $\checkmark$
Yang et al. [6] 2025 ADHOC K IoT Traffic Identification Classification Accuracy, Resource Efficiency $\checkmark$ $\checkmark$
GKT [61] 2024 ACL K LLM Inference Inference Speed, Accuracy, Cost $\checkmark$ $\checkmark$
OTO [62] 2024 IEEE TPDS K Reinforcement Learning Convergence Speed, Reward $\checkmark$ $\checkmark$
Zhang et al. [63] 2024 ESWA K Manufacturing Processes Fault Detection Accuracy $\checkmark$ $\checkmark$
Note for ’Type’: P = Pruning, Q = Quantization, K = Knowledge Distillation.
III-A1 Pruning
Strategically eliminating redundant model parameters, pruning techniques reduce size and complexity in ECCC, enabling efficient AI model deployment on resource-constrained edge devices [52]. We explore various strategies targeting different architectural aspects, such as pruning entire neurons or filters based on performance contribution, as in the two-phase split computing framework (TSCF) [51], or focusing on individual weights. Applied during or after training, dynamic and static pruning balance accuracy and computational cost. For instance, Hybrid SD [52] employs structural pruning of the Stable Diffusion Model’s U-Net to maintain image quality on edge devices, while the adaptive pruning-split federated learning (PSFed) method [53] optimizes semantic coder updates in a SemCom-assisted SEC framework, reducing training delay and energy consumption.
III-A2 Quantization
Lower-precision representation of model parameters effectively reduces both memory footprint and computational cost [55], making quantized models particularly suitable for resource-constrained edge devices. Instead of 32-bit floating-point representations, quantization employs lower precision formats such as 8-bit integers, consequently diminishing memory requirements and accelerating computations on these lower-precision numbers. For example, [54] explored model quantization for edge implementation in arrhythmia monitoring to reduce energy consumption. Additionally, QuAsyncFL [56] integrates quantization with asynchronous federated learning to enhance communication efficiency in cloud-edge-terminal collaborations within the AIoT domain.
III-A3 Knowledge Distillation
Model compression through knowledge distillation (KD) trains a smaller ”student” model to mimic a larger ”teacher” model’s behavior [61]. This process enables the student model to achieve comparable performance with reduced complexity, facilitating efficient deployment in resource-constrained edge-cloud environments [49]. Instead of learning from ground truth labels, the student model learns the teacher model’s output distribution, providing richer information than hard labels and enabling more nuanced representations and better generalization [58]. KD benefits ECCC by placing the teacher model in the resource-rich cloud and the distilled student model on edge devices [6], optimizing overall system performance. agglomerative federated learning (FedAgg) [60] uses a bridge sample based online distillation protocol (BSBODP) for knowledge transfer between nodes across different tiers, enabling larger model training while maintaining privacy and flexibility. Guidance-based knowledge transfer (GKT) [61] employs a larger LLM as a ”teacher” to generate guidance prompts for a smaller ”student” model, enhancing accuracy and speed without fine-tuning. Growth-adaptive distillation [6] automatically adjusts the student model structure for efficient, low-power edge deployment. In cloud-edge collaborative reinforcement learning, algorithm-independent knowledge distillation [62] accelerates online RL agent convergence by leveraging pre-trained cloud models and offline agent interactions. A federated bi-directional knowledge distillation strategy [63] enhances fault detection in manufacturing by integrating quality knowledge from both cloud and edge, strengthening interaction between sub-processes and hierarchical levels.
III-B Model Adaptation
Figure 4: Illustration of model adaptation, detailing (a) transfer learning, (b) federated learning, and (c) continual learning.
TABLE III: Summary of TL, FL, and CL techniques for enhancing AI model adaptability in ECCC systems.
Method Year Venue Type Domain Edge Cloud ECC Model Comp.
CoLLaRS [64] 2024 FGCS CL AIoT Applications $\checkmark$ $\checkmark$ $\checkmark$
Li et al. [65] 2024 IEEE TNSM TL Manufacturing $\checkmark$ $\checkmark$ $\checkmark$
Pan et al. [66] 2024 IEEE Netw. FL Large Model Services $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
MADDPG [67] 2024 IEEE IoTJ TL 6G Networks $\checkmark$ $\checkmark$ $\checkmark$
DTS-KD [49] 2024 IEEE IoTJ TL Industrial IoT $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
Wang et al. [58] 2024 CVPR CL Multimodal Large Language Models $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
Wang et al. [68] 2024 ESWA FL Rotating Machinery Fault Diagnosis $\checkmark$ $\checkmark$ $\checkmark$
FedAgg [60] 2024 INFOCOM FL End-Edge-Cloud Systems $\checkmark$ $\checkmark$ $\checkmark$
OTO [62] 2024 IEEE TPDS TL Cloud-Edge Collaborative DRL $\checkmark$ $\checkmark$ $\checkmark$
U-Federator [69] 2024 JNCA FL Cloud-Edge-Fog Computing $\checkmark$ $\checkmark$ $\checkmark$
OTFAC [70] 2024 IEEE IoTJ FL Cloud-Edge Collaborative IoT $\checkmark$ $\checkmark$ $\checkmark$
FLAT [71] 2024 IEEE ISCC FL Cloud-Edge-Client Continuum $\checkmark$ $\checkmark$ $\checkmark$
FedITA [72] 2024 AEI FL Industrial Motor Fault Diagnosis $\checkmark$ $\checkmark$ $\checkmark$
Khuen et al. [73] 2024 IEEE Access FL Precision Aquaculture $\checkmark$ $\checkmark$ $\checkmark$
Kim et al. [74] 2024 IEEE IoTJ FL Cloud-Edge-Terminal IoT Networks $\checkmark$ $\checkmark$ $\checkmark$
QuAsyncFL [56] 2024 IEEE IoTJ FL Cloud-Edge-Terminal AIoT $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
Loconte et al. [75] 2024 FGCS FL Cloud-to-Edge-IoT Continuum $\checkmark$ $\checkmark$ $\checkmark$
BAFL-DT [76] 2024 IEEE TSC FL Digital Twin Edge-Cloud Continuum $\checkmark$ $\checkmark$ $\checkmark$
FDRL [77] 2024 IJPR FL Cloud-Edge Manufacturing Systems $\checkmark$ $\checkmark$ $\checkmark$
Wei et al. [78] 2024 IEEE TPDS FL Edge Clouds $\checkmark$ $\checkmark$ $\checkmark$
Yang et al. [79] 2024 IEEE IoTJ FL AIoT Data Prediction $\checkmark$ $\checkmark$ $\checkmark$
FedCAE [80] 2024 IEEE TIE FL Machine Fault Diagnosis $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
FlexibleFL [81] 2024 INS FL Cloud-Edge Federated Learning $\checkmark$ $\checkmark$ $\checkmark$
PSPFL [82] 2024 IEEE TMC FL Mobile Edge-Cloud Networks $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
CCSFLF [83] 2024 CCPE FL Federated Learning Systems $\checkmark$ $\checkmark$ $\checkmark$
Zhou et al. [13] 2024 ADHOC FL AIoT Quality of Service $\checkmark$ $\checkmark$ $\checkmark$
CADCL-DTM [84] 2024 CAC CL Rolling Bearings Monitoring $\checkmark$ $\checkmark$
Li et al. [85] 2024 IEEE TMC CL Edge Computing Service Provisioning $\checkmark$ $\checkmark$ $\checkmark$
Lian et al. [86] 2024 SDI CL ITS Object Detection $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
EdgeC3 [87] 2024 IEEE TMC CL Distributed Deep Learning $\checkmark$ $\checkmark$ $\checkmark$
Note for ’Type’: TL = Transfer Learning, FL = Federated Learning, CL = Continual Learning
Effectively adapting AI models to new tasks, datasets, or evolving data distributions is crucial within the dynamic landscape of ECCC systems. To achieve this adaptability, three primary learning paradigms are widely employed: transfer learning (TL), which accelerates learning on target tasks by leveraging knowledge from pre-trained models; federated learning (FL), which facilitates privacy-preserving, distributed training across multiple clients; and continual learning (CL), which enables models to learn sequentially from evolving data streams without catastrophic forgetting. The interplay and distinct roles of these paradigms in fostering adaptability within ECCC systems are conceptually depicted in Fig. 4. Furthermore, Table III presents a comparative summary of recent research efforts utilizing TL, FL, and CL, highlighting their specific approaches, application domains, and system characteristics within ECCC. Following this, the subsequent sections delve into the transfer learning in Sec. III-B1, federated learning in Sec. III-B2, and continual learning in Sec. III-B3.
III-B1 Transfer Learning
Within edge-cloud environments, adapting pre-trained models to new tasks or datasets often leverages transfer learning [65]. This approach reduces training time and data requirements by building upon existing knowledge, thereby improving deployment efficiency, especially in resource-constrained systems [58]. For example, in surface defect detection, a model pre-trained on a general image dataset can be fine-tuned with a smaller, specialized dataset to identify specific defects [65]. Similarly, in cloud-device collaborative learning for multimodal large language models (MLLMs), techniques like adapter-based knowledge distillation (AKD) facilitate knowledge transfer from larger cloud models to smaller, compressed device models [58], maintaining performance while minimizing computational burden. In cloud-edge collaborative fault diagnosis, transfer learning addresses data heterogeneity and distribution shifts by fine-tuning source models trained on diverse user data to bridge distribution gaps and enable effective diagnosis on target datasets [68].
III-B2 Federated Learning
As a pivotal technique for distributed model training, federated learning [82] preserves data privacy by enabling clients to train local models on their own datasets and aggregate updates (e.g., gradients) on a central server to refine a global model [79]. This collaborative approach, by keeping sensitive data localized, not only addresses privacy concerns but also enhances the global model’s generalization capabilities by leveraging diverse datasets [73]. In edge-cloud systems, FL efficiently and securely leverages distributed computational resources [75], as exemplified by frameworks like FedCAE [80] for machine fault diagnosis and QuAsyncFL [56] for AIoT applications. Specifically, SFETEC [88] optimizes FL for thing-edge-cloud environments by splitting model training to reduce communication overhead and address non-IID data. For enhanced security against poisoning attacks, FlexibleFL [81] evaluates uploaded parameters and participant contributions. Additionally, optimizing participant selection and learning topologies, as in [78], improves FL efficiency, while integrating FL with technologies like blockchain [76] further bolsters security and trustworthiness in collaborative learning.
III-B3 Continual Learning
In ECCC, adapting models to new data or tasks incrementally is crucial for lifelong learning [89]. This adaptability enables models to evolve with dynamic data and emerging tasks, addressing the core challenge of catastrophic forgetting, where new information disrupts previously acquired knowledge [89]. Several strategies, including regularization, replay, and dynamic architectures, mitigate this issue, ensuring robust and adaptable systems. Specifically, continual learning enables digital twins to synchronize with their evolving physical counterparts and sensor data in edge computing, providing accurate predictions [85, 84]. For instance, in intelligent transportation systems (ITS), cloud-edge collaborative continual adaptation frameworks, leveraging visual prompts and knowledge distillation, enhance the generalization of lightweight edge models for object detection in dynamic environments [86]. Similarly, online management frameworks, such as EdgeC3, optimize model aggregation and dynamic offloading for continuous data streams in edge-cloud collaborative continual learning [87]. Cross-edge federated continual learning (Cross-FCL) frameworks, employing parameter decomposition and various cross-edge strategies, address knowledge retention across devices in multiple federated learning systems, balancing memory and adaptation for new tasks [90].
III-C Resource-Aware Model Optimization
Figure 5: Resource allocation strategies in edge-cloud system, which illustrates three approaches: (1) Dynamic resource allocation, (2) collaborative resource management, and (3) reinforcement learning optimization.
Dynamic adaptation to available resources and operational constraints emerges as a critical factor for model optimization in edge-cloud systems. As illustrated in Fig. 5, resource allocation integrates three complementary strategies to enhance ECCC performance across diverse environments. Dynamic resource allocation mechanisms continuously adapt to fluctuating workload demands and resource availability, orchestrating real-time service placement and task offloading decisions to maximize system efficiency. Building on this foundation, collaborative resource management techniques orchestrate allocation across multiple edge nodes and cloud resources, employing federated learning and hybrid expansion strategies to ensure cohesive operation. Further advancing this framework, reinforcement learning-based optimization refines system performance by integrating data-driven decisions informed by feedback on metrics such as latency and energy consumption, utilizing sophisticated methods like deep Q-networks and multi-agent coordination. Collectively, these strategies establish a versatile framework applicable to diverse domains, including IoT networks, drone communications, vehicular edge computing, and 5G environments.
III-C1 Dynamic Model Selection
Edge-cloud systems leverage distributed computing to optimize resource allocation and enhance application performance across environments. In ECCC, optimizing resource utilization and achieving desired performance necessitates selecting the most suitable AI model from a pool of candidates. Such dynamic model selection (DMS) process considers factors such as available resources at the edge and cloud (computation, communication bandwidth, and energy), task complexity, and desired QoS metrics, including latency and accuracy. Resource-constrained edge devices may require smaller, less computationally intensive models, potentially sacrificing some accuracy, while ample resources allow deploying more complex models for maximized performance [91]. Stringent latency requirements may prioritize models optimized for fast inference [46]. Additionally, DMS must adapt to the dynamic nature of edge-cloud environments, accommodating fluctuations in network connectivity or available resources to maintain optimal performance and resource efficiency [92]. Several research efforts have explored DMS strategies. Some leverage online learning techniques for dynamic adjustments based on real-time feedback [19]. Others employ reinforcement learning (RL) to learn optimal model selection policies for maximizing long-term performance [93]. Trust-aware DMS strategies incorporate the trustworthiness of edge resources into selection decisions [94].
III-C2 Model Splitting and Offloading
In ECCC, a crucial optimization technique involves partitioning a complex AI model into smaller sub-models and strategically deploying them across the edge-cloud continuum [95]. This distributed execution paradigm effectively leverages the heterogeneous nature of these systems, optimizing resource utilization and minimizing latency [96]. Computationally intensive components can be offloaded to the resource-rich cloud, while less demanding or latency-sensitive parts can be executed at the edge, closer to the data source [96]. For example, within a digital twin enabled 6G industrial IoT environment, this approach allows for efficient management of computationally intensive and delay-sensitive tasks within an end-edge-cloud architecture [97].
Several factors influence the optimization of model splitting and offloading decisions, including the computational capabilities of edge devices, communication bandwidth limitations, and application-specific latency requirements [98]. Resource-constrained edge devices often benefit from offloading complex layers of a CNN to the cloud [99]. Dynamic model selection and adaptation techniques, along with algorithms like multi-agent deep deterministic policy gradient (MADDPG), further enhance efficiency by tailoring the model architecture and execution strategy to the current system state and workload demands [100, 95]. Additionally, user mobility introduces further complexity, especially in heterogeneous edge-cloud computing scenarios with distributed resources [96].
III-C3 Reinforcement Learning for Model Optimization
In dynamic edge-cloud systems, RL provides a powerful mechanism for optimizing model parameters and offloading decisions [101]. RL agents learn optimal policies by leveraging environmental feedback, adapting model configurations and resource allocation strategies to improve system adaptability and performance. Specifically, techniques like deep Q-learning [102] and asynchronous proximal policy Optimization (APPO) [103] guide agents to explore actions and learn strategies that maximize cumulative rewards. For example, a study in [18] combined deep reinforcement learning with edge computing to optimize real-time monitoring and control in IoT environments, ultimately improving system stability and reducing failure rates. Additionally, RL enables sophisticated offloading mechanisms that consider factors such as latency, energy consumption, and task dependencies [104]. Through dynamic parameter adjustment, such as the number of layers or computational precision, RL optimizes the trade-off between model accuracy and computational cost [105]. Similarly, RL-based task offloading strategies dynamically determine optimal execution locations (edge or cloud) by adapting to real-time network and system conditions [103], simultaneously maximizing resource efficiency and reducing latency for time-critical applications.
III-D Neural Architecture Search
For automating the design of optimal neural network architectures in edge-cloud systems, NAS methods aim to enhance both model performance and efficiency. NAS techniques are broadly categorized into three primary approaches: reinforcement learning-based, evolutionary algorithm-based, and gradient-based. In reinforcement learning-based NAS, an agent learns to select optimal architectures based on performance feedback. Evolutionary algorithm-based NAS, conversely, employs evolutionary algorithms to evolve populations of network architectures towards optimal solutions. Gradient-based NAS leverages gradient descent to directly search the architecture space for optimal configurations.
III-D1 Reinforcement Learning-based NAS
RL-NAS employs RL algorithms to automate the neural architecture search process. Specifically, an RL agent interacts with an environment representing the architecture search space and receives rewards based on the generated architectures’ performance, typically measured by validation set accuracy while considering edge-cloud deployment resource constraints [106]. Through iterative exploration and exploitation, the agent learns to select actions that result in high-performance architectures tailored to edge-cloud systems. [105]. Consequently, this approach can discover architectures that outperform manually designed ones, especially in resource-constrained edge devices [107]. For example, an RL agent can be trained to optimize for both accuracy and latency, balancing these conflicting objectives [108]. By incorporating resource constraints into the reward function, RL-NAS can directly target architectures suitable for edge-cloud environments with limited resources like computational power, memory, and energy budgets [109]. This adaptability to diverse edge environments makes RL-NAS promising for optimizing model deployment in ECCC [67].
III-D2 Evolutionary Algorithm-based NAS
For ECCC, neural architecture search (NAS) can leverage evolutionary algorithms (EAs) to discover optimal architectures by mimicking natural selection [110]. EAs encode network architectures as individuals within a population, subjecting them to iterative selection, mutation, and crossover to enhance performance [111]. Selection favors architectures demonstrating superior accuracy, latency, or energy efficiency [112], while mutation introduces random architectural changes to explore new design possibilities [113]. Crossover combines elements of high-performance architectures, potentially yielding even better solutions [19]. This iterative refinement identifies architectures suited for the resource constraints and performance needs of edge-cloud environments [111]. For instance, the self-adaptive discrete particle swarm optimization algorithm with genetic algorithm operators (GA-DPSO) optimizes data transmission time for scientific workflows in edge-cloud systems [114]. The inherent adaptability of EAs proves particularly valuable in dynamic edge-cloud environments characterized by fluctuating resource availability and workload demands [115].
III-D3 Gradient-based NAS
Leveraging gradient descent, gradient-based NAS methods efficiently explore and optimize the architecture search space [19]. Unlike earlier discrete search methods, this approach formulates architecture search as a continuous optimization problem, enabling the use of established gradient-based optimization techniques [116]. Specifically, the architecture search space is represented differentiably, allowing gradient descent to update architectural parameters and iteratively refine the model for optimal performance and efficiency. Consequently, search time is significantly reduced compared to discrete search methods, facilitating exploration of complex architecture spaces and tailoring models for diverse edge-cloud environments [107]. For example, DARTS [116] constructs a differentiable search space and optimizes it using gradient descent, achieving high-performance architectures rapidly. Despite these advantages, challenges remain, such as handling strong correlations among operations, which can lead to suboptimal architectures. In addition, techniques like weight sharing and channel bottlenecks can further enhance efficiency by reducing the computational cost of evaluating different architectures [117]. Critically, incorporating device constraints, including model size and floating-point operations, directly into the search process is essential for deploying optimized models on resource-constrained edge devices [117].
IV AI-Driven Resource Management
IV-A Task Offloading Mechanisms
Figure 6: Task offloading mechanisms, including latency-aware, energy-efficient, and dependency-aware offloading strategies.
TABLE IV: Summary of task offloading mechanisms optimizing latency, energy efficiency, and dependency management.
Method Domain Performance Metrics Performance Results Edge Cloud ECC Task Off.
Zhang et al. [19] Edge Cloud Architecture Time, Cost, Load balance, User satisfaction, Trust measurement Found more than half of the best solutions on benchmark functions $\checkmark$ $\checkmark$
Fan et al. [118] Edge-Cloud Cooperative Computing Task processing delay, Task queuing stability Superior performance compared to 4 other schemes $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
Qi et al. [92] Edge-Cloud Service Provisioning Service quality, Cost optimization Effective cost-service quality balance demonstrated $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
Abouaomar et al. [119] Edge Computing for IoT Latency, Resource consumption Outperforms benchmark approaches in latency and resource utilization $\checkmark$ $\checkmark$
Li et al. [120] Edge-Cloud Cooperation Long-term network performance, Energy constraints Converges to approximate optimal solution in polynomial time $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
Luckow et al. [121] Edge-to-Cloud Applications Performance impact of task placement 5-65% performance improvement depending on scenario $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
Cai et al. [122] Data-Intensive Edge Services Throughput, Delay, Resource consumption Superior performance compared to state-of-the-art algorithms $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
Pang et al. [123] Vehicle-Edge-Cloud Latency, Energy consumption Significant decrease in latency and energy consumption $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
PRAISE [124] Multi-Access Edge Cloud System benefits, Resource costs Outperforms baselines with higher benefits and lower costs $\checkmark$ $\checkmark$ $\checkmark$
Khalafi et al. [125] Edge-Cloud Orchestrated Networks System cost, QoS, Delay violations Reduced costs associated with network slicing $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
Peng et al. [126] Edge Computing Systems Job computing time, Job blocking probability Optimal replication factor varies with distribution parameters and arrival rate $\checkmark$ $\checkmark$ $\checkmark$
DRLIS [127] Edge-Fog-Cloud Computing Load balancing, Response time, Weighted cost Reduces execution cost by up to 55% (load balancing), 37% (response time), 50% (weighted cost) $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
Hua et al. [96] Heterogeneous Edge-Cloud Computing Energy consumption Significantly reduced energy consumption of mobile devices $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
Arroba et al. [128] Sustainable Edge Computing Energy efficiency, Economic viability Early results demonstrate benefits for future computing environments $\checkmark$ $\checkmark$
PDCO [129] Cloud-Edge Collaboration Total energy consumption, Resource utilization Approaches optimal performance in energy consumption and resource utilization $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
He et al. [130] Wireless Powered MEC Systems Amount of processed data Enhances computation performance compared to benchmark methods $\checkmark$ $\checkmark$
iGATS [8] IoT-Edge-Cloud Networks Response latency, Temporal satisfaction, Load-balancing Outperforms baseline techniques for reducing response latency $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
ETCRA [94] Edge-Cloud Workflows Comprehensive system function, Execution time, Energy consumption, Reliability Significantly outperforms other baselines on four key measurements $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
Hu et al. [131] MEC-enabled Mobile IoT Energy efficiency, Service delay Achieves [O(1/V), O(V)] tradeoff between EE and service delay $\checkmark$ $\checkmark$
Alopex-DE [111] Local-Edge-Cloud Computing Latency, Energy consumption Superior effectiveness compared to three state-of-the-art algorithms $\checkmark$ $\checkmark$ $\checkmark$
DT-MADDPG [97] Digital Twin Enabled 6G IIoT Task success rate, Latency, Energy consumption Significantly improved task success rate with reduced latency and energy consumption $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
Fan et al. [132] Cloud-Edge-End Cooperative Networks Task processing delay, Energy consumption, Block mining speed Superior performance compared to 5 reference schemes in 7 scenarios $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
Zhang et al. [133] Edge-Cloud Systems Task completion time, Cost, Energy consumption, System reliability Outperforms existing design approaches $\checkmark$ $\checkmark$ $\checkmark$
Qi et al. [92] Edge-Cloud Collaboration Cost, Service quality Effective in both synthetic and real-world data experiments $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
EdgeOptimizer [134] Kubernetes-Based Edge-Cloud Clusters Success rate of time-critical tasks, Overall overload Improved success rate for time-critical tasks $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
Effective resource utilization and performance optimization in ECCC depend critically on sophisticated task offloading mechanisms. As depicted in Fig. 6, a framework outlines three core strategies—latency-aware, energy-efficient, and dependency-aware offloading—detailing their interactions with edge devices, cloud servers, and supporting systems to improve task placement and scheduling. Specifically, these mechanisms address distinct objectives such as real-time performance, energy consumption, and task dependencies, thereby boosting overall system efficiency. To further illustrate their practical impact, Table IV summarizes recent task offloading methods in edge-cloud environments. Notably, the table reveals significant performance gains, such as up to a 33.42× reduction in delay for sensitive workloads [135], alongside improvements in energy consumption and system reliability, highlighting the versatility of these strategies across diverse application domains.
IV-A1 Latency-Aware Offloading
Minimizing response time through latency-aware offloading represents a crucial strategy for time-sensitive applications in edge-cloud systems, with numerous approaches demonstrating significant performance gains. According to the comparative analysis in Table IV, methods like iGATS [8] and DRLIS [127] have achieved substantial reductions in processing delays through intelligent placement decisions. For instance, Lyapunov optimization techniques employed by Abouaomar et al. [119] enable dynamic resource allocation at the edge, while Poularakis et al. [15] address multidimensional constraints through joint optimization of service placement and request routing. Similarly, Fan et al. [118] propose a collaborative framework that minimizes total processing delay while ensuring queuing stability in multi-task scenarios. When considering delay-sensitive workloads specifically, Zhai et al. [135] demonstrate a remarkable 33.42× reduction in latency compared to traditional cloud deployments by intelligently distributing tasks based on their characteristics. Additionally, the double deep Q-network (DDQN) approach introduced by Pang et al. [123] effectively manages mobility while optimizing offloading decisions, while Qi et al. [92] present a two-timescale scheduling algorithm that balances service provider costs with stringent latency requirements.
IV-A2 Energy-Efficient Offloading
Strategic energy management across edge-cloud systems requires balancing computation and communication costs while optimizing resource allocation through specialized techniques. In ECCC, prolonging the battery life of edge devices necessitates balancing computation and communication energy costs, especially for battery-powered devices [131]. As shown in Table IV, numerous approaches including PDCO [129], Hua et al. [96], and ETCRA [94] have demonstrated significant improvements in energy efficiency across edge-cloud systems. Particularly for battery-powered devices, strategic decision-making have to consider both computational and communication energy costs, weighing factors like task complexity, network conditions, and hardware energy profiles [131, 130]. While offloading computationally intensive tasks to cloud resources can reduce device energy consumption, communication costs under limited bandwidth or unstable networks often counterbalance these benefits, necessitating sophisticated optimization approaches. Beyond basic consumption metrics, advanced energy-efficient mechanisms consider additional dimensions such as power control, transmission scheduling [96], and user mobility patterns [128]. Furthermore, in heterogeneous IoT deployments with diverse device capabilities and constraints, efficient resource allocation becomes increasingly crucial for sustainable service delivery [111], with some approaches even incorporating renewable energy sources to create more environmentally responsible edge-cloud ecosystems [135]. Energy-efficient offloading mechanisms strive to minimize overall energy consumption by jointly considering computation and communication costs. Additionally, optimizing power control and transmission scheduling can further enhance energy efficiency [96]. User mobility introduces another layer of complexity to energy optimization, as dynamic network conditions and varying proximity to edge servers impact energy trade-offs [128]. In heterogeneous IoT environments with diverse device capabilities and energy constraints, efficient resource allocation becomes even more crucial for energy-efficient service delivery [111]. Prioritizing workloads based on their energy requirements and leveraging renewable energy sources can contribute to a more sustainable edge-cloud ecosystem [135].
IV-A3 Dependency-Aware Offloading
Complex applications often consist of multiple interconnected subtasks with dependencies, forming a dependency graph where some subtasks must finish before others can start [124]. Such interdependence introduces complexities in task scheduling and resource allocation, potentially leading to suboptimal performance due to idle resources and increased latency if ignored. Therefore, efficient offloading mechanisms must consider these precedence constraints to optimize overall application execution. For example, the PRAISE algorithm in [124] addresses triangular dependencies by reconfiguring the task call graph into sequential layers for efficient resource allocation. Similarly, modeling task dependencies as a directed acyclic graph (DAG) optimizes offloading decisions, particularly in vehicle-edge-cloud environments [123]. Additionally, deep Q-networks (DQN) for real-time offloading in cloud-edge computing benefit from incorporating dependency awareness [136], enabling dynamic adaptation to changing environments without pre-set task priorities.
IV-B Resource Allocation Strategies
TABLE V: Summary of dynamic resource allocation methods.
Method Domain Performance Metrics Cloud ECC Model Comp. Res. Alloc.
DySCo [137] Edge-Cloud Infrastructure Deployment Cost, Delay
ADPRL [138] Collaborative Edge Computing Task Completion Time, Energy Consumption
Das et al. [139] Drone Communication Networks Latency, Scalability, Resource Utilization
FRL [74] Cloud-Edge-Terminal IoT Networks Learning Speed, Adaptation
DSCF [140] Distributed Serverless Edge Clouds Inference Latency, Resource Consumption
CADCL-DTM [84] Rolling Bearings Digital Twin Forgetting Performance, RUL Prediction Accuracy
A2C-DRL [141] Edge-Cloud Environments Reward Value, Task Rejection, Load-balancing
Lu et al. [142] IoV Edge Cloud Quality of Service
IntellectReq [143] Device-Cloud Recommendation Communication Efficiency, Recommendation Accuracy
AC-DRL [101] 5G Edge-Cloud Continuum QoE, Energy Consumption
Tian et al. [144] Things-Edge-Cloud Computing Throughput, Delay Utility
FDRL [77] Cloud-Edge Collaborative Manufacturing Makespan, Energy Consumption
DESKG [145] Edge-Cloud Collaborative PSS Service Efficiency, OEE
DRL-MTCN [146] Cloud-Fog-Edge Multi-Tier Networks Revenue, Revenue-Cost Ratio, Acceptance Rate
Fan et al. [118] Edge-Cloud Cooperative Computing Task Processing Delay
OTFAC [70] Cloud-Edge Collaborative IoT Delay, Energy Consumption
Wang et al. [147] IoT-Based Distribution Grid Convergence Rate, Performance Stability
DaCM [93] Edge-Cloud Computing Long-term Profit, Computation Delay
RL [105] Edge-Cloud Networks Execution Time, Network Usage, Migration Delay, Energy Consumption
DODQ [136] Cloud-Edge Computing Task Completion Time
DA-HADRL [148] SAGINs with Cloud-Edge Computing Energy Consumption, Latency
GRLDO [149] Client-Edge-Cloud Environment User Satisfaction
MARL-SF [142] IoV Edge Cloud Quality of Service
AC-DRL [101] 5G Edge-Cloud Continuum QoE, Energy Consumption
DA-DDQN [123] Vehicle-Edge-Cloud Collaboration Latency, Energy Consumption
DRL-LOA [150] Healthcare Cloud-Edge Networks Power Usage, Latency
MADDPG [67] End-Edge-Cloud in 6G Networks Delay, Energy Consumption
MARL-CA [151] Industrial Edge-Cloud Web 3.0 Communication Rate, Stability
OTO [62] Cloud-Edge Collaborative DRL Convergence Speed, Reward
HDMO [133] Edge-Cloud Collaboration Task Completion Time, Cost, Energy Consumption, System Reliability
LsiA3CS [152] Large-Scale IIoT Task Completion Time, Energy Consumption
RIS-DRL [153] RIS-Aided Cloud-Edge Computing Wireless Capacity, Latency, Energy Consumption
QRSRL [154] Train-Edge-Cloud Smart Systems Task Processing Delay
Informer-A3C [155] Vehicular Edge Cloud Computing Content Access Cost
Strategic allocation of computational resources across the edge-cloud continuum remains fundamental for maximizing performance in AI-driven collaborative environments. Accordingly, we explore three pivotal strategies: dynamic resource allocation, collaborative resource management, and reinforcement learning-based optimization, each designed to enhance task execution and system efficiency. By effectively distributing resources, these approaches address the challenges of varying workloads and resource constraints in distributed environments. To illustrate their practical impact, Table V provides a comparative analysis of recent dynamic resource allocation methods in ECCC systems. Notably, the table reveals a wide range of application domains—from IoT networks to vehicular edge clouds—alongside key performance metrics like latency, energy consumption, and quality of service, highlighting the adaptability and effectiveness of these strategies in diverse scenarios.
IV-B1 Dynamic Resource Allocation
We highlight key approaches for optimizing ECCC systems through effective task offloading strategies. Edge-cloud systems leverage dynamic resource allocation to address fluctuating workloads and optimize resource utilization, crucial for handling unpredictable real-world applications [48]. Approaches include dynamic NUMA node selection and core consolidation [137], optimizing service placement across the edge-cloud continuum by considering factors like delay and resource constraints. Dynamic task offloading mechanisms further optimize task placement and scheduling based on latency, energy consumption, and task dependencies [138].
In drone communication networks, dynamic resource allocation considers drone mobility, communication loads, and latency constraints for efficient real-time applications [139]. Similarly, for cloud-edge-terminal IoT networks, collaborative policy learning using federated reinforcement learning (FRL) enables dynamic task scheduling, adapting to changing workload demands [74]. Distributed serverless edge clouds utilize dynamic split computing frameworks to minimize inference latency while adhering to resource constraints by dynamically determining splitting points and maintaining container instance warm status [140]. Deep reinforcement learning techniques optimize task offloading decisions in 5G edge-cloud environments, maximizing quality-of-experience (QoE) while meeting latency requirements [101]. Additionally, multi-target-aware dynamic resource scheduling algorithms in multi-tier computing networks consider diverse QoS requirements, including computing, storage, bandwidth, and delay, to enhance resource flexibility and efficiency [146].
IV-B2 Collaborative Resource Management
Efficient ECCC requires coordinating resource allocation across edge nodes and the cloud. Such coordination is essential given the heterogeneous nature of edge devices and dynamic workload demands. For instance, in IoT-based distribution grids, resource optimization across cloud, edge, and end devices spans multiple timescales, with decisions at each level influencing others [147], necessitating sophisticated, interdependent resource management strategies. Dynamic resource allocation is also crucial for adapting to fluctuating demands and optimizing real-time performance [147].
One approach leverages optimal transport (OT) and FL principles. Specifically, an OT-based offline optimization model can initially allocate bandwidth and computation resources, followed by online optimization using federated actor-critic learning (OTFAC) [70]. In this dual-driven approach, edge servers can act as local aggregators within the federated learning process, improving efficiency and reducing delays [156]. Additionally, collaborative resource management benefits from a hybrid approach where edge servers not only collaborate with the cloud but also expand their local resources [93]. This strategy allows edge servers to dynamically purchase cloud resources and expand local capacity for long-term cost optimization [93]. Consequently, this adaptability manages dynamic workloads and varying demands. Edge-cloud and edge-edge cooperation further enhances performance through efficient network-wide resource utilization [118].
IV-B3 Reinforcement Learning for Resource Management
In dynamic edge-cloud systems, RL provides a powerful mechanism for optimizing resource allocation. By interacting with the edge-cloud environment and receiving rewards for performance-enhancing actions, RL agents learn adaptive strategies to improve metrics such as latency [104], energy consumption [105], and task completion time [138]. Specifically, these agents adapt to changing workloads and resource availability, leading to more efficient resource utilization. For example, RL optimizes task offloading decisions in vehicular edge-cloud computing (VECC) to minimize latency while maintaining quality of service [104], and addresses online dependent task offloading (ODTO) in collaborative edge computing by jointly optimizing network flow scheduling and resource allocation [138]. In addition, RL dynamically adjusts model parameters and offloading decisions, improving the adaptability and performance of edge-cloud systems [148]. Researchers have demonstrated RL’s effectiveness in various scenarios, including optimizing microservice placement [105], managing resources for delay-sensitive tasks [104], and handling dependent and parallel tasks [136]. Integrating RL with techniques like graph neural networks (GNNs) [149] and deep Q-networks (DQNs) [136] further enhances the agents’ ability to capture complex relationships and optimize offloading decisions. Multi-agent RL extends these capabilities by enabling decentralized learning and coordinated resource management across multiple servers [141].
IV-C Energy Efficiency Optimization
Power consumption is a critical concern in AI-driven ECCC. We explore strategies for minimizing this consumption without sacrificing performance, focusing on power consumption modeling in Sec. IV-C1, energy-aware scheduling (Sec. IV-C2), and joint optimization techniques in Sec. IV-C3. Sec. IV-C1 provides the foundation for understanding energy dynamics by investigating power consumption modeling within the edge-cloud ecosystem. Building upon these models, Sec. IV-C2 presents energy-aware scheduling algorithms to optimize task allocation and execution for minimal energy use. Subsequently, Sec. IV-C3 explores joint optimization techniques, recognizing the interdependence of energy efficiency and other performance metrics, such as latency, to achieve a balanced and sustainable system operation.
IV-C1 Power Consumption Modeling
Developing energy-efficient edge-cloud systems requires accurate power consumption models that capture power dynamics at both the device and system levels, considering the relationship between workloads and energy consumption. At the device level, models must account for diverse hardware components and their individual power characteristics, adapting to the heterogeneity of edge devices, which range from resource-constrained IoT sensors to powerful edge servers [11]. For example, [157] introduces a comprehensive energy model for the Arm Cortex-M0, considering factors like frequency and wait states that influence energy consumption. Specialized models, such as those in [158] for video transcoding and deep learning on SoC-Clusters, offer insights into workload-specific power dynamics.
System-level modeling encompasses individual device power usage and the energy costs of communication and data transfer between edge, cloud, and edge servers. Dynamic workload fluctuations and task offloading further complicate this process, requiring models to consider the interplay between computation, communication, and storage demands. Research like [159] explores energy-aware task migrations, while [96] addresses resource allocation in heterogeneous environments, optimizing power control and offloading to minimize energy consumption. Incorporating renewable energy and pricing, as in [135], allows for further optimization, while understanding the limitations of cloud-centric approaches in edge environments, as highlighted in [128], is crucial for developing effective models.
IV-C2 Energy-Aware Scheduling
Optimizing the energy efficiency of ECCC hinges on dynamic resource management, including adjusting hardware parameters and employing energy minimization strategies. Dynamic voltage and frequency scaling (DVFS) [160] enables real-time adaptation of processor voltage and frequency according to workload, achieving energy savings by lowering these parameters during periods of low activity. For example, DVFO [160] co-optimizes DVFS and offloading decisions to minimize energy consumption on edge devices during DNN inference. Additionally, sleep-wake mechanisms [161] offer further optimization by selectively powering down idle components. A Co-HDRL algorithm [161] demonstrates this in RIS-aided HetNets by jointly optimizing sleep and power control. Green computing strategies [128] broaden the scope of energy efficiency by considering hardware choices, software optimization, and power management policies. Specifically, approaches like two-phase immersion cooling and energy-centric federated management, explored in [128], contribute to sustainable edge infrastructure.
IV-C3 Joint Optimization Techniques
We examine recent work integrating collaborative intelligence in edge-cloud architectures across diverse application domains. Resource management in edge-cloud systems necessitates a holistic approach, jointly optimizing multiple objectives to balance conflicting goals like minimizing energy consumption and latency. Offloading tasks to the cloud reduces energy consumption on edge devices but introduces latency due to data transfer [96]. Multi-objective optimization (MOO) addresses this by finding Pareto optimal solutions representing the best trade-offs [98]. For example, [19] proposes a multi-objective resource scheduling model considering time, cost, load balance, user satisfaction, and trust, while [162] frames service deployment and task offloading as a multi-task, high-dimensional MOO problem, incorporating network reliability and load balancing. Joint optimization dynamically adjusts power consumption based on real-time workload demands, improving energy efficiency without sacrificing performance [112]. Specifically, [111] proposes an energy-efficient resource allocation scheme for local-edge-cloud collaboration, optimizing latency, resource allocation, and energy consumption using an alopex-based differential evolution algorithm. Similarly, [163] develops a hybrid multi-objective optimization algorithm for resilience in error-prone environments, balancing energy and reliability while meeting quality of service standards. In addition, [135] designs an edge-cloud collaboration system for mixed workloads, aiming to meet delay requirements while reducing carbon emissions and improving profits.
V Privacy and Security Enhancement
V-A Data Privacy Protection
Robust mechanisms are essential for safeguarding sensitive information throughout its lifecycle in the collaborative edge-cloud paradigm. We explore methods and techniques addressing this critical challenge, ensuring data protection within distributed environments. We begin by examining privacy-preserving learning methods, such as federated learning and secure aggregation, enabling collaborative model training without direct data sharing. Subsequently, we delve into secure multi-party computation (SMPC), which allows joint function computation on private inputs. Additionally, differential privacy techniques, adding noise to protect individual data points while preserving aggregate insights, are considered. Finally, data anonymization approaches, transforming data to prevent individual identification, are explored.
V-A1 Privacy-preserving learning methods
To enhance privacy in FL, techniques like differential privacy (DP) and secure aggregation are employed [164]. DP adds carefully calibrated noise to model updates before aggregation, making it difficult to infer individual data points [165]. Meanwhile, secure aggregation uses cryptographic protocols like multi-party computation (MPC) to aggregate updates without revealing individual contributions to the server [166]. Additionally, zero-concentrated differential privacy (zCDP) offers tighter privacy guarantees than traditional DP in FL [167], and CaPC learning combines MPC, homomorphic encryption, and privately aggregated teacher models to achieve both confidentiality and privacy [168]. Lightweight encryption and aggregation techniques further enhance security and efficiency in federated learning systems [169].
V-A2 Secure Multi-Party Computation
Multiple parties can collaboratively compute a function over their private inputs without revealing them through SMPC [170]. SMPC protocols ensure each party learns only the function’s output, safeguarding data privacy and security, crucial for edge-cloud collaborative AI model training. For example, multiple hospitals could train a joint model on patient data without sharing raw data due to privacy regulations [168]. Additionally, SMPC can decompose complex AI tasks into smaller sub-tasks distributed across edge devices and the cloud, improving training efficiency [171]. Robust management frameworks, such as FlexSMC [172], address the dynamic edge-cloud environment by facilitating node discovery, establishing trust, and ensuring robustness against failures. Coded computation within SMPC, as in coded-MPC (CMPC) [173], further enhances performance through resource optimization, with adaptive gap entangled (AGE) polynomial codes [174] as a promising approach. Ensuring security against malicious attacks during MPC-enabled privacy-preserving training is also addressed by approaches like the one in [175], which provides active security even with a majority of malicious parties. Systems like Helen [176] demonstrate practical applications of maliciously secure cooperative learning for linear models.
V-A3 Differential Privacy Techniques
In data analysis and machine learning, DP offers a robust framework for safeguarding individual privacy by introducing carefully calibrated noise into data or computations [177]. This noise injection allows meaningful insights to be extracted while preserving privacy, making DP particularly relevant in edge-cloud collaborative AI systems. Specifically, DP mechanisms ensure negligible impact on analysis outputs from the presence or absence of a single individual’s data, preventing identification or inference of sensitive information [178, 179]. Common approaches include adding noise directly to raw data before analysis or perturbing the results of computations, such as model parameters, with the noise calibrated based on the computation’s sensitivity [178, 179]. For instance, in federated learning, DP protects individual device data during collaborative training by adding noise to local model updates before aggregation, preventing private data reconstruction from shared updates [180]. Similarly, in collaborative inference, DP can protect intermediate outputs and mitigate membership information leakage [181]. While noise addition inevitably impacts data utility and model accuracy [182], a key challenge is balancing privacy and utility, especially given the heterogeneous nature of edge devices and distributed edge-cloud systems [183, 184].
V-A4 Data Anonymization Approaches
In AI-driven ECCC environments, data anonymization techniques are crucial for maintaining privacy by transforming data to prevent individual identification [185]. Such environments, characterized by distributed data processing and model training, present unique privacy challenges, necessitating anonymization to safeguard sensitive information while enabling collaborative AI development. Several approaches achieve this goal, including StyleID, which anonymizes faces in image datasets through identity disentanglement [186]. Specifically, StyleID projects images onto a GAN latent space, manipulating specific features to protect identities while preserving other image characteristics. Another key technique, DP, adds noise to data or computations to protect individual privacy while preserving useful insights [79]. For instance, differentially private federated tensor completion enables secure cloud-edge collaborative AIoT data prediction [79]. Dynamic anonymization techniques, often leveraging machine learning, adapt to specific contexts and data characteristics for more flexible and effective privacy protection [185]. Data anonymization in edge-cloud collaborative AI systems, where data may be distributed across multiple devices and locations, can be combined with other privacy-preserving techniques. These techniques include SMPC and federated learning to create a more comprehensive privacy framework [79]. As an example, in cloud-edge collaborative inference, anonymization complements methods like PrivaScissors, which reduces mutual information leakage between intermediate model outputs and device data [181].
V-B Model Security
TABLE VI: Summary of model integrity protection, attack mitigation, and secure update mechanisms.
Method Type Domain Performance Metrics Edge Cloud ECC Privacy Security
Li et al. [187] Data Auditing Next-Generation Consumer Electronics Data Integrity, Privacy
Alaverdyan et al. [188] Security Framework Edge-Cloud Networks Security, Data Protection
Chi et al. [189] Trusted Architecture AIoT Model Performance, Security
FlexibleFL [81] Defense Framework Cloud-Edge Federated Learning Defense Performance, Model Quality
Yao et al. [190] Data Sharing Terminal-Edge-Cloud Network Data Security, Access Control, Integrity
Yang et al. [79] Privacy Framework Cloud-Edge AIoT Data Prediction Accuracy, Privacy
Joshi et al. [191] Edge-AI Framework Landslide Monitoring Response Latency, Energy Consumption, Prediction Accuracy
Khuen et al. [73] Federated Learning Precision Aquaculture Bandwidth Usage, Data Privacy
Nguyen et al. [192] Resource Allocation Space-Air-Ground Integrated Networks (SAGIN) Energy Consumption, Task Delay
1D CNN-LSTM [193] Machine Learning Smart Education Systems Learning Effectiveness, Content Personalization
Shi et al. [194] Resource Management IIoT Energy Consumption, Processing Accuracy, Service Delay
Zeng et al. [195] QoS Testing Satellite Edge Clouds QoS Metrics for Microservices
RLDA [196] Data Aggregation Edge-Cloud IoT Systems Overhead with Growing Dropped Users, Security
Al-Hawawreh et al. [197] Security Framework IIoT Attack Detection Efficiency
Chen et al. [198] Security Framework Distributed Cyber-Physical Systems Attack Detection Efficiency, Resource Utilization
Fan et al. [199] Intrusion Detection Vehicular Networks Attack Detection Accuracy
GS-SNC [200] Secure Data Transmission AIoT Security, Computational Complexity, Encryption Capacity
Ke et al. [201] Reputation Assessment Edge VANETs Precision, Recall
Li et al. [202] Attack Detection Distribution Networks Detection Accuracy
ERLG [203] Game Theory Sensor Edge Cloud Alarm Message Reduction
Qiu et al. [204] Service Chain Provisioning Edge Intelligence Cloud Throughput, Deployment Cost
CD-DOACIR [205] Access Control Cloud-Edge IoT Effectiveness, Efficiency
Yu et al. [206] Game Theory 5G Distributed Networks Defense Effectiveness, Resource Allocation
KubeEdge-Sedna [207] Lifelong Learning Edge AI Adaptability to Heterogeneous Data
Liu et al. [208] Blockchain Framework Edge AIGC Product Protection, Trustworthiness
Abdelmoniem [209] Model-Centric Design Edge AI/ML Scalability, Collaboration Efficiency
TMP [210] Confidential Inference Cloud AI Service Information Leakage Prevention, Performance
Tidjon and Khomh [211] Framework AI-based Systems Trust Properties, Verification
Hutiri and Ding [212] Unified Framework Voice-activated Edge Services Privacy, Security, Fairness
CHAI-T [213] Process Framework Human-AI Collaboration Trust Evolution, Team Processes
U-Federator [69] Authentication Cloud, Edge, and Fog Computing Authentication Time
Hutiri et al. [212] Framework Edge Intelligence, Voice-activated Services Privacy, Security, Fairness
Distributed and collaborative nature of edge-cloud systems introduces significant security vulnerabilities for the AI models deployed within them, potentially exposing them to unauthorized modifications, adversarial manipulations, or inference attacks. Consequently, establishing robust security measures throughout the model lifecycle is paramount for ensuring trustworthy and reliable AI operations in ECCC environments. We explore critical security challenges through three key dimensions: model integrity, adversarial attack mitigation, and secure model updates. Initially, Sec. V-B1 investigates techniques to protect model integrity from tampering and corruption. Following this, Sec. V-B2 analyzes methods to detect and counter adversarial attacks and model poisoning that aim to degrade performance or manipulate behavior. Concluding the discussion, Sec. V-B3 focuses on secure protocols for model updates across the edge-cloud continuum, ensuring authenticity and minimizing vulnerabilities. To further illustrate these security mechanisms, Table VI compares recent approaches in ECCC systems. Notably, the table reveals a wide range of domains—from AIoT to vehicular networks—alongside consistent emphasis on privacy and security, with methods like FlexibleFL [81] demonstrating robust defense performance in federated learning scenarios.
V-B1 Model Integrity Protection
Model integrity protection is crucial in edge-cloud collaborative AI systems to ensure that models are not tampered with or corrupted, either accidentally or maliciously [81]. This protection encompasses various methods, focusing on both the storage and the usage of the AI models. Secure model storage mechanisms aim to safeguard models from unauthorized access and modification while at rest. Cryptographic hashing, for instance, can be employed to generate a unique fingerprint of the model [187]. Any subsequent alteration to the model will result in a different hash value, thereby enabling the detection of tampering. Furthermore, secure enclaves and trusted execution environments (TEEs) offer hardware-based isolation and protection for model storage and execution, preventing unauthorized access and modification [188]. Beyond secure storage, ensuring the integrity of models during operation is equally critical. Techniques such as code signing and model authentication can verify the origin and authenticity of models before they are loaded and executed [189]. Moreover, runtime monitoring can detect anomalies in model behavior, potentially indicating a compromise or corruption [190]. These methods, collectively, contribute to establishing a robust security framework for protecting the integrity of AI models in edge-cloud environments.
V-B2 Attack Detection and Prevention
AI models in edge-cloud systems are vulnerable to attacks, necessitating robust defense mechanisms. Adversarial attacks introduce subtle perturbations to input data, causing misclassification or incorrect outputs [199]. Detecting such attacks involves analyzing input data for anomalies; meanwhile, prevention strategies include adversarial training with adversarial examples and input sanitization to neutralize perturbations. Auto-updating mechanisms can further enhance adaptability to evolving attack strategies [199]. Model poisoning attacks, conversely, inject malicious data into training datasets, manipulating model learning and creating exploitable backdoors [81]. Detecting these attacks requires careful examination of training data for inconsistencies. Federated learning offers inherent protection by distributing training [79], and contribution-based mitigation strategies in frameworks like FlexibleFL [81] enhance resilience by evaluating uploaded model parameters. Secure network coding schemes [200] can protect data transmission, while trust management mechanisms [201] ensure model and update reliability.
V-B3 Secure Model Updates
Maintaining the integrity and trustworthiness of AI models in edge-cloud collaborative systems necessitates secure model updates [214]. Authentication mechanisms are essential for verifying the update source and preventing unauthorized modifications [215]. Specifically, digital signatures, cryptographic hashes, or blockchain-based solutions [208] can be employed for this purpose. Additionally, verification of update integrity ensures that updates remain untampered during transmission or storage [187], leveraging techniques such as checksums, message authentication codes (MACs), or Merkle trees. Secure transmission through encrypted channels using protocols like TLS or SSH [210] protects against eavesdropping. A trusted execution environment (TEE) [210] can further secure the update process itself from runtime attacks. Rollback mechanisms to a previous stable state in case of faulty updates are also crucial [207].
V-B4 Trust Management Mechanisms
In edge-cloud collaborative systems, ensuring the reliability and trustworthiness of AI models and their updates is crucial [216]. Specifically, this necessitates establishing mechanisms for evaluating and maintaining trust in the components and processes involved in model development, deployment, and operation. Given the distributed and collaborative nature of these environments, trust management must also address potential risks associated with interactions between edge devices, cloud servers, and the AI models themselves. Reputation systems are vital for establishing trust within collaborative AI environments [211]. They track and aggregate feedback on the performance and behavior of different entities, providing a basis for assessing trustworthiness. For example, an edge device’s reputation could be based on its history of providing accurate data and reliably executing tasks, while an AI model’s reputation could reflect its accuracy, fairness, and robustness [217]. These reputation scores subsequently inform decisions about trust allocation, resource distribution, and task assignment. Trust evaluation methods quantify and assess the trustworthiness of AI models and other system components [217]. Such methods can involve quantitative metrics like accuracy, latency, and energy efficiency, combined with qualitative assessments such as user feedback and expert reviews [213]. Context-specific application requirements also influence trust evaluation. For instance, safety-critical applications, like autonomous driving, prioritize robustness and reliability [218], whereas applications like personalized healthcare prioritize privacy and fairness [212]. Robust and context-aware trust evaluation is therefore crucial, including consideration of the trustworthiness of AI engineering schemes themselves, as exemplified by platforms like KubeEdge-Sedna [207].
V-C Communication Security
Securing communication channels is paramount in edge-cloud collaborative systems to protect data transmission from unauthorized access and tampering. This protection involves secure data transmission methods, including encryption and secure communication protocols, safeguarding data in transit. Authentication and access control mechanisms, such as multi-factor authentication and role-based access control, prevent unauthorized access to resources and data. Various encryption schemes, including symmetric and asymmetric encryption, secure data storage and transmission in edge-cloud environments. Additionally, blockchain technology offers potential for enhancing security by providing decentralized security, data integrity, and transparent audit trails.
V-C1 Secure Data Transmission
Protecting sensitive information in transit between edge devices and cloud servers requires robust security mechanisms in ECCC systems [219]. Various encryption methods serve critical roles in securing data transmission, with symmetric encryption offering efficiency for large data transfers through a single key approach, while asymmetric encryption enhances security for key exchange by employing separate encryption and decryption keys [220, 221]. Secure communication protocols including transport layer security (TLS) and its predecessor secure sockets layer (SSL) establish protected channels that ensure data confidentiality and integrity by utilizing cryptographic algorithms to encrypt data during transmission [222]. Additional security layers can be implemented through virtual private networks (VPNs) that create secure connections over public networks [223], while blockchain technology provides decentralized and immutable frameworks for creating transparent audit trails that prevent unauthorized modifications and strengthen the security of edge-cloud communication [190, 188].
V-C2 Authentication and Access Control
In edge-cloud collaborative systems, robust security is paramount due to their distributed nature and the potential involvement of resource-constrained devices. Specifically, authentication and access control mechanisms are essential for preventing unauthorized access to data and resources [224]. Without these safeguards, malicious actors could compromise sensitive data or disrupt system operations, while inadequate access control could lead to data breaches or system instability due to authorized users accessing resources beyond their permitted scope. Authentication verifies user or device identity attempting system access, while access control governs permitted resources and actions [205]. This combined approach is crucial for maintaining security and integrity in these environments. For example, DOACIR [205] and P-MOD [225] offer enhanced security through controlled data sharing and hierarchical privilege management, respectively. Multi-factor authentication (MFA) strengthens security by requiring multiple forms of identification, mitigating risks even if one factor is compromised. Additionally, role-based access control (RBAC) [226] streamlines access management by assigning permissions to roles, improving efficiency and reducing errors, especially in complex systems. Integrating blockchain technology can further enhance RBAC by providing a secure platform for managing access control policies [226].
V-C3 Encryption Schemes
Robust encryption schemes are essential for protecting data both in transit and at rest within edge-cloud collaborative environments. These schemes are broadly classified as symmetric and asymmetric encryption, each possessing distinct advantages and applications. Symmetric encryption uses a single key for both encryption and decryption, offering high performance suitable for securing large data volumes, such as data streams between devices and edge nodes [190]. This efficiency is particularly beneficial for resource-constrained edge devices, although secure key distribution and management present challenges. In contrast, asymmetric, or public-key, encryption utilizes separate keys for encryption and decryption, simplifying key distribution and facilitating secure communication across the edge-cloud hierarchy [188]. Asymmetric encryption also enables fine-grained access control, as exemplified by CP-ABE schemes, where data decryption is contingent on possessing specific attributes [205]. While offering enhanced security, asymmetric encryption typically incurs higher computational overhead. Consequently, edge-cloud systems often combine symmetric and asymmetric encryption to leverage the strengths of both, potentially integrating hybrid models with blockchain technology for enhanced data integrity and auditability in cloud storage [190].
V-C4 Blockchain-based Security Solutions
We explore how blockchain technology enhances security in edge-cloud collaborative environments through decentralized architecture and immutable data records. In edge-cloud collaborative systems, blockchain offers a promising approach to enhancing security [227]. Its decentralized nature, based on cryptographic principles, distributes security responsibility across a network of nodes, eliminating the single point of failure vulnerability inherent in centralized systems [228]. Consequently, this distributed architecture enhances resilience against attacks and ensures system availability. Blockchain’s immutability guarantees data integrity, preventing unauthorized modifications and ensuring data provenance [229, 230]. Specifically, once a transaction is recorded on the blockchain, it becomes tamper-proof. Additionally, blockchain facilitates transparent audit trails by recording all transactions on a public ledger [231]. This characteristic enables verifiable tracking of data flow and system operations, promoting accountability and facilitating efficient auditing for security and compliance [190].
VI Applications
VI-A Autonomous Driving
Real-time processing of complex sensor data is crucial for safe and efficient autonomous driving. Edge-cloud collaboration offers a promising solution by distributing computational workloads between resource-constrained in-vehicle edge devices and powerful cloud servers. Specifically, edge devices can handle time-sensitive tasks like object detection and perception, potentially leveraging lightweight frameworks like Edge YOLO [232], which combines pruned and compressed networks for efficient multi-scale prediction on resource-limited platforms. In addition, the collaborative nature of these systems enables integration of crowdsourced data, enhancing perception capabilities by providing a more comprehensive environmental view, as demonstrated by LiveMap [233, 234], which efficiently processes and integrates multi-vehicle data for real-time dynamic map generation. Beyond perception, LLMs are emerging as powerful tools for motion planning and decision-making [235]. For example, systems like EC-Drive utilize LLMs across both cloud and edge, adapting to data drifts and optimizing communication for reduced latency. Similarly, collaborative policy learning through DRL [233] offers a mechanism for dynamic scheduling and resource allocation, optimizing vehicle scheduling and offloading decisions for improved system robustness and adaptability. Adaptive edge-cloud collaboration, aided by large vision models like in LAECIPS [236], further enhances performance and adaptability of IoT-based perception systems through efficient hard input mining. Concurrently, the fusion of DRL and edge computing facilitates real-time monitoring and control optimization in IoT environments [18], enabling predictive system state analysis and dynamic resource allocation for enhanced control.
VI-B Smart Healthcare
Figure 7: Integration of federated learning, edge-cloud platforms, and Internet of medical things (IoMT) optimization in smart healthcare.
Advancements in AI and edge-cloud computing are transforming smart healthcare systems into efficient, secure, and personalized ecosystems, as shown in Fig.7. We explore how AI-driven edge-cloud collaboration revolutionizes healthcare delivery and improves patient outcomes by investigating FL for medical image analysis, edge-cloud platforms for remote patient monitoring, and resource allocation optimization for Internet of medical things (IoMT) devices. FL offers a promising technique for training AI models on distributed medical datasets without compromising patient privacy, allowing collaborative model training across multiple healthcare providers while keeping sensitive data localized [237]. Specifically, FL enables robust diagnostic model development, including COVID-19 diagnosis using multimodal data like X-rays and ultrasound images [238]. Edge-cloud platforms are crucial for enabling remote patient monitoring and personalized healthcare [239], leveraging edge device computation and cloud storage for real-time monitoring and personalized interventions. For example, wearable sensors and IoMT devices collect patient data, processed at the edge for immediate feedback and transmitted to the cloud for analysis and storage [240]. AI algorithms deployed on edge devices analyze this data, providing personalized recommendations and improving intervention effectiveness [241]. Efficient resource allocation is also essential for optimizing IoMT device performance and ensuring timely healthcare delivery [242]. Consequently, AI-driven task offloading and resource management dynamically allocate tasks and resources between edge and cloud based on real-time demands and availability, minimizing latency and improving system efficiency. Hierarchical FL approaches further enable multi-party collaboration and efficient model training for anomaly detection in IoMT data, leveraging digital twins and edge cloudlets for real-time patient health insights and timely interventions [243]. Integrating FL within the healthcare Metaverse enhances these capabilities by improving privacy, scalability, and interoperability [244].
VI-C Industrial Automation
The integration of ECCC delivers significant advantages to industrial automation processes. Edge intelligence enables real-time fault diagnosis and predictive maintenance [18], while anomaly detection techniques deployed on AIoT devices enhance system reliability by identifying unusual patterns and potential failures before they cascade throughout the system [243, 16]. Leveraging both cloud computational power and edge real-time processing capabilities, industrial automation systems achieve greater agility through optimized production processes and resource allocation [63], where cloud-based platforms analyze historical data to generate optimized production schedules while edge devices adapt these schedules in real-time based on local conditions. Distributed intelligence across multiple edge devices handles complex manufacturing tasks more efficiently and enhances fault tolerance since individual device failures do not necessarily compromise the entire system [245]. As edge intelligence, anomaly detection, cloud-edge collaboration, and distributed intelligence converge, industrial automation systems evolve toward more sophisticated and autonomous manufacturing processes [49].
VI-D Smart Cities
ECCC enhances urban services in smart cities by enabling intelligent traffic management and environmental monitoring [246, 128]. Edge devices collect and analyze real-time data for adaptive traffic control and congestion reduction, while distributed sensor networks monitor environmental factors like air quality and noise levels. Subsequently, cloud platforms aggregate and analyze this data to inform urban planning and environmental protection. In addition, AI-driven resource allocation optimizes smart grid energy consumption [247]. Specifically, edge devices monitor energy usage, and cloud-based algorithms dynamically adjust energy distribution to minimize waste and improve grid stability. FL is crucial for privacy-preserving data analysis and decision-making in these applications [13, 248, 249]. Through FL, stakeholders such as government agencies and private companies can collaboratively train models on combined data without sharing sensitive information, leading to more accurate and robust models for urban planning and public safety.
VII Performance Analysis
VII-A Evaluation Metrics
Figure 8: Evaluation metrics, which are categorized into four primary groups: System performance, model efficiency, quality of service, and security and privacy. Each group contains specific evaluation dimensions (e.g., latency, inference time, availability, and data breach frequency), enabling comprehensive assessment of collaborative performance, resource usage, user experience, and trustworthiness.
Performance evaluation of edge-cloud collaborative systems requires a comprehensive assessment across multiple dimensions. We categorize the relevant metrics into four primary groups: system performance, model efficiency, quality of service, and security and privacy. System performance metrics. Fig. 8 shows comprehensive evaluation spans system-level, model-level, service-level, and trust-level metrics, considering of real-world constraints like network variability, compute heterogeneity, and regulatory demands.
VII-A1 System Performance
Throughput, latency, and scalability are crucial metrics for evaluating the effectiveness and responsiveness of ECCC systems [91, 17, 250]. Throughput measures the volume of data processed or tasks completed within a given timeframe, with higher throughput indicating efficient handling of substantial workloads [251]. Conversely, latency quantifies the delay between request initiation and completion; minimizing latency is paramount for real-time and interactive applications where responsiveness is critical [250]. Scalability, in turn, refers to a system’s ability to adapt to increasing workloads or expanding infrastructure without performance degradation, accommodating growth in user demands or data volume while maintaining acceptable throughput and latency [227]. The distributed and heterogeneous nature of edge-cloud environments amplifies the significance of these metrics, as the interplay between edge and cloud resources necessitates careful optimization. For example, the study in [91] benchmarks a container-based edge computing system, demonstrating the impact of data synchronization on scalability across different deployment strategies. Similarly, the SPEC-RG reference architecture in [17] provides a framework for analyzing resource management and performance, while [250] highlights the potential for performance inversion due to edge queuing delays. Evaluating these metrics is essential for understanding system efficiency and informing design choices for optimized performance in resource-constrained and dynamic edge-cloud environments [46].
VII-A2 Model Efficiency
Efficient AI models are essential for effective deployment and operation within the resource-constrained edge-cloud continuum. Key metrics for quantifying model efficiency include computational cost, memory usage, and inference time, informing model selection and deployment strategies. Computational cost, typically measured in FLOPS [252], quantifies processing demands. Higher FLOPS values indicate greater computational requirements and potentially increased energy consumption, impacting resource-limited edge devices and overall system sustainability [253, 254]. Memory usage reflects a model’s storage footprint; large models may exceed the limited capacity of edge devices [255], necessitating compression techniques like pruning, quantization, or knowledge distillation [256]. Inference time, the duration for processing input and generating output, directly impacts application responsiveness [255]. In latency-sensitive applications like autonomous driving or real-time video analytics, minimizing inference time is crucial [17]. This optimization often involves balancing model accuracy and computational efficiency, potentially employing strategies like model partitioning across different devices [253, 257].
VII-A3 Quality of Service
Evaluating the performance of edge-cloud collaborative systems necessitates an analysis of QoS metrics, which directly reflect the system’s capacity to fulfill user expectations and provide consistent, dependable service [195]. Key QoS metrics include availability, reliability, and user satisfaction. Availability quantifies the proportion of time a system remains operational and accessible, a crucial aspect in edge-cloud systems where interruptions significantly impact user experience and application performance [17]. High availability relies on robust system design, redundant infrastructure, and effective fault tolerance mechanisms, further enhanced by dynamic resource management and adaptive optimization techniques [258]. Reliability, in turn, assesses the system’s ability to consistently deliver the expected service level over time, encompassing fault tolerance, error recovery, and performance maintenance under varying workloads [259]. Dependency-aware offloading and collaborative resource management contribute to improved reliability by optimizing task placement and resource allocation. While often subjective, user satisfaction remains a crucial QoS indicator, reflecting the overall user experience, including performance, ease of use, and perceived value [260]. In edge-cloud systems, factors such as latency, throughput, and the quality of AI-driven services influence user satisfaction. Specifically, QoE serves as a proxy for user satisfaction, directly linked to QoS parameters like latency and loss in applications like video streaming [261]. Similarly, in industrial automation, the relationship between QoS and QoE is critical for optimizing network resource management and ensuring customer satisfaction.
VII-A4 Security and Privacy
In edge-cloud collaborative systems, robust security and privacy are crucial for maintaining trust and protecting sensitive information, especially given the sensitive nature of processed data [262]. Therefore, evaluating these systems requires metrics focusing on data breaches, encryption strength, and regulatory compliance. Data breach frequency directly indicates system vulnerability by quantifying unauthorized accesses within a given timeframe [221]. A lower frequency suggests stronger security, and analyzing breach characteristics can inform improved protocols, potentially leveraging frameworks like the transparency, accountability, and blockchain framework [262]. Encryption strength assesses cryptographic robustness, including key length, algorithms, and resistance to attacks. Strong encryption is paramount in distributed edge-cloud environments [190], and combining techniques like lightweight symmetric encryption and ciphertext-policy attribute-based encryption can enhance security [190]. Additionally, compliance with regulations like GDPR and CCPA is essential. Metrics in this area assess adherence to data anonymization, user consent, and data retention policies [263]. Evaluating privacy protectability, along with techniques like PrivaScissors, can further enhance privacy protection [264, 181].
VII-B Benchmarking Methods
Performance benchmarking of edge-cloud collaborative systems requires standardized evaluation approaches to facilitate meaningful comparisons across diverse system designs and optimization strategies. Standardized datasets and workloads, discussed in Sec. VII-B1, provide a common basis for evaluation, minimizing variability stemming from data discrepancies. Well-defined testing environments, explored in Sec. VII-B2, further enhance reliability and reproducibility by controlling hardware and software configurations, network setups, and other influential environmental factors. Sec. VII-B3 introduces performance comparison frameworks, offering structured methodologies and tools for identifying performance bottlenecks and areas for improvement in edge-cloud systems. Additionally, Sec. VII-B4 examines various evaluation methodologies, including simulation, emulation, and real-world testing, analyzing the strengths and limitations of each to guide researchers in selecting the most appropriate approach for their specific benchmarking requirements.
VII-B1 Standard Datasets and Workloads
Consistent and comparable performance evaluation across different studies of ECCC systems requires the use of standard datasets and workloads [265]. Such standardization clarifies the strengths and weaknesses of various architectures, optimization techniques, and algorithms, enabling researchers to build upon existing work and accelerate the development of innovative solutions. For example, Edge AIBench [266] offers benchmarks modeling typical edge scenarios like ICU patient monitoring and autonomous vehicles, focusing on data distribution and workload collaboration across client devices, the edge, and the cloud for comprehensive end-to-end performance evaluation. Similarly, EdgeBench [267] provides a workflow-based benchmarking approach, allowing customization of workflow logic, data storage, and the distribution of workflow stages across computing tiers to tailor benchmarks to specific application scenarios. Appropriate workload selection is also crucial [17]. Specifically, workloads should reflect real-world application characteristics, including data intensity, computational complexity, and communication patterns, particularly in distributed intelligence scenarios where data and model distribution, as well as communication overhead for model training and inference, are critical factors. KheOps [268] further emphasizes reproducibility in edge-to-cloud experiments by providing a collaborative environment with an experiment repository, a notebook environment, and a multi-platform methodology, promoting transparency and rigor in performance evaluation. The availability of standard datasets and workloads also facilitates the development of new optimization techniques [269], allowing researchers to evaluate their methods on established benchmarks [257].
VII-B2 Testing Environments
Robust benchmarking of edge-cloud collaborative systems mandates controlled testing environments [270]. Specifically, accurately capturing performance characteristics requires careful configuration and control due to the complex interactions among edge devices, network infrastructure, and cloud servers. This process involves meticulously configuring hardware, software, and network parameters to minimize external influences and ensure reproducibility. The diverse hardware platforms in edge deployments, from resource-constrained devices to powerful edge servers, necessitate standardized testing environments for fair comparisons across different system architectures [266]. For example, BenchFaaS and KheOps employ heterogeneous testbeds, encompassing virtual machines, Raspberry Pis, and various cloud platforms, to evaluate performance under diverse hardware and network conditions [271, 268]. Additionally, accurate measurement of key performance indicators (KPIs), such as latency, throughput, and energy consumption, requires precise control over network conditions, including bandwidth, latency, and packet loss. Standardized software configurations, including operating systems, libraries, and runtime environments, are also essential for eliminating inconsistencies and ensuring comparability across experiments. Representative benchmarks that accurately reflect real-world scenarios, encompassing data volume, processing requirements, and communication patterns, are necessary for comprehensive performance evaluation.
VII-B3 Performance Comparison Frameworks
Benchmarking edge-cloud collaborative systems relies on performance comparison frameworks, which provide structured methodologies for evaluating diverse architectures and optimization techniques. Such frameworks offer a standardized approach to performance assessment, enabling researchers and practitioners to identify bottlenecks and derive best practices. Inspired by frameworks like OpenPerf for benchmarking open-source projects [272] and a cloud infrastructure evaluation framework [265], a robust framework for edge-cloud systems should encompass diverse workloads and metrics, facilitating systematic analysis of system configurations and algorithms for improved designs and optimized deployments. Specifically, a comprehensive framework should define a multifaceted set of evaluation metrics encompassing system performance (throughput, latency, scalability), model efficiency, quality of service, and security and privacy [273]. Additionally, it should specify standardized datasets and workloads representative of real-world applications [270], alongside guidelines for establishing controlled testing environments for reproducible results [274]. Finally, incorporating tools and methodologies for result analysis, including statistical analysis and visualization techniques [17], allows for identifying performance bottlenecks and understanding design trade-offs.
VII-B4 Evaluation Methodologies
Benchmarking ECCC systems relies on diverse methodologies, each with inherent strengths and limitations. These methodologies are broadly categorized into simulation, emulation, and real-world testing. Simulation creates a virtual model of the entire system, encompassing hardware, network characteristics, and workloads [17]. This approach offers cost-effectiveness and flexibility, enabling exploration of various scenarios and configurations without physical deployment. However, reliance on simplified models and assumptions may not accurately reflect real-world complexities, potentially leading to discrepancies between simulated and actual performance. Specifically, accurately modeling dynamic network conditions or heterogeneous edge devices remains challenging. Emulation deploys real software components on virtualized infrastructure, often incorporating realistic network emulation tools [121, 275]. Consequently, this approach bridges the gap between simulation and real-world testing by providing a more realistic yet controlled environment. Researchers can emulate network characteristics like latency, bandwidth, and packet loss using tools like NetEm or eBPF-based solutions [275], enabling more accurate performance assessments. Nevertheless, emulation still introduces virtualization overhead and may not fully capture hardware-software interactions; additionally, scaling network emulation for large deployments presents significant challenges [275]. Real-world testing deploys the system on physical infrastructure [265, 91], offering the most accurate performance assessment by capturing the full system complexity and real-world interactions. This approach is essential for validating simulation and emulation findings and demonstrating solution practicality. However, real-world testing can be expensive, time-consuming, and difficult to control, especially for large-scale distributed systems. Ensuring reproducibility and replicability also remains challenging due to variations in hardware, software, and environmental factors [268]. Tools and platforms like KheOps [268] aim to facilitate reproducible experiments across different infrastructures, partially addressing these challenges.
VII-C System Design Evaluation
We analyze the performance of various ECCC paradigms by comparing different architectural approaches and exploring their strengths and weaknesses across diverse application scenarios. Additionally, we examine the effectiveness of optimization techniques, such as load balancing, resource allocation, and task scheduling, for enhancing system performance. Inherent trade-offs in ECCC, including the balance between performance and energy consumption, latency and accuracy, and security and usability, are also discussed. Specifically, Sec. VII-C1 compares different edge-cloud architectures and their suitability for various applications. Sec. VII-C2 then analyzes how various optimization techniques impact system performance and efficiency. Following this, Sec. VII-C3 explores the inherent trade-offs in edge-cloud systems, providing insights into design considerations.
VII-C1 Comparison of Different Architectures
Edge-cloud collaborative architectures balance centralized and decentralized processing, impacting performance across various applications. Cloud-centric architectures, relying heavily on cloud resources, offer scalability [276], but may introduce latency, especially in real-time applications [250]. In contrast, edge-centric architectures, such as Armada [276], prioritize local processing, minimizing latency but facing limitations in computational resources and data synchronization. This trade-off is further explored in [91], which highlights the performance benefits of localized edge processing and the scalability of cloud processing for synchronized applications. Hybrid architectures combine both paradigms. For example, AppealNet [115] uses predictive offloading to optimize accuracy and efficiency, while EdgeMatrix [277] employs a networked multi-agent actor-critic algorithm for resource management and service orchestration. Similarly, ACE [38] advocates for a unified platform to manage edge and cloud resources transparently. Architectural choices also influence model optimization. Model compression techniques suit resource-constrained edge devices [115]. Federated learning enables collaborative training across distributed edge devices while preserving privacy [38]. The SPEC-RG reference architecture [17] provides a framework for analyzing these techniques, and reinforcement learning approaches are increasingly used for dynamic resource management in complex edge-cloud systems [105].
VII-C2 Analysis of Optimization Techniques
Load balancing, resource allocation, and task scheduling are essential optimization techniques in edge-cloud collaborative systems for enhanced performance and efficiency [278]. Load balancing distributes workloads across edge devices and the cloud to mitigate congestion and improve resource utilization, as exemplified by the two-timescale optimization framework in [120] for balancing edge and cloud loads. Similarly, [118] introduces a collaborative approach for service placement and task scheduling, incorporating both edge-cloud and edge-edge cooperation for improved task offloading. Resource allocation strategies optimize the use of computational resources, bandwidth, and storage capacity. Dynamic resource allocation adapts to real-time demands and availability [163], while collaborative resource management coordinates allocation across multiple edge nodes and the cloud [19], with reinforcement learning emerging as a promising approach for adaptive resource management [46]. Finally, task scheduling determines task execution order and timing to minimize latency and improve system performance. Latency-aware and energy-efficient offloading mechanisms prioritize tasks with deadlines or minimize energy consumption [120], while dependency-aware offloading considers task dependencies for complex applications [118].
VII-C3 Trade-off Analysis
Optimizing ECCC systems requires careful consideration of various trade-offs inherent in their design. A primary concern is balancing performance, often measured by throughput and processing speed, against energy consumption [128]. While deploying computationally intensive tasks at the edge minimizes latency and improves responsiveness, it can strain the limited energy resources of edge devices; conversely, offloading to the cloud conserves edge energy but introduces network latency and potential bandwidth bottlenecks, particularly affecting latency-sensitive applications [279]. Achieving an optimal balance necessitates sophisticated resource allocation and task offloading strategies [129]. Another key trade-off lies between latency and accuracy, especially in AI-driven applications. Complex, accurate AI models demand substantial computational resources, potentially increasing latency, especially on edge devices, whereas simpler models offer lower latency but potentially reduced accuracy. Model compression techniques, such as pruning, quantization, and knowledge distillation, aim to mitigate this trade-off [135], but the optimal balance remains application-dependent. For example, real-time applications like autonomous driving prioritize minimal latency [250], potentially accepting slightly lower accuracy, while applications like medical diagnosis prioritize accuracy even at the cost of increased latency. Additionally, security and usability present another trade-off. Robust security measures, like encryption and authentication, protect data integrity but can introduce complexity and overhead, impacting usability. Specifically, complex authentication procedures, while enhancing security, can increase user login time, and data encryption adds computational overhead, potentially affecting application performance. The distributed nature of these systems further complicates security, particularly in managing trust and secure communication [46].
VIII Challenges and Future Directions
VIII-A Technical Challenges
Figure 9: Technical challenges, which span across model optimization, system integration, and performance bottlenecks, with domain-specific requirements shaping solution design.
Edge-cloud system faces key technical hurdles that impede its widespread adoption and optimal performance. As illustrated in Fig. 9, the deployment of AI in ECCC systems faces multifaceted technical challenges, including model optimization under resource constraints, integration of heterogeneous hardware/software, overcoming performance bottlenecks, and addressing domain-specific concerns such as real-time processing and scalability. Model optimization for the unique resource constraints of edge-cloud environments requires careful consideration of the trade-offs between accuracy, efficiency, and scalability. Additionally, integrating diverse hardware and software components necessitates effective heterogeneity management, seamless interoperability, and standardized protocols. Performance bottlenecks, including communication overhead, computational and storage limitations, and the demands of real-time processing, further challenge the efficiency and responsiveness of these distributed systems. Addressing these obstacles is crucial for unlocking the full potential of this promising paradigm.
VIII-A1 Model Optimization Issues
Optimizing AI models for deployment in edge-cloud collaborative systems presents significant challenges in balancing model accuracy with computational efficiency given the resource constraints of edge devices [280]. Specifically, complex and accurate models often demand substantial resources, unsuitable for edge deployment, while smaller models may sacrifice accuracy for performance [280]. Resource limitations across the edge-cloud continuum, including limited processing power, memory capacity, and energy availability, especially for IoT devices [128, 281], further complicate model deployment. Additionally, scaling AI models across distributed systems introduces complexities in maintaining performance and efficiency with growing system size and data volumes [12]. Such scaling challenges are exacerbated by communication overhead, resource contention, and device heterogeneity [282, 283]. Therefore, effective model optimization necessitates techniques like pruning, quantization, knowledge distillation [32], and resource-aware strategies like dynamic model selection and offloading [46, 23], while also considering energy-aware scheduling [284] and adaptive model partitioning [207] to address resource constraints and scalability limitations in these dynamic environments. The growing use of LLMs further emphasizes the need for efficient inference and deployment techniques on resource-constrained edge devices [285]. For example, Shoggoth [286] utilizes online knowledge distillation to enhance accuracy while managing data drift and offloading computationally intensive tasks.
VIII-A2 System Integration Challenges
Effectively integrating the diverse hardware and software of edge-cloud collaborative systems presents significant challenges, demanding effective heterogeneity management, robust interoperability solutions, and standardized protocols [287]. Managing heterogeneous hardware and software across resource-constrained edge nodes and powerful cloud servers requires addressing variations in processing capabilities, memory capacities, network bandwidths, and operating systems [254]. Ensuring seamless communication and data exchange between these disparate components is crucial, as differing data formats, communication protocols, and software interfaces can create interoperability issues [288]. Efficient workload orchestration across heterogeneous resources necessitates sophisticated scheduling algorithms that consider each device’s capabilities and limitations [121], while application portability demands flexible deployment mechanisms. Approaches like containerization (e.g., Docker [287]) and WebAssembly  [288] offer portability solutions. AI-driven techniques [289, 290] aim to optimize task placement and resource management. Additionally, interoperability challenges arise from the heterogeneous mix of devices with distinct operating systems, communication protocols, and data formats [37]. Variations in data representation and communication protocols can lead to compatibility issues, requiring complex data transformations [290]. Standardized communication protocols and well-defined interfaces are essential for seamless data exchange, simplified application development, and enhanced performance and scalability [258, 289]. Such standardization promotes interoperability between vendor solutions, fostering a more open ecosystem [277, 291]. Standard protocols ensure consistent and reliable communication, facilitating efficient data exchange and coordinated operation across diverse components [288, 218]. Standardized data exchange formats promote interoperability, enabling tasks like distributed model training and real-time analytics [289]. Consequently, adopting standard protocols simplifies system integration, reduces development complexity, and accelerates the deployment of edge-cloud solutions [290, 258]. This is particularly critical for applications like autonomous driving [37] and smart healthcare [128] where reliable, real-time data exchange is paramount [292].
VIII-A3 Performance Bottlenecks
ECCC faces significant performance bottlenecks stemming from resource limitations. Communication overhead introduces latency and bandwidth consumption between edge devices and the cloud, impacting real-time applications [91, 250, 293]. Strategies like optimized task offloading, latency-aware strategies, data compression, and efficient communication protocols can mitigate these issues [294, 48, 279, 128]. Additionally, edge devices, often constrained by limited processing power, memory, and energy, struggle to execute complex AI models and handle large datasets, especially with the growing data volume from IoT devices [50, 295, 296, 297]. This computational bottleneck is further complicated by device heterogeneity [17, 257]. Limited storage capacity at the edge, exacerbated by data volume and device heterogeneity, necessitates efficient data management solutions like intelligent caching, data compression, distributed storage (e.g., EdgeKV [47]), and data lifecycle management [32, 298, 299, 191, 128, 300]. Furthermore, real-time applications, such as autonomous driving, industrial automation, and smart healthcare, demand ultra-low latency [236, 301, 18]. Consequently, future research should focus on minimizing latency through techniques like data compression, model partitioning, and adaptive resource allocation [191, 18].
VIII-B Future Research Directions
A critical challenge lies in scaling AI model deployments across expansive edge-cloud systems while preserving performance and efficiency as device numbers and data volumes surge. Promising research avenues include investigating advanced AI techniques such as neural architecture search, automated machine learning, and large language models. Additionally, exploring AI agents, embodied AI, neuromorphic and quantum edge computing, along with 6G network integration, offers potential for revolutionizing these systems. Equally crucial is improving system-level aspects, including dynamic resource management, adaptive optimization strategies, and robust fault tolerance mechanisms to enhance reliability.
VIII-B1 Advanced AI Techniques
Figure 10: Overview of advanced AI techniques, including NAS, AutoML, LLMs, AI agents, embodied AI, neuromorphic computing, quantum edge computing, and 6G integration.
As illustrated in Fig. 10 shows the summary of advanced AI technologies in future research directions, including NAS, AutoML, and LLMs, contributes to enhancing system efficiency and adaptability. All these modules operate synergistically under the orchestration of an edge-cloud control unit, which facilitates dynamic task allocation, resource optimization, and intelligent decision-making across heterogeneous infrastructures.
$\bullet$ Neural Architecture Search  NAS automates the design and optimization of neural network architectures, aiming to discover models that outperform manually designed ones in accuracy and efficiency for edge-cloud collaborative environments [302]. NAS algorithms explore a vast search space of potential architectures, evaluating their performance to identify architectures tailored to the specific constraints of edge-cloud systems, such as limited resources and bandwidth [98, 245].  Approaches like reinforcement learning, evolutionary algorithms, and gradient-based NAS show promise in finding optimal architectures for specific tasks and resource constraints [42], and combining NAS with model optimization techniques like compression and adaptation further improves performance and efficiency [303, 304].  This integrated approach addresses the challenges of deploying complex AI models in heterogeneous edge-cloud environments [305].
$\bullet$ AutoML Automated machine learning (AutoML) streamlines AI model development and deployment in edge-cloud systems by automating processes like model selection, hyperparameter tuning, and deployment [306, 307]. This automation is crucial in edge-cloud environments where diverse hardware and software coupled with resource constraints make manual optimization complex [308].  For example, techniques like FAST-DAD [309] distill complex models into smaller, deployable ones for edge devices, while frameworks like H2O AutoML simplify model selection and hyperparameter optimization [307].  Language-driven systems like AutoMMLab [310] further democratize AI development.  However, challenges remain in transparency, flexibility for complex scenarios, and consistent coverage of the entire ML workflow [311]. In large-scale environments, sampling-based AutoML and human-centered systems like OmniForce [312, 308] show promise, while automatically generating data-dependent configurations improves model performance [313].
$\bullet$ Large Language Models Optimizing LLM inference for edge devices involves addressing limited resources and real-time performance, necessitating research on optimized inference techniques and model compression methods like Agile-Quant [314, 315].  Approaches like expert-wise bitwidth adaptation in EdgeMoE [314] can further enhance on-device efficiency.  Integrating LLMs for edge-cloud orchestration, including dynamic resource allocation and task offloading, promises more efficient systems, as demonstrated by VELO and EC-Drive [316, 235].  Developing specialized LLM architectures for edge-cloud environments, including low-rank adaptation (NetGPT [317]), model partitioning, and split learning/inference [285], is a critical research direction.
$\bullet$ AI Agents Intelligent agents capable of autonomous decision-making and learning within edge-cloud systems enhance adaptability and efficiency [12, 318].  AI agents facilitate continuous learning on edge devices, addressing bandwidth limitations and connectivity issues through approaches like NeuroEvolutionary learning [303]. Collaborative learning improves model accuracy, particularly in heterogeneous networks [11, 10], and these agents optimize system performance through dynamic task offloading and resource allocation [283, 305].  Integration with lifelong learning frameworks like KubeEdge-Sedna further enhances automation [207].
$\bullet$ Embodied AI Embodied AI, focusing on agents that interact with their environment, requires new algorithms and frameworks tailored for resource-constrained edge devices [319].  Algorithms for robot control and coordination must consider latency and bandwidth limitations [320], while efficient task offloading and resource allocation optimize performance [277].  Exploring embodied AI for multi-robot collaboration and dynamic visual servoing offers significant opportunities [321, 320].  Developing robust and adaptive algorithms that leverage distributed resources is a critical research direction [322].
$\bullet$ Neuromorphic Computing Inspired by the biological brain, neuromorphic computing promises energy-efficient AI models for edge-cloud environments [323, 324]. Neuromorphic hardware, like SNNs, offers potential energy savings compared to traditional hardware [325], while inherent fault tolerance improves robustness [326].  Integrating neuromorphic hardware into data centers presents challenges in algorithms, software, and benchmarks [327].  However, its potential benefits for edge robotics drive research towards ultra-low-power designs and 3D integration technologies [322, 328].
$\bullet$ Quantum Edge Computing Quantum computing offers revolutionary potential for edge-cloud applications by tackling classically intractable problems [329]. Consequently, developing edge-tailored quantum algorithms [330] and robust quantum communication protocols [331], potentially in hybrid approaches, are key research avenues. However, significant security [332] and privacy concerns [333] necessitate quantum-resistant techniques. Investigating the associated societal impacts [14] is also crucial for responsible deployment.
$\bullet$ 6G Network Integration 6G networks, with ultra-low latency and high bandwidth, promise to enhance edge-cloud systems [334].  Low latency enables real-time processing, while high bandwidth facilitates seamless data transfer, supporting resource-intensive AI tasks [335, 336].  Advanced features like network slicing and cell-free MIMO optimize performance [336], and integration with digital twins and pervasive AI unlocks new possibilities [337, 338].  Addressing challenges in network architecture, resource allocation, and security through AI-driven solutions is crucial [339, 340].
VIII-B2 System Improvements
Efficient and dependable AI-driven ECCC necessitates crucial system improvements, primarily focusing on dynamic resource management, adaptive optimization, and fault tolerance.
$\bullet$ Dynamic Resource Management Dynamic resource management optimizes performance and resource utilization in the fluctuating demands of edge-cloud environments through efficient, real-time resource allocation and reallocation.  Key aspects include algorithms that dynamically adjust resource allocation based on real-time workload characteristics and available resources [118], such as the energy-aware and trust-collaboration cross-domain resource allocation (ETCRA) algorithm, which minimizes system functions while adhering to latency and trust constraints [94].  Collaborative resource management coordinates resource allocation across multiple edge nodes and the cloud for enhanced system efficiency [12].  Additionally, integrating RL allows systems to learn optimal resource allocation strategies based on environmental feedback [18], while dynamic algorithms with initial allocation and real-time updating based on trust assessment improve performance [35]. Multi-timescale algorithms address the varying optimization periods of different sub-problems like service placement and task scheduling within edge-cloud cooperation [118].
$\bullet$ Adaptive Optimization  Adaptive optimization dynamically adjusts system parameters and optimization strategies in response to real-time data and evolving conditions to maximize efficiency. For instance, the dynamic resource allocation mechanism in [18] ensures efficient scheduling of edge resources for global optimization.  Adaptive training with small batches, as proposed in Shoggoth [286], allows edge model adaptation under limited computing power.  Similarly, incremental learning approaches periodically retrain edge AI models using real-time data [191], optimizing prediction accuracy and reducing cloud dependency.  Dynamic model selection based on resources and application requirements [12] and decentralized adaptive scheduling using A3C learning and R2N2 [341] are other crucial aspects.  Additionally, the two-stage iterated greedy optimization (TIGO) algorithm addresses changing user requirements for dynamic microservice placement [110], while adaptive switching between fog and cloud instances based on response time enhances performance [342].
$\bullet$ Fault Tolerance  Fault tolerance ensures reliable operation by implementing fault detection, isolation, and recovery strategies.  DeepFT [343], for example, uses a deep surrogate model for proactive fault prediction and optimized task scheduling.  Ensuring service resilience for critical nodes is addressed by confidence-aware resilience models like CAROL [344], which leverages generative neural networks for QoS prediction and system recovery.  Programmable asset orchestration at the edge contributes to operational resilience, with systems like RECS [292] dynamically managing services and resources during incidents.  Finally, distributed data auditing schemes, often incorporating techniques like FSS [187], ensure data integrity for reliable AI applications.
IX Conclusion
In this survey, we comprehensively examine AI-driven edge-cloud collaborative computing, focusing on distributed intelligence and model optimization. We first outline foundational architectures, then categorize key techniques into model optimization, resource management, and privacy-security enhancements. Additionally, we highlight trends in AI-driven resource allocation and model adaptation. Despite notable progress, challenges in system heterogeneity, security, and application diversity persist, urging further research into advanced AI techniques and emerging technologies like 6G, LLM integration, AI agents, embodied AI, edge quantum computing, etc.