### v4

# Edge-Cloud Collaborative Computing on Distributed Intelligence and Model Optimization: A Survey

Abstract
Edge-cloud collaborative computing (ECCC) has emerged as a pivotal paradigm for addressing the computational demands of modern intelligent applications, integrating cloud resources with edge devices to enable efficient, low-latency processing across distributed communication networks. Recent advancements in AI, particularly deep learning and large language models (LLMs), have dramatically enhanced the capabilities of these networked systems, yet introduce significant challenges in model deployment, network resource management, and cross-layer optimization. In this survey, we comprehensively examine the intersection of distributed intelligence and model optimization within edge-cloud environments, providing a structured tutorial on fundamental architectures, communication protocols, and network-aware computing frameworks. Additionally, we systematically analyze model optimization approaches, including compression, adaptation, and neural architecture search, alongside AI-driven resource management strategies that balance performance, energy efficiency, and communication overhead across heterogeneous networks. We further explore critical aspects of privacy protection and security enhancement within ECCC systems and examine practical deployments through diverse networked applications, spanning autonomous driving, healthcare, and industrial automation. Performance analysis and benchmarking techniques are also thoroughly explored to establish evaluation standards for these complex distributed systems. Furthermore, the review identifies critical research directions including LLMs deployment, 6G integration, neuromorphic computing, and quantum computing, offering a roadmap for addressing persistent challenges in heterogeneity management, real-time processing, and scalability. By bridging theoretical advancements in communications with practical deployments, this survey offers researchers and practitioners a holistic perspective on leveraging AI to optimize distributed computing environments over next-generation communication networks, fostering innovation in intelligent networked systems.
Index Terms:
Edge-Cloud Collaborative Computing, Distributed Intelligence, Model Optimization, AI, Survey
I Introduction
I-A Background
The increasing prevalence of Internet of things (IoT) devices and the demand for real-time applications have propelled a shift from centralized cloud computing to a distributed edge-cloud paradigm [1, 2, 3]. While cloud computing initially offers scalability and cost-effectiveness, limitations such as latency and bandwidth constraints spur the emergence of edge computing, bringing computation closer to data sources [4, 5, 6]. Consequently, edge-cloud collaborative computing (ECCC) arises, integrating the strengths of both paradigms to facilitate efficient resource utilization and enable complex AI tasks across the edge-cloud continuum [7, 8]. AI integration within ECCC systems addresses the inherent complexity of distributed environments, enabling optimized resource allocation and strategic model deployment for enhanced system performance [9]. In addition, AI empowers intelligent decision-making at the edge, enabling real-time responses. However, realizing distributed intelligence in ECCC systems introduces challenges, including efficient model partitioning and resource-aware model optimization, along with ensuring seamless edge-cloud collaboration and addressing security and privacy concerns [10, 11].
Technical foundations underlying ECCC span multiple critical domains that require comprehensive understanding. Model optimization represents a fundamental challenge in deploying sophisticated AI models across resource-constrained edge devices while maintaining computational efficiency and performance [12]. Model compression techniques including pruning, quantization, and knowledge distillation, alongside adaptive learning approaches such as transfer learning, federated learning, and continual learning enable models to evolve with changing data distributions [6]. Resource management in distributed edge-cloud environments involves complex orchestration of computational, communication, and storage resources across heterogeneous infrastructure, requiring intelligent task offloading strategies, dynamic resource allocation algorithms, and energy-efficient optimization techniques to balance performance, cost, and sustainability [13]. Security and privacy protection mechanisms are paramount in ECCC systems due to the distributed nature of data processing and the involvement of multiple stakeholders, necessitating advanced cryptographic protocols, privacy-preserving machine learning techniques, and secure multi-party computation approaches to protect sensitive information while enabling collaborative intelligence [14].
ECCC has emerged as a crucial paradigm to support the growing demands of real-time, intelligent applications, particularly with the proliferation of artificial intelligence of things (AIoT) devices [15]. However, inherent limitations in edge devices, such as constrained computational power and limited battery life, hinder their ability to independently handle computationally intensive AI workloads [16]. Additionally, the dynamic nature of workloads in these environments requires sophisticated resource allocation and task offloading strategies to maintain performance and prevent bottlenecks [17]. Applications like autonomous driving [18] and industrial automation [19] further necessitate low latency and high reliability, demanding intelligent resource management and optimization [20]. Consequently, AI offers a powerful set of tools, including deep reinforcement learning (DRL) for dynamic resource allocation [21] and multi-objective optimization for task scheduling [22], to address these challenges and improve the efficiency and performance of ECCC systems. For instance, federated learning enables collaborative model training across distributed devices while preserving privacy [16], ultimately contributing to enhanced quality of service (QoS) by optimizing resource utilization and minimizing latency.
Communication network evolution plays a pivotal role in enabling effective edge-cloud collaboration by providing the foundational infrastructure that connects distributed computational resources. Recent advancements in network technologies, particularly software-defined networking (SDN) [23] and network function virtualization (NFV) [24], have significantly enhanced the flexibility and programmability of communication infrastructures supporting ECCC. Advanced technologies facilitate dynamic network resource allocation and traffic engineering, which are essential for optimizing data flows between edge devices and cloud servers [25]. Furthermore, emerging wireless communication paradigms, including millimeter-wave communications, massive MIMO, and non-orthogonal multiple access (NOMA) [26], are revolutionizing the capacity, reliability, and efficiency of edge-cloud connectivity. The integration of network slicing techniques [27] enables the creation of virtualized, isolated network segments with tailored quality-of-service guarantees for diverse edge-cloud applications. Moreover, cross-layer optimization approaches that jointly consider communication, computation, and caching resources [28] are becoming increasingly important for minimizing end-to-end latency and maximizing system throughput. As communication networks evolve toward 6G, the synergistic optimization of networking and computing resources will be instrumental in addressing the growing demands of distributed intelligence across the edge-cloud continuum [29].
The selection and optimization of communication protocols significantly impact the performance, security, and reliability of edge-cloud collaborative systems [30]. Protocol optimization encompasses multiple layers with distinct considerations: (1) Application-layer protocols such as HTTP/HTTPS remain fundamental for web-based edge services, while MQTT and CoAP provide lightweight alternatives specifically designed for resource-constrained IoT devices with intermittent connectivity [31]. MQTT’s publish-subscribe model offers excellent scalability for sensor data aggregation scenarios, whereas CoAP’s REST-based approach provides seamless integration with existing web infrastructures, though MQTT introduces broker dependency that can create single points of failure, while CoAP’s UDP-based nature may require additional mechanisms for reliable delivery in lossy networks [32]; (2) Network-layer protocols including IPv6 and emerging protocols like named data networking (NDN) address the scalability challenges of massive IoT deployments. IPv6 provides abundant address space essential for large-scale edge device connectivity, while NDN’s content-centric approach enables efficient data distribution and caching across edge networks, nevertheless IPv6 adoption remains inconsistent across heterogeneous edge environments, and NDN requires significant infrastructure changes that may limit near-term deployment [33]; and (3) Transport-layer innovations such as QUIC and specialized protocols like BBR congestion control offer improved performance over traditional TCP in high-latency, variable-bandwidth edge-cloud connections [34]. QUIC’s built-in encryption and reduced connection establishment overhead particularly benefit mobile edge scenarios, while BBR’s model-based congestion control adapts better to the diverse network conditions encountered in edge deployments. The choice of appropriate protocols depends on specific application requirements, device capabilities, and network characteristics, with optimal edge-cloud systems often employing protocol stacks that dynamically adapt to changing conditions and requirements [35].
I-B Related Works
Over the past few years, several surveys have explored various aspects of edge computing and cloud collaboration, each with distinct scopes and technical emphases. Early works like Abbas et al. [36] and Ferrer et al. [37] establish a foundational understanding of edge computing and decentralized cloud architectures respectively, with their primary focus on architectural foundations rather than AI-driven model optimization. Similarly, Liu et al. [38] provide valuable insights into edge computing systems and tools, while Wang et al. [39] contribute significantly to offloading algorithms, with their emphasis primarily on different aspects of edge-cloud integration rather than comprehensive AI technique integration. More recent surveys have begun exploring particular dimensions of this space. Cao et al. [40] examine edge-cloud assisted cyber-physical systems with strong coverage of resource management, with their primary emphasis on system-level considerations rather than model optimization techniques. Domain-specific surveys like Arthurs et al. [41] focus on intelligent transportation systems with specialized depth in this domain. Despite promising advancements, emerging paradigms such as large language models (LLMs) at the network edge remain an active area of exploration, with Qu et al. [42] providing a dedicated investigation into mobile edge intelligence for LLMs, focusing primarily on resource-efficient techniques within their specialized scope. Additionally, specialized applications of edge intelligence for digital twin-enabled metaverse environments, as examined by Van Huynh et al. [3], offer valuable insights into ultra-reliable and low-latency communications with their emphasis on domain-specific considerations.
The most relevant contemporary works include Yang et al. [4], who extensively analyze reinforcement learning methodologies for mobile edge computing through an advanced exploration of optimization techniques, with their work emphasizing network dynamics as their primary contribution. Similarly, the comprehensive works by Yao et al. [11], Gu et al. [43], and Wang et al. [2] each make valuable contributions with distinct emphases. Specifically, [11] provides excellent coverage of edge-cloud collaboration and AI techniques with particular strength in these core areas. [43] delivers strong analysis of AI-enhanced cloud-edge-terminal networks with good application coverage, emphasizing network architecture aspects. [2] offers comprehensive model optimization coverage across the end-edge-cloud spectrum with particular depth in optimization techniques.
Two surveys are particularly relevant to our work and warrant detailed comparison. Yang et al. [4] provides a comprehensive survey of reinforcement learning approaches for mobile edge computing with excellent coverage of resource management and autonomous driving applications. Their work has a specific focus that differs from our survey in several key aspects: (1) Specialized computing paradigm emphasis, while they focus extensively on edge computing and resource management, our work provides comprehensive coverage of cloud integration and edge-cloud collaboration frameworks; (2) Domain-specific technical focus, their emphasis on reinforcement learning, while thorough, represents a specialized approach compared to our comprehensive treatment of model optimization techniques such as compression, adaptation, and neural architecture search that are central to modern ECCC systems; (3) Different scope for privacy and security, their work emphasizes other technical aspects, while our survey provides systematic attention to privacy-preserving mechanisms and security enhancement for real-world ECCC deployments; (4) Specialized application coverage, while strong in autonomous driving, our work provides comprehensive coverage of manufacturing and healthcare applications that increasingly rely on edge-cloud collaboration. Qu et al. [42] presents a focused investigation into mobile edge intelligence for large language models with strong coverage of LLMs, privacy considerations, and healthcare applications. Their work complements our survey with different emphases: (1) Specialized technical scope, while excellent in LLM-specific optimizations, our survey provides systematic coverage of general model optimization techniques applicable across diverse AI models and architectures; (2) Mobile edge intelligence focus, their focus on mobile edge intelligence represents a specialized area, while our work emphasizes comprehensive edge-cloud collaboration frameworks and distributed intelligence paradigms; (3) Specialized application domain, while strong in healthcare applications, our survey provides detailed analysis of manufacturing and industrial IoT scenarios that represent major application domains for ECCC; (4) Domain-specific security analysis, their privacy coverage addresses specific requirements, while our work provides comprehensive treatment of security mechanisms and adversarial considerations essential for robust ECCC systems.
In contrast to these works, our survey uniquely provides: (1) Comprehensive computing continuum coverage spanning edge, cloud, edge-cloud computing, and collaborative frameworks; (2) Holistic technical integration systematically covering model optimization (compression, adaptation, NAS), resource management, privacy, security, and emerging LLM paradigms; (3) Diverse application analysis with detailed coverage across manufacturing, autonomous driving, and healthcare domains; (4) Unified framework approach that synthesizes distributed intelligence principles with practical implementation considerations. As summarized in Table I, these surveys collectively address different aspects of ECCC, with each making valuable contributions to specific areas of the field. Our survey builds upon these foundational works by providing integrated coverage of model optimization, privacy, security, and diverse applications within a unified framework. Most importantly, existing works have established strong foundations in specific technical aspects (e.g., model optimization or resource management) or particular application domains (e.g., transportation or healthcare), while our work provides a holistic synthesis that spans the entire computing continuum. Consequently, our survey contributes by providing comprehensive analysis that encompasses advanced technical focuses including LLMs, and practical applications across manufacturing, autonomous driving, and healthcare, thereby complementing and extending the existing literature.
TABLE I: Comparative analysis of existing surveys related to AI-driven edge-cloud collaborative computing. Our work uniquely provides comprehensive coverage across the distributed computing continuum, technical focuses, and applications.
Ref. Year Topic Computing Continuum Technical Focuses Applications
Edge Cloud ECCo ECC Model Opt. Task Off. Res. Mgmt. Privacy Security LLMs Manuf. Auto. Drive Health
[36] 2018 Mobile Edge Comp., Network Architecture
[37] 2019 Decentralized Cloud, Ad Hoc Computing
[38] 2019 Edge Computing, Systems, Tools
[39] 2019 Edge Cloud, Offloading Algorithms
[40] 2021 Edge Computing, ECC, CPS
[41] 2022 EC Computing, ITS, Connected Vehicles
[3] 2022 Edge Intel., URLLC, Digital Twin
[44] 2023 Offloading, Federated Cloud-Edge-Fog
[45] 2023 Edge Computing, Sensor-Cloud
[11] 2023 EC Polarization, Collab. AI, Edge Comp.
[46] 2024 AIGC, Mobile Networks, EC Generative AI
[43] 2024 AI, CET Network, Collaborative Computing
[2] 2024 DL, Distributed Computing, Model Opt.
[4] 2025 RL, Mobile Edge Comp., Res. Mgmt.
[42] 2025 Mobile Edge Intel., LLMs, Res. Mgmt.
Ours 2025 ECCC, AI, Distributed Intel., Model Opt.
[…]
I-C Scope and Contributions
Figure 1: Overall structure of this survey. The framework is organized into five layers: (1) Foundations & architecture covering introduction and edge-cloud fundamentals, (2) technical implementation focusing on model optimization and resource management techniques, (3) protection framework addressing privacy and security enhancement, (4) practical deployment discussing applications and performance analysis, and (5) research vision outlining challenges and future directions.
As illustrated in Fig. 1, we provide a systematic and multi-layered examination of ECCC, addressing the critical intersection of distributed intelligence and model optimization. Edge-cloud systems face unprecedented challenges due to heterogeneous infrastructures, resource limitations, and growing computational demands from AI applications, making a comprehensive review of this rapidly evolving field both timely and necessary.
Our survey establishes a comprehensive taxonomy that organizes ECCC into five interconnected layers, each addressing specific aspects of distributed intelligence deployment. Foundations layer encompasses architectural paradigms, distributed intelligence frameworks, and fundamental resource constraints that govern system design. Technical implementation layer focuses on core enabling technologies including model optimization techniques (compression, adaptation, neural architecture search) and AI-driven resource management strategies (task offloading, resource allocation, energy optimization). Protection framework layer addresses critical security and privacy concerns through advanced cryptographic protocols, privacy-preserving learning, and secure computation techniques. Practical deployment layer examines real-world applications across diverse domains and comprehensive performance analysis methodologies. Finally, the Research vision layer identifies current challenges and future research directions that will shape the evolution of ECCC systems. Our taxonomic framework provides researchers and practitioners with a structured understanding of how different components interact within the broader ECCC ecosystem, enabling systematic development and optimization of distributed intelligent systems.
By systematically analyzing ECCC architectures, resource management strategies, and performance metrics, the survey bridges existing gaps in understanding how AI can enhance decentralized computing environments. Moreover, it investigates critical enabling technologies such as model compression, adaptation, resource-aware optimization, and security mechanisms, which are pivotal for overcoming current limitations. The scope spans from architectural foundations and optimization techniques to practical applications and future research directions, offering researchers and practitioners a holistic perspective of this transformative paradigm.
The three primary technical categories of ECCC each address distinct challenges and exhibit specific applicability constraints. Model optimization techniques (Sec. III) primarily target the fundamental challenge of deploying sophisticated AI models across resource-constrained edge devices while maintaining computational efficiency. Model optimization approaches are particularly suitable for scenarios involving deep learning model deployment, privacy-sensitive distributed training, and real-time inference applications. However, their effectiveness is constrained by the inherent trade-offs between model accuracy and compression ratios, the heterogeneity of edge device capabilities, and the complexity of managing distributed learning processes. Model compression methods excel in reducing computational overhead but may compromise accuracy, while federated learning approaches provide privacy preservation at the cost of increased communication complexity and convergence challenges with non-IID data distributions. AI-driven resource management strategies (Sec. IV) focus on optimizing resource utilization and task orchestration across the heterogeneous edge-cloud continuum. Resource management approaches are most effective in scenarios with dynamic workloads, multi-tenant environments, and applications requiring intelligent load balancing across diverse computational resources. The primary limitations include the computational complexity of multi-objective optimization problems, the difficulty in predicting system behavior under varying conditions, and scalability challenges in large-scale deployments. While task offloading mechanisms can significantly reduce latency and energy consumption, they require sophisticated algorithms to handle dependency relationships and may introduce communication overhead that negates performance benefits. Privacy and security enhancement frameworks (Sec. V) address the critical requirements for protecting sensitive data and ensuring system trustworthiness in distributed environments. Security techniques are essential for healthcare, financial, and government applications where regulatory compliance and data protection are paramount. Nevertheless, privacy-preserving mechanisms often introduce computational overhead and complexity in key management, while security measures must continuously evolve to address emerging threats in heterogeneous edge environments.
  • First Comprehensive Survey. To the best of our knowledge, this is the first comprehensive survey that specifically targets the intersection of distributed intelligence and model optimization within broader field of AI-driven edge-cloud collaborative computing, reviewing recent advances across architectures, techniques, and security.
  • Unified Architectural Framework. Our survey establishes a cohesive framework for understanding edge-cloud collaborative computing architectures, resource constraints, and distributed intelligence models. Through systematic categorization, researchers gain a structured understanding of how computational resources can be effectively distributed across the edge-cloud continuum.
  • Multi-domain Application Analysis. Through detailed applications spanning autonomous driving, smart healthcare, industrial automation, and smart cities, our work demonstrates the practical impact of AI-driven edge-cloud computing across diverse domains and identifies domain-specific requirements and challenges.
  • Future Research Roadmap. By identifying technical challenges and promising research directions, including advanced AI techniques, system improvements, and emerging applications, we provide a comprehensive roadmap for researchers and practitioners to advance the field of AI-driven edge-cloud collaborative computing.
I-D Research Methodology and Organization
Figure 2: Publication trends and citation analysis of research in distributed computing paradigms from 2021 to 2025.
Recent analysis reveals significant research momentum in ECCC, as illustrated in Fig. 2, where publication trends and citation patterns demonstrate the evolving landscape across four interrelated domains (The statistics data is collected as of 2025/03/05). While traditional edge computing maintains consistent publication volumes with steadily increasing citation impact, cloud computing exhibits characteristics of a mature research area with decreasing publication numbers yet sustained high citation rates. Most notably, edge-cloud computing shows clear upward trajectories in both publications and citations, signaling growing research interest, whereas the specialized subfield of edge-cloud collaboration displays exponential citation growth peaking in 2024, indicating its emerging importance. Our systematic review methodology employed comprehensive literature searches across prominent scholarly databases including IEEE Xplore, ACM Digital Library, ScienceDirect, and Web of Science, utilizing carefully selected keywords such as ”edge-cloud computing”, ”distributed intelligence”, ”model optimization”, ”AI”, and ”machine learning”. Following retrieval, selected papers pertinent to ECCC were carefully reviewed and categorized based on their primary contributions aligned with the scope of this work.
The remainder of this survey provides a comprehensive overview of AI-driven ECCC. We begin with edge-cloud computing fundamentals including architecture and distributed intelligence in Sec. II. Sec. III examines model optimization techniques including compression and neural architecture search. Sec. IV explores AI-driven resource management covering task offloading and energy efficiency, while Sec. V addresses privacy and security enhancement. Sec. VI presents practical applications spanning autonomous driving, healthcare, and smart cities. Sec. VII delivers performance analysis through benchmarking and evaluation, followed by challenges and future directions in Sec. VIII, with conclusions in Sec. IX.
II Architecture and Fundamentals
Modern intelligent applications demand architectures capable of processing vast data volumes with minimal latency, driving innovation toward distributed computing paradigms that bridge resource-rich cloud infrastructure with proximity-based edge processing. Fig. 3 provides a conceptual illustration of AI-driven edge-cloud collaborative systems, depicting the core infrastructure interplay between cloud servers and edge nodes, alongside essential AI-enabled techniques (e.g., model adaptation, resource-aware optimization) that facilitate distributed intelligence across various application scenarios. Building upon this visual overview, the subsequent subsections delve into the foundational aspects of this architecture. We begin by outlining the core computing paradigm in Sec. II-A, subsequently examining the distributed intelligence framework enabling collaboration in Sec. II-B. Following this, an analysis of crucial resource types and their inherent constraints is presented in Sec. II-C.
II-A Edge-Cloud Computing Paradigm
Distributed computing architectures optimize performance by positioning computational and storage resources closer to data sources, implementing a three-tier hierarchy encompassing cloud infrastructure, edge nodes, and end devices [47]. The cloud tier provides extensive computational and storage resources for large-scale data processing and model training, while the edge tier, comprising edge servers and gateways, offers localized processing to reduce latency [48]. End devices, such as smartphones and IoT sensors, generate and consume data at the network edge. Effective resource management and orchestration are crucial for efficient resource allocation across these tiers. Distributing intelligence across the network, a key characteristic of this paradigm, can be represented by a function $D(R_{i},L_{j})$, quantifying the proportion of resource $R_{i}$ allocated to layer $L_{j}$ [49]. Edge-cloud systems aim to optimize resource allocation, considering factors like application requirements and network conditions, to maximize a utility function $U(P,C,L)$ encompassing performance $P$, cost $C$, and latency $L$, subject to resource constraints [50]. Consequently, processing data closer to its source reduces latency and bandwidth consumption. Additionally, distributing workloads enhances scalability and fault tolerance, improving the system’s ability to handle increasing data volumes and withstand node failures [51]. Edge federation models further enable collaborative resource sharing, optimizing service delivery across a federated network [52], especially for mission-critical applications [53].
II-B Distributed Intelligence Framework
Distributed intelligence frameworks optimize performance by strategically coordinating computational resources across heterogeneous infrastructure, leveraging both proximity-based edge processing and scalable cloud capabilities to address unique constraints and opportunities inherent in collaborative environments. The distributed intelligence framework in ECCC adheres to three key principles specifically adapted for edge-cloud deployment: task parallelism, data parallelism, and model parallelism.
Task parallelism in edge-cloud contexts executes different parts of complex AI tasks concurrently on separate devices, considering edge device limitations and cloud processing capabilities [54], represented as $T=\{t_{1}^{edge},t_{2}^{cloud},...,t_{n}^{hybrid}\}$, where tasks are strategically assigned based on latency requirements, computational complexity, and data locality. Data parallelism trains identical models on different data subsets distributed across multiple edge devices while maintaining privacy [11], expressed as $D=\{d_{1}^{local},d_{2}^{local},...,d_{m}^{local}\}$, where data remains decentralized to address privacy concerns and network bandwidth limitations. Model parallelism deploys different parts of large models on separate devices within the edge-cloud continuum [55], formulated as $M=\{m_{1}^{edge},m_{2}^{fog},...,m_{p}^{cloud}\}$, where model partitioning considers device capabilities, network latency, and energy constraints.
Key design challenges include managing heterogeneous device capabilities across the edge-cloud spectrum, where edge devices have limited computational power but offer low latency, while cloud resources provide extensive computation but introduce communication delays. Network reliability and bandwidth variability require adaptive model partitioning strategies that can dynamically adjust based on real-time network conditions. Energy efficiency becomes paramount for battery-powered edge devices, necessitating intelligent workload distribution that balances computational load with energy consumption. Privacy and data locality requirements demand sophisticated federated learning approaches that enable collaborative intelligence without centralizing sensitive data.
Edge intelligence focuses on real-time processing with sub-10ms latency requirements [56], making it suitable for time-critical applications like autonomous driving and industrial control. Cloud intelligence handles complex computations and large-scale training using virtually unlimited resources [57], enabling sophisticated model development and global knowledge aggregation. Model partitioning strategies must balance workloads by considering model complexity, data size, network topology, and resource constraints specific to edge-cloud deployments [58]. Resource allocation algorithms dynamically assign computational, storage, and communication resources based on real-time demands, device capabilities, and edge-cloud topology [59]. Performance optimization techniques like model compression enhance efficiency across the distributed architecture by reducing communication overhead and enabling deployment on resource-constrained devices [60]. Additionally, fault tolerance mechanisms ensure system reliability despite potential edge device failures, network partitions, or cloud service interruptions, requiring redundancy strategies and graceful degradation protocols.
II-C Resource Types and Constraints
Figure 3: Concepts of architecture. The diagram provides a holistic view of the technical scope covered in this survey, where the architecture integrates centralized cloud resources, decentralized edge nodes, and diverse mobile/edge devices. AI-enabled technologies include model compression, adaptation, and resource-aware optimization, whose components jointly support distributed intelligence and efficient model execution across multiple application scenarios.
Fundamental system resources, including computation, communication bandwidth, and storage capacity, present inherent constraints that significantly impact distributed architecture design and performance optimization across heterogeneous infrastructure [48]. Specifically, computation resources, denoted as $C_{k}$ for node $k$, represent processing power; communication bandwidth, $B_{ij}$, limits data transfer rates between nodes $i$ and $j$, directly affecting latency; and storage capacity, $S_{k}$, dictates the data volume node $k$ can store, influencing data-intensive applications [61]. Managing these constrained and heterogeneous resources across diverse devices, under dynamic workloads and with limited edge resources, especially energy for battery-powered devices [49], presents significant challenges.
Dynamic resource allocation addresses these challenges by adapting to workload fluctuations, aiming to find an optimal mapping of tasks $\mathcal{T}$ to nodes $\mathcal{N}$ that minimizes a cost function (e.g., latency) subject to resource constraints [62]. Mathematically, this involves minimizing $\sum_{t\in\mathcal{T}}Cost(M(t),t)$ subject to $\forall n\in\mathcal{N}:\sum_{t:M(t)=n}R(t)\leq C_{n}$, where $Cost(M(t),t)$ is the cost of executing task $t$ on node $M(t)$, and $R(t)$ is the resource requirement of task $t$. In addition, energy-efficient offloading minimizes energy consumption, $E_{ij}(t)=E_{comm}(t,B_{ij})+E_{comp}(t,C_{j})$, by strategically offloading tasks, where $E_{comm}$ and $E_{comp}$ represent communication and computation energy, respectively [39]. Efficient edge storage management, leveraging techniques like data caching and prefetching [63], is also crucial, especially for data-intensive applications, alongside integrating QoS considerations [64].
III Model Optimization for Edge-Cloud Systems
Collaborative computing environments present distinctive optimization challenges that fundamentally differ from traditional centralized or purely cloud-based approaches. Distributed environments impose specific constraints that significantly impact optimization strategies: (1) Extreme resource heterogeneity, edge devices range from smartphones with limited processing power to industrial gateways with substantial computational capabilities, necessitating adaptive optimization techniques that can dynamically scale across this spectrum; (2) Intermittent connectivity, network links between edge and cloud are often unreliable, requiring optimization methods that can function during disconnection periods and efficiently synchronize when connectivity is restored; (3) Real-time processing requirements, many edge applications demand ultra-low latency responses (<< 10ms), making traditional batch optimization approaches unsuitable; (4) Privacy-preserving constraints, sensitive data often cannot leave edge devices, requiring optimization techniques that work with limited data visibility; (5) Energy limitations, battery-powered edge devices necessitate energy-aware optimization that balances computational efficiency with power consumption. Consequently, model optimization techniques must be specifically adapted to address these edge-cloud specific constraints, moving beyond general-purpose optimization to develop specialized approaches that leverage the unique characteristics of the edge-cloud continuum while mitigating its inherent limitations.
III-A Model Compression
Resource-constrained environments demand sophisticated model compression techniques to enable deployment of complex AI models while maintaining acceptable performance levels despite limited computational power, memory, and energy resources. Advanced compression approaches [2, 65, 66] have emerged as critical solutions to reduce model size and computational demands while preserving performance. Table II provides a comparison of recent approaches across three primary compression categories: pruning, quantization, and knowledge distillation.
Model compression techniques in ECCC systems can be systematically categorized based on their compression mechanisms and deployment characteristics: (1) Structural modification approaches (pruning) that reduce model complexity by eliminating redundant parameters or architectural components, particularly effective for over-parameterized networks commonly deployed in edge scenarios; (2) Numerical precision optimization (quantization) that maintains model structure while reducing parameter precision requirements, enabling efficient deployment on specialized edge hardware with quantization support; (3) Knowledge transfer and distillation that creates compact models through supervised learning from larger teacher models, providing architectural flexibility and optimal accuracy-efficiency trade-offs for edge deployment. The classification framework reflects both the underlying technical mechanisms and practical deployment considerations essential for edge-cloud environments.
Each compression category exhibits distinct advantages and limitations for edge-cloud deployment scenarios. Pruning techniques excel in reducing model size and computational complexity by eliminating redundant parameters, making them particularly effective for convolutional neural networks where structured pruning can remove entire channels or layers. However, pruning may require fine-tuning to recover accuracy losses and can be sensitive to the pruning ratio selection [67]. Quantization methods provide significant memory reduction and accelerated inference by reducing parameter precision, with minimal accuracy degradation when carefully implemented. They are especially suitable for deployment on specialized hardware with quantization support, though they may face challenges with extremely low-bit quantization or complex model architectures [68]. Knowledge distillation offers the unique advantage of training compact models from scratch while leveraging teacher model knowledge, enabling flexible student model architectures. However, it requires access to both teacher models and training data, making it computationally intensive during the training phase [69].
The choice of compression technique should consider several factors: (1) Resource constraints, pruning is preferred for memory-limited devices, quantization for computational acceleration, and knowledge distillation for balanced performance-efficiency trade-offs; (2) Model complexity, structured pruning works well for over-parameterized models, quantization suits models with redundant precision, and knowledge distillation excels for complex teacher models; (3) Deployment requirements, pruning enables immediate deployment, quantization requires hardware compatibility, and knowledge distillation demands training infrastructure; (4) Performance sensitivity, knowledge distillation typically achieves the best accuracy preservation, followed by quantization and pruning respectively.
Notably, pruning techniques demonstrate significant improvements in inference latency and energy efficiency by strategically eliminating less important parameters [70], as further detailed in Sec. III-A1. Meanwhile, quantization methods, explored in Sec. III-A2, reduce parameter precision requirements, with implementations achieving remarkable energy efficiency in healthcare applications. Knowledge distillation approaches, discussed in Sec. III-A3, enable smaller models to effectively mimic larger counterparts, with recent innovations specifically optimized for efficient LLM inference across edge-cloud boundaries. Across these diverse techniques, a common pattern emerges wherein most successful implementations leverage both edge and cloud resources collaboratively, highlighting the synergistic nature of modern compression strategies in distributed computing environments.
TABLE II: Summary of model compression techniques.
Method Year Venue Type Domain Performance Metrics Edge Cloud
TSCF [67] 2024 IEEE IoTJ P Edge-Cloud Systems Inference Latency
Hybrid SD [71] 2024 ArXiv P Image Synthesis Parameter Efficiency, Image Quality
PSFed [72] 2024 IEEE JSAC P Satellite Edge Cloud Latency, Energy, Privacy
3CE2C [73] 2024 ASC Q Healthcare/ECG Accuracy, Energy Efficiency
CDC [74] 2020 IJCAI P Edge-Cloud Systems Bandwidth Efficiency
QuAsyncFL [68] 2024 IEEE IoTJ Q AIoT Communication Efficiency
Zhang et al. [75] 2025 IEEE JSAC Q Aerial Edge Computing Mean Average Precision (mAP)
DTS-KD [65] 2024 IEEE IoTJ K Industrial IoT F1 Score, Accuracy, Precision, Recall
Wang et al. [69] 2024 CVPR K Multimodal LLMs Generalization Performance
Wu et al. [76] 2024 IEEE TII P Aviation Manufacturing Edge Detection Accuracy
FedAgg [77] 2024 INFOCOM K End-Edge-Cloud Computing Model Accuracy, Convergence Rate
Yang et al. [6] 2025 ADHOC K IoT Traffic Identification Classification Accuracy, Resource Efficiency
GKT [78] 2024 ACL K LLM Inference Inference Speed, Accuracy, Cost
OTO [79] 2024 IEEE TPDS K Reinforcement Learning Convergence Speed, Reward
Zhang et al. [80] 2024 ESWA K Manufacturing Processes Fault Detection Accuracy
Note for ’Type’: P = Pruning, Q = Quantization, K = Knowledge Distillation.
III-A1 Pruning
Parameter elimination strategies reduce model complexity by removing redundant components, enabling efficient AI deployment on resource-constrained devices through targeted removal of less critical weights or architectural elements [71]. We explore various strategies targeting different architectural aspects, such as pruning entire neurons or filters based on performance contribution, as in the two-phase split computing framework (TSCF) [67], or focusing on individual weights. Applied during or after training, dynamic and static pruning balance accuracy, and computational cost. For instance, Hybrid SD [71] employs structural pruning of the Stable Diffusion Model’s U-Net to maintain image quality on edge devices, while the adaptive pruning-split federated learning (PSFed) method [72] optimizes semantic coder updates in a SemCom-assisted SEC framework, reducing training delay and energy consumption.
III-A2 Quantization
Precision reduction techniques effectively minimize memory footprint and computational cost by employing lower-precision representations of model parameters [74], making quantized models particularly suitable for resource-constrained devices. Instead of 32-bit floating-point representations, quantization employs lower precision formats such as 8-bit integers, consequently diminishing memory requirements and accelerating computations on these lower-precision numbers. For example, [73] explored model quantization for edge implementation in arrhythmia monitoring to reduce energy consumption. Additionally, QuAsyncFL [68] integrates quantization with asynchronous federated learning to enhance communication efficiency in cloud-edge-terminal collaborations within the AIoT domain.
III-A3 Knowledge Distillation
Knowledge transfer approaches enable compact ”student” models to achieve comparable performance with reduced complexity by learning from larger ”teacher” models [78], facilitating efficient deployment in resource-constrained environments [65]. Instead of learning from ground truth labels, the student model learns the teacher model’s output distribution, providing richer information than hard labels and enabling more nuanced representations and better generalization [69]. KD benefits ECCC by placing the teacher model in the resource-rich cloud and the distilled student model on edge devices [6], optimizing overall system performance. Agglomerative federated learning (FedAgg) [77] uses a bridge sample based online distillation protocol (BSBODP) for knowledge transfer between nodes across different tiers, enabling larger model training while maintaining privacy and flexibility. Guidance-based knowledge transfer (GKT) [78] employs a larger LLM as a ”teacher” to generate guidance prompts for a smaller ”student” model, enhancing accuracy and speed without fine-tuning. Growth-adaptive distillation [6] automatically adjusts the student model structure for efficient, low-power edge deployment. In cloud-edge collaborative reinforcement learning, algorithm-independent knowledge distillation [79] accelerates online RL agent convergence by leveraging pre-trained cloud models and offline agent interactions. A federated bi-directional knowledge distillation strategy [80] enhances fault detection in manufacturing by integrating quality knowledge from both cloud and edge, strengthening interaction between sub-processes and hierarchical levels.
III-B Model Adaptation
Figure 4: Illustration of model adaptation, detailing (a) transfer learning, (b) federated learning, and (c) continual learning.
Adaptive learning paradigms enable AI models to effectively respond to new tasks, datasets, or evolving data distributions within dynamic distributed computing environments. To achieve this adaptability, three primary learning paradigms are widely employed: transfer learning (TL), which accelerates learning on target tasks by leveraging knowledge from pre-trained models; federated learning (FL), which facilitates privacy-preserving, distributed training across multiple clients; and continual learning (CL), which enables models to learn sequentially from evolving data streams without catastrophic forgetting. The interplay and distinct roles of these paradigms in fostering adaptability within ECCC systems are conceptually depicted in Fig. 4.
While some taxonomies treat federated learning as a distinct resource management or privacy technique, our classification positions FL within model adaptation because it fundamentally addresses how models adapt to distributed data patterns and heterogeneous environments, which is a core adaptation challenge in ECCC systems. Specifically: (1) Domain adaptation techniques (transfer learning) enable models to adapt across different domains or tasks by leveraging pre-existing knowledge; (2) Distributed adaptation methods (federated learning) enable models to adapt collaboratively across multiple data sources while preserving privacy and handling non-IID distributions; (3) Temporal adaptation approaches (continual learning) enable models to adapt incrementally to evolving data and new tasks over time. The classification emphasizes the adaptation aspect of each technique while acknowledging that FL also serves privacy and resource distribution functions, making it inherently multi-faceted in ECCC systems.
Each learning approach offers unique advantages for different edge-cloud deployment scenarios and data distribution characteristics. Transfer learning excels in scenarios with limited target domain data by leveraging pre-trained models from related domains, achieving faster convergence and reduced computational requirements during training [81]. However, transfer learning performance heavily depends on the similarity between source and target domains, potentially struggling with domain shifts or highly specialized tasks. Federated learning provides exceptional privacy preservation by keeping data distributed across edge devices while enabling collaborative model training, making it ideal for scenarios with sensitive data or regulatory constraints [77]. Nevertheless, FL faces challenges with non-IID data distributions, communication overhead, and convergence difficulties in heterogeneous environments [82]. Continual learning offers superior adaptability to evolving data distributions and new tasks without requiring access to historical data, enabling lifelong learning capabilities crucial for dynamic edge environments [83]. However, CL techniques often struggle with catastrophic forgetting and require sophisticated regularization mechanisms to maintain performance on previous tasks.
The choice of learning paradigm should align with specific data distribution characteristics: (1) Homogeneous data distributions, transfer learning works well when source and target domains share similar characteristics, while federated learning benefits from consistent data patterns across participants; (2) Non-IID data distributions, specialized FL techniques (e.g., FedProx, FedNova) address statistical heterogeneity, while transfer learning may require domain adaptation methods; (3) Temporal data shifts, continual learning excels in adapting to evolving data patterns, while transfer learning may need periodic retraining; (4) Limited local data, transfer learning provides quick adaptation with few samples, while FL can leverage collective knowledge across devices even with sparse local datasets. The optimal selection often involves hybrid approaches that combine multiple paradigms to address specific edge-cloud requirements and constraints. Furthermore, Table III presents a comparative summary of recent research efforts utilizing TL, FL, and CL, highlighting their specific approaches, application domains, and system characteristics within ECCC. Following this, the subsequent sections delve into the transfer learning in Sec. III-B1, federated learning in Sec. III-B2, and continual learning in Sec. III-B3.
TABLE III: Summary of TL, FL, and CL techniques for enhancing AI model adaptability in ECCC systems.
Method Year Venue Type Domain Edge Cloud ECC MC
CoLLaRS [84] 2024 FGCS CL AIoT Applications
Li et al. [81] 2024 IEEE TNSM TL Manufacturing
Pan et al. [85] 2024 IEEE Netw. FL Large Model Services
MADDPG [86] 2024 IEEE IoTJ TL 6G Networks
DTS-KD [65] 2024 IEEE IoTJ TL Industrial IoT
Wang et al. [69] 2024 CVPR CL Multimodal Large Language Models
Wang et al. [87] 2024 ESWA FL Rotating Machinery Fault Diagnosis
FedAgg [77] 2024 INFOCOM FL End-Edge-Cloud Systems
OTO [79] 2024 IEEE TPDS TL Cloud-Edge Collaborative DRL
U-Federator [88] 2024 JNCA FL Cloud-Edge-Fog Computing
OTFAC [89] 2024 IEEE IoTJ FL Cloud-Edge Collaborative IoT
FLAT [90] 2024 IEEE ISCC FL Cloud-Edge-Client Continuum
FedITA [91] 2024 AEI FL Industrial Motor Fault Diagnosis
Khuen et al. [92] 2024 IEEE Access FL Precision Aquaculture
Kim et al. [93] 2024 IEEE IoTJ FL Cloud-Edge-Terminal IoT Networks
QuAsyncFL [68] 2024 IEEE IoTJ FL Cloud-Edge-Terminal AIoT
Loconte et al. [94] 2024 FGCS FL Cloud-to-Edge-IoT Continuum
BAFL-DT [82] 2024 IEEE TSC FL Digital Twin Edge-Cloud Continuum
FDRL [95] 2024 IJPR FL Cloud-Edge Manufacturing Systems
Wei et al. [96] 2024 IEEE TPDS FL Edge Clouds
Yang et al. [97] 2024 IEEE IoTJ FL AIoT Data Prediction
FedCAE [98] 2024 IEEE TIE FL Machine Fault Diagnosis
FlexibleFL [99] 2024 INS FL Cloud-Edge Federated Learning
PSPFL [100] 2024 IEEE TMC FL Mobile Edge-Cloud Networks
CCSFLF [101] 2024 CCPE FL Federated Learning Systems
Zhou et al. [16] 2024 ADHOC FL AIoT Quality of Service
CADCL-DTM [102] 2024 CAC CL Rolling Bearings Monitoring
Li et al. [103] 2024 IEEE TMC CL Edge Computing Service Provisioning
Lian et al. [104] 2024 SDI CL ITS Object Detection
EdgeC3 [105] 2024 IEEE TMC CL Distributed Deep Learning
Note for ’Type’: TL = Transfer Learning, FL = Federated Learning, CL = Continual Learning, MC = Model Compression.
III-B1 Transfer Learning
Pre-trained model adaptation to new tasks or datasets leverages transfer learning principles to reduce training time and data requirements, thereby improving deployment efficiency in resource-constrained environments [81]. For example, in surface defect detection, a model pre-trained on a general image dataset can be fine-tuned with a smaller, specialized dataset to identify specific defects [81]. Similarly, in cloud-device collaborative learning for multimodal large language models (MLLMs), techniques like adapter-based knowledge distillation (AKD) facilitate knowledge transfer from larger cloud models to smaller, compressed device models [69], maintaining performance while minimizing computational burden. In cloud-edge collaborative fault diagnosis, transfer learning addresses data heterogeneity and distribution shifts by fine-tuning source models trained on diverse user data to bridge distribution gaps and enable effective diagnosis on target datasets [87].
III-B2 Model Adaptation in Federated Learning
The explosive growth of FL research has made it an indispensable paradigm in distributed AI systems, driving critical need to explore how model adaptation mechanisms can be effectively implemented within federated frameworks [100]. As one of the most active research domains in distributed machine learning, FL cannot be overlooked when examining model adaptation techniques, particularly given its unique challenges in handling heterogeneous data distributions and diverse edge environments [92]. We focus specifically on model adaptation strategies within FL contexts, exploring how adaptive mechanisms enable models to effectively respond to distributed, non-IID data patterns while maintaining collaborative training capabilities [94]. Recent frameworks demonstrate sophisticated adaptation approaches, such as FedCAE [98] for machine fault diagnosis that adapts models to diverse industrial environments, and QuAsyncFL [68] for AIoT applications that enables adaptive synchronization strategies across heterogeneous devices. Advanced adaptation mechanisms like SFETEC [106] optimize model adaptation for thing-edge-cloud environments by implementing adaptive model splitting strategies that dynamically adjust to varying communication constraints and device capabilities, addressing non-IID data challenges through intelligent partitioning. Security-aware adaptation frameworks such as FlexibleFL [99] integrate adaptive evaluation mechanisms that assess uploaded parameters and participant contributions, enabling models to adapt against evolving poisoning attacks while maintaining collaborative learning effectiveness. Furthermore, adaptive participant selection and topology optimization strategies [96] enhance FL efficiency by dynamically adjusting collaboration patterns, while blockchain integration approaches [82] provide adaptive security mechanisms that evolve with emerging threats, demonstrating how model adaptation principles are fundamental to advancing FL systems in distributed edge-cloud environments.
III-B3 Continual Learning
Incremental adaptation strategies enable models to evolve with dynamic data and emerging tasks while addressing catastrophic forgetting challenges through lifelong learning capabilities [83]. Several strategies, including regularization, replay, and dynamic architectures, mitigate this issue, ensuring robust and adaptable systems. Specifically, continual learning enables digital twins to synchronize with their evolving physical counterparts and sensor data in edge computing, providing accurate predictions [103, 102]. For instance, in intelligent transportation systems (ITS), cloud-edge collaborative continual adaptation frameworks, leveraging visual prompts and knowledge distillation, enhance the generalization of lightweight edge models for object detection in dynamic environments [104]. Similarly, online management frameworks, such as EdgeC3, optimize model aggregation and dynamic offloading for continuous data streams in edge-cloud collaborative continual learning [105]. Cross-edge federated continual learning (Cross-FCL) frameworks, employing parameter decomposition and various cross-edge strategies, address knowledge retention across devices in multiple FL systems, balancing memory and adaptation for new tasks [107].
III-B4 Split Learning
Distributed learning paradigms partition neural network models across multiple computational nodes to enable collaborative training without sharing raw data, addressing computational limitations while preserving privacy [108]. Unlike FL where complete models are trained locally, split learning divides the neural network architecture at specific cut layers, with edge devices typically handling initial feature extraction layers while cloud or edge servers process deeper computational layers [109]. Strategic architectural division effectively addresses the computational limitations of resource-constrained IoT devices while preserving data privacy, as demonstrated in eHealth applications where heterogeneous workload-based resource recommendation models leverage federated split learning (FSL) for smart city deployments [110]. Advanced frameworks like MADRL-based model partitioning systems optimize the collaboration between cloud-edge-device architectures by jointly considering model aggregation control and resource allocation, addressing device heterogeneity through adaptive partitioning strategies rather than uniform approaches [111].
The optimal selection of split points significantly impacts system performance across multiple dimensions including computational load distribution, communication overhead, and privacy preservation. Efficient parallel split learning (EPSL) frameworks demonstrate how intelligent cut layer selection, combined with subchannel allocation and power control optimization, can minimize per-round training latency in resource-constrained wireless edge networks [112]. The SFETEC framework exemplifies optimized split learning for thing-edge-cloud environments, where base models are trained at client devices while core models are processed at edge servers and aggregated at the cloud, significantly reducing communication overhead and training duration compared to traditional federated approaches [106]. Advanced split learning architectures like Split-and-Bridge address class incremental learning challenges by adaptively partitioning neural networks into specialized segments for handling new and old tasks, demonstrating superior performance in knowledge distillation-based continual learning scenarios [113].
Split learning implementations must carefully balance communication efficiency with model performance, particularly in resource-constrained environments. The CSE-FSL framework introduces auxiliary networks for local weight updates while maintaining single server-side models, dramatically reducing gradient transmission requirements and server storage overhead [114]. For wearable IoMT devices, FedSL frameworks enable collaborative healthcare analytics by allowing even low-end devices to participate in model training through strategic model partitioning, while maintaining privacy by avoiding raw data transmission over wireless networks [115]. Comprehensive evaluations comparing FL and split learning in IoT settings reveal that while split learning provides better learning performance under imbalanced data distributions, FL demonstrates superior communication efficiency, highlighting the importance of application-specific optimization strategies [108]. Optimization considerations collectively enable split learning deployment across diverse 6G edge networks, facilitating connected intelligence through distributed model training architectures [109].
III-C Resource-Aware Model Optimization
Resource-aware model optimization represents a critical advancement in ECCC systems, focusing on dynamic adaptation to heterogeneous and constrained computational environments. Advanced optimization strategies encompass three primary approaches: dynamic model selection based on real-time resource availability and application requirements, model splitting and offloading that strategically partition components across edge-cloud infrastructure, and reinforcement learning-based optimization for adaptive resource allocation. Dynamic model selection considers computational constraints, communication bandwidth, and energy limitations while maintaining QoS requirements, with techniquess ranging from online learning approaches to trust-aware selection strategies [116, 62]. Model partitioning leverages heterogeneous system capabilities by offloading computationally intensive components to resource-rich cloud while executing latency-sensitive parts at the edge, with sophisticated algorithms like MADDPG enhancing efficiency through workload-aware adaptation [23, 117]. Reinforcement learning mechanisms optimize resource allocation through environmental feedback, employing techniques like deep Q-learning and APPO to balance computational efficiency with system performance, achieving improved stability and reduced failure rates in IoT monitoring applications [21, 118]. Consequently, integrated optimization strategies establish a versatile framework for diverse domains including IoT networks, vehicular computing, and 5G environments, addressing the fundamental challenges of resource heterogeneity and dynamic operational constraints in distributed computing systems. For comprehensive technical details, evaluation methodologies, and implementation strategies, please see LABEL:sec:supp-tech-comparison of the Supplementary Material.
III-D Neural Architecture Search
Neural architecture search (NAS) automates the design and optimization of neural network architectures to discover optimal configurations for edge-cloud collaborative environments, systematically exploring architectural configurations tailored to distributed computing constraints. Advanced automated methodologies encompass three primary paradigms: reinforcement learning-based NAS where agents learn optimal architecture selection through performance feedback and reward systems that incorporate edge-cloud deployment constraints; evolutionary algorithm-based NAS that employs natural selection principles to evolve architecture populations through iterative selection, mutation, and crossover operations; and gradient-based NAS that formulates architecture search as continuous optimization problems enabling efficient exploration through established gradient descent techniques [119, 120, 22]. Reinforcement learning approaches demonstrate particular effectiveness in resource-constrained environments by directly incorporating computational, memory, and energy constraints into reward functions, enabling discovery of architectures specifically optimized for edge device limitations [25, 121]. Evolutionary algorithms provide exceptional adaptability to dynamic edge-cloud environments through inherent population diversity and exploration capabilities, with applications ranging from scientific workflow optimization to adaptive resource allocation [122, 123]. Gradient-based methods achieve superior search efficiency by representing architecture spaces differentiably, reducing search time significantly while enabling exploration of complex architectural configurations, though challenges remain in handling operation correlations and incorporating device constraints directly into the search process [124, 125]. These complementary approaches collectively enable automated discovery of neural architectures that outperform manually designed alternatives while meeting the stringent performance and resource requirements of distributed edge-cloud deployments. For detailed technical analysis, comparative evaluation, and implementation guidelines, please see LABEL:sec:supp-resource-nas of the Supplementary Material.
IV AI-Driven Resource Management
Effective resource management in ECCC systems requires intelligent orchestration of computational, communication, and storage resources across heterogeneous infrastructure to optimize system performance while meeting diverse application requirements. The section examines three complementary yet distinct aspects of AI-driven resource management. Task offloading mechanisms (Sec. IV-A) focus on intelligent decision-making for task placement across the edge-cloud continuum, considering factors such as latency requirements, energy constraints, and task dependencies to determine optimal execution locations. Resource allocation strategies (Sec. IV-B) address the dynamic distribution and management of computational resources (CPU, memory, storage) across edge nodes and cloud servers, emphasizing adaptive allocation algorithms and collaborative resource sharing approaches. Energy efficiency optimization (Sec. IV-C) concentrates on minimizing power consumption throughout the system lifecycle, including power modeling, energy-aware scheduling, and joint optimization techniques that balance performance with sustainability requirements.
IV-A Task Offloading Mechanisms
Figure 5: Task offloading mechanisms, including latency-aware, energy-efficient, and dependency-aware offloading strategies.
TABLE IV: Summary of task offloading mechanisms, categorized by strategy type and performance optimization targets.
Type Method Metrics Achievements Gain ECC
Latency-Aware
iGATS [8] Response latency, Load balancing Outperforms baseline techniques for latency reduction High
Fan et al. [126] Task processing delay, Stability Superior performance vs. 4 comparison schemes Medium
Zhai et al. [127] Response delay, Carbon emissions 33.42× delay reduction for sensitive workloads Very High
Luckow et al. [128] Performance impact, Task placement 5-65% performance improvement across scenarios High
Energy-Efficient
PDCO [129] Energy consumption, Resource util. Approaches optimal performance in energy metrics High
Hua et al. [117] Energy consumption of mobile devices Significantly reduced energy consumption High
ETCRA [130] Energy, Time, Reliability Outperforms baselines on four key measurements High
Hu et al. [131] Energy efficiency, Service delay Achieves [O(1/V), O(V)] EE-delay tradeoff Medium
Dependency-
Aware
PRAISE [132] System benefits, Resource costs Higher benefits and lower costs vs. baselines Medium
Peng et al. [133] Computing time, Blocking probability Optimal replication with varying parameters Medium
Multi-Objective
Zhang et al. [22] Time, Cost, Load balance, Trust Found >>50% of best solutions on benchmarks High
DRLIS [134] Load balancing, Response time, Cost Up to 55% cost reduction across metrics Very High
DT-MADDPG [135] Success rate, Latency, Energy Improved success rate with reduced resource usage High
Zhang et al. [136] Time, Cost, Energy, Reliability Outperforms existing design approaches Medium
EdgeOptimizer [137] Success rate, Overall overload Improved success rate for time-critical tasks Medium
Notes: Performance Gain: Very High (>>50%), High (20-50%), Medium (5-20%).
Effective resource utilization and performance optimization in ECCC depend critically on sophisticated task offloading mechanisms. As depicted in Fig. 5, a framework outlines three core strategies: latency-aware, energy-efficient, and dependency-aware offloading, detailing their interactions with edge devices, cloud servers, and supporting systems to improve task placement and scheduling. Specifically, these mechanisms address distinct objectives such as real-time performance, energy consumption, and task dependencies, thereby boosting overall system efficiency. To further illustrate their practical impact, Table IV summarizes recent task offloading methods in edge-cloud environments.
Notably, the table reveals significant performance gains, such as up to a 33.42× reduction in delay for sensitive workloads [127], alongside improvements in energy consumption and system reliability, highlighting the versatility of these strategies across diverse application domains.
IV-A1 Latency-Aware Offloading
Minimizing response time through latency-aware offloading represents a crucial strategy for time-sensitive applications in edge-cloud systems, with numerous approaches demonstrating significant performance gains. According to the comparative analysis in Table IV, methods like iGATS [8] and DRLIS [134] have achieved substantial reductions in processing delays through intelligent placement decisions. For instance, Lyapunov optimization techniques employed by Abouaomar et al. [138] enable dynamic resource allocation at the edge, while Poularakis et al. [18] address multidimensional constraints through joint optimization of service placement and request routing. Similarly, Fan et al. [126] propose a collaborative framework that minimizes total processing delay while ensuring queuing stability in multi-task scenarios. When considering delay-sensitive workloads specifically, Zhai et al. [127] demonstrate a remarkable 33.42× reduction in latency compared to traditional cloud deployments by intelligently distributing tasks based on their characteristics. Additionally, the double deep Q-network (DDQN) approach introduced by Pang et al. [139] effectively manages mobility while optimizing offloading decisions, while Qi et al. [140] present a two-timescale scheduling algorithm that balances service provider costs with stringent latency requirements.
IV-A2 Energy-Efficient Offloading
Strategic energy management across edge-cloud systems requires balancing computation and communication costs while optimizing resource allocation through specialized techniques.
In ECCC, prolonging the battery life of edge devices necessitates balancing computation and communication energy costs, especially for battery-powered devices [131]. As shown in Table IV, numerous approaches including PDCO [129], Hua et al. [117], and ETCRA [130] have demonstrated significant improvements in energy efficiency across edge-cloud systems. Particularly for battery-powered devices, strategic decision-making has to consider both computational and communication energy costs, weighing factors like task complexity, network conditions, and hardware energy profiles [131, 141]. While offloading computationally intensive tasks to cloud resources can reduce device energy consumption, communication costs under limited bandwidth or unstable networks often counterbalance these benefits, necessitating sophisticated optimization approaches. Beyond basic consumption metrics, advanced energy-efficient mechanisms consider additional dimensions such as power control, transmission scheduling [117], and user mobility patterns [142]. Furthermore, in heterogeneous IoT deployments with diverse device capabilities and constraints, efficient resource allocation becomes increasingly crucial for sustainable service delivery [122], with some approaches even incorporating renewable energy sources to create more environmentally responsible edge-cloud ecosystems [127].
Energy-efficient offloading mechanisms strive to minimize overall energy consumption by jointly considering computation and communication costs. Additionally, optimizing power control and transmission scheduling can further enhance energy efficiency [117]. User mobility introduces another layer of complexity to energy optimization, as dynamic network conditions and varying proximity to edge servers impact energy trade-offs [142]. In heterogeneous IoT environments with diverse device capabilities and energy constraints, efficient resource allocation becomes even more crucial for energy-efficient service delivery [122]. Prioritizing workloads based on their energy requirements and leveraging renewable energy sources can contribute to a more sustainable edge-cloud ecosystem [127].
IV-A3 Dependency-Aware Offloading
Complex applications often consist of multiple interconnected subtasks with dependencies, forming a dependency graph where some subtasks must finish before others can start [132]. Complex interdependence introduces complexities in task scheduling and resource allocation, potentially leading to suboptimal performance due to idle resources and increased latency if ignored. Therefore, efficient offloading mechanisms must consider these precedence constraints to optimize overall application execution. For example, the PRAISE algorithm in [132] addresses triangular dependencies by reconfiguring the task call graph into sequential layers for efficient resource allocation. Similarly, modeling task dependencies as a directed acyclic graph (DAG) optimizes offloading decisions, particularly in vehicle-edge-cloud environments [139]. Additionally, deep Q-networks (DQN) for real-time offloading in cloud-edge computing benefit from incorporating dependency awareness [143], enabling dynamic adaptation to changing environments without pre-set task priorities.
IV-A4 Multi-Objective Optimization Strategies
Contemporary edge-cloud systems increasingly require sophisticated offloading strategies that simultaneously optimize multiple performance metrics, reflecting the complex trade-offs inherent in distributed computing environments. Unlike single-objective approaches that focus on isolated metrics, multi-objective optimization techniques explicitly address the conflicts between competing goals such as minimizing latency while reducing energy consumption, maximizing throughput while maintaining quality of service, and balancing computational load while ensuring system reliability [22]. Multi-objective approaches typically employ Pareto optimization methods to identify solution sets that represent optimal trade-offs between conflicting objectives, enabling system designers to select configurations that best match their specific application requirements and operational constraints [144].
Recent research demonstrates the effectiveness of multi-objective approaches across diverse scenarios. For instance, Zhang et al. [22] propose a many-objective resource scheduling model that simultaneously considers time, cost, load balance, user satisfaction, and trust factors, employing advanced evolutionary algorithms to navigate the complex solution space. Similarly, Guo et al. [145] frame service deployment and task offloading as a multi-task, high-dimensional optimization problem that incorporates network reliability and load balancing considerations alongside traditional performance metrics. Furthermore, joint optimization techniques that combine energy efficiency with latency minimization have proven particularly effective in battery-powered edge environments [117], while approaches that integrate security and privacy considerations with performance optimization address the growing demands of trust-sensitive applications [130]. The challenge lies in developing efficient algorithms that can handle the increased complexity of multi-dimensional solution spaces while maintaining real-time decision-making capabilities essential for dynamic edge-cloud environments.
IV-B Resource Allocation Strategies
TABLE V: Summary of resource allocation methods.
Type Method Domain Metrics MC RA ECC
Deep Reinforcement Learning A2C-DRL [13] Edge-Cloud Environments Reward Value, Load-balancing
AC-DRL [146] 5G Edge-Cloud Continuum QoE, Energy Consumption
DA-HADRL [147] SAGINs with Cloud-Edge Energy, Latency
GRLDO [148] Client-Edge-Cloud Environment User Satisfaction
MADDPG [86] End-Edge-Cloud in 6G Networks Delay, Energy Consumption
Federated Learning FRL [93] Cloud-Edge-Terminal IoT Learning Speed, Adaptation
FDRL [95] Cloud-Edge Manufacturing Makespan, Energy Consumption
MARL-SF [149] IoV Edge Cloud Quality of Service
MARL-CA [150] Industrial Edge-Cloud Web3 Communication Rate, Stability
Optimization-Based DySCo [24] Edge-Cloud Infrastructure Deployment Cost, Delay
OTFAC [89] Cloud-Edge Collaborative IoT Delay, Energy Consumption
HDMO [136] Edge-Cloud Collaboration Time, Cost, Energy, Reliability
DaCM [151] Edge-Cloud Computing Long-term Profit, Delay
Model-Aware DSCF [152] Distributed Serverless Edge Inference Latency, Res. Consump.
DRL-LOA [153] Healthcare Cloud-Edge Power Usage, Latency
Wang et al. [154] IoT-Based Distribution Grid Convergence Rate, Stability
Domain-Specific Das et al. [155] Drone Communication Networks Latency, Scalability, Resource Util.
DA-DDQN [139] Vehicle-Edge-Cloud Latency, Energy Consumption
QRSRL [156] Train-Edge-Cloud Smart Systems Task Processing Delay
Informer-A3C [157] Vehicular Edge Cloud Computing Content Access Cost
Notes: RA = Resource Allocation.
Strategic allocation of computational resources across the edge-cloud continuum remains fundamental for maximizing performance in AI-driven collaborative environments. Accordingly, we explore three pivotal strategies: dynamic resource allocation, collaborative resource management, and reinforcement learning-based optimization, each designed to enhance task execution and system efficiency. By effectively distributing resources, advanced allocation approaches address the challenges of varying workloads and resource constraints in distributed environments. To illustrate their practical impact, Table V provides a comparative analysis of recent dynamic resource allocation methods in ECCC systems.
Notably, the table reveals a wide range of application domains, from IoT networks to vehicular edge clouds, alongside key performance metrics like latency, energy consumption, and quality of service, highlighting the adaptability and effectiveness of these strategies in diverse scenarios.
Resource allocation strategies in ECCC systems can be systematically organized according to their adaptation mechanisms and coordination approaches: (1) Temporal adaptation strategies (dynamic resource allocation) that respond to time-varying workloads and resource availability through real-time monitoring and reactive adjustment mechanisms; (2) Spatial coordination strategies (collaborative resource management) that optimize resource distribution across geographically distributed edge nodes and cloud infrastructure through inter-node cooperation and hierarchical coordination; (3) Learning-based optimization strategies (reinforcement learning approaches) that continuously improve resource allocation decisions through environmental feedback and experience accumulation. The classification framework emphasizes the complementary nature of these approaches, where temporal adaptation handles immediate demand fluctuations, spatial coordination optimizes cross-infrastructure efficiency, and learning-based methods provide long-term performance improvement through adaptive policy refinement.
IV-B1 Dynamic Resource Allocation
Adaptive resource management addresses fluctuating workloads and optimizes resource utilization through real-time monitoring and adjustment mechanisms, crucial for handling unpredictable application demands [64]. Approaches include dynamic NUMA node selection and core consolidation [24], optimizing service placement across the edge-cloud continuum by considering factors like delay and resource constraints. Dynamic task offloading mechanisms further optimize task placement and scheduling based on latency, energy consumption, and task dependencies [158].
In drone communication networks, dynamic resource allocation considers drone mobility, communication loads, and latency constraints for efficient real-time applications [155]. Similarly, for cloud-edge-terminal IoT networks, collaborative policy learning using federated reinforcement learning (FRL) enables dynamic task scheduling, adapting to changing workload demands [93]. Distributed serverless edge clouds utilize dynamic split computing frameworks to minimize inference latency while adhering to resource constraints by dynamically determining splitting points and maintaining container instance warm status [152].
Deep reinforcement learning techniques optimize task offloading decisions in 5G edge-cloud environments, maximizing quality-of-experience (QoE) while meeting latency requirements [146]. Additionally, multi-target-aware dynamic resource scheduling algorithms in multi-tier computing networks consider diverse QoS requirements, including computing, storage, bandwidth, and delay, to enhance resource flexibility and efficiency [159].
IV-B2 Collaborative Resource Management
Coordinated resource allocation across distributed infrastructure requires sophisticated management strategies that account for heterogeneous device capabilities and dynamic workload demands. For instance, in IoT-based distribution grids, resource optimization across cloud, edge, and end devices spans multiple timescales, with decisions at each level influencing others [154], necessitating sophisticated, interdependent resource management strategies. Dynamic resource allocation is also crucial for adapting to fluctuating demands and optimizing real-time performance [154].
One approach leverages optimal transport (OT) and FL principles. Specifically, an OT-based offline optimization model can initially allocate bandwidth and computation resources, followed by online optimization using federated actor-critic learning (OTFAC) [89]. In such dual-driven approach, edge servers can act as local aggregators within the federated learning process, improving efficiency and reducing delays [160]. Additionally, collaborative resource management benefits from a hybrid approach where edge servers not only collaborate with the cloud but also expand their local resources [151]. Such strategies allow edge servers to dynamically purchase cloud resources and expand local capacity for long-term cost optimization [151]. Consequently, such adaptability manages dynamic workloads and varying demands. Edge-cloud and edge-edge cooperation further enhance performance through efficient network-wide resource utilization [126].
IV-B3 Reinforcement Learning for Resource Management
Adaptive optimization mechanisms leverage reinforcement learning to optimize resource allocation by learning strategies through environmental feedback and performance-enhancing actions [161]. Specifically, these agents adapt to changing workloads and resource availability, leading to more efficient resource utilization. For example, RL optimizes task offloading decisions in vehicular edge-cloud computing (VECC) to minimize latency while maintaining quality of service [161], and addresses online dependent task offloading (ODTO) in collaborative edge computing by jointly optimizing network flow scheduling and resource allocation [158]. In addition, RL dynamically adjusts model parameters and offloading decisions, improving the adaptability and performance of edge-cloud systems [147]. Researchers have demonstrated RL’s effectiveness in various scenarios, including optimizing microservice placement [162], managing resources for delay-sensitive tasks [161], and handling dependent and parallel tasks [143]. Integrating RL with techniques like graph neural networks (GNNs) [148] and deep Q-networks (DQNs) [143] further enhances the agents’ ability to capture complex relationships and optimize offloading decisions. Multi-agent RL extends these capabilities by enabling decentralized learning and coordinated resource management across multiple servers [13].
IV-C Energy Efficiency Optimization
Sustainable power management across edge-cloud environments requires integrated optimization strategies balancing energy consumption with performance through comprehensive modeling and adaptive scheduling mechanisms [163, 11]. Energy-aware scheduling algorithms deliver dynamic adaptation via techniques like DVFS and intelligent sleep-wake management, with DVFO achieving 40% energy savings through co-optimization while Co-HDRL demonstrates 50-70% power reductions in IoT scenarios [164, 165]. Furthermore, joint optimization addresses fundamental trade-offs between efficiency and competing metrics through multi-objective approaches employing Pareto methods to balance conflicting goals [22, 117]. Consequently, advanced strategies demonstrate substantial impact, with green computing achieving 30-45% reductions in data center consumption through renewable integration and intelligent load balancing [142, 122]. Comprehensive analysis is detailed in LABEL:sec:supp-energy of the Supplementary Material.
V Privacy and Security Enhancement
Privacy and security mechanisms in ECCC systems form an interconnected protection ecosystem where techniques often complement and build upon each other rather than operating in isolation. Our classification organizes these mechanisms according to their protection scope and technical approach: (1) Data-centric privacy protection (Sec. V-A) encompassing techniques that protect raw data and sensitive information throughout the distributed learning process, including privacy-preserving learning, secure multi-party computation, differential privacy, and data anonymization; (2) Model-centric security protection (Sec. V-B) addressing threats to AI models themselves, including integrity protection, adversarial attack mitigation, secure updates, and trust management; (3) Communication-centric security (Sec. V-C) securing data transmission and inter-node communication across the edge-cloud continuum.
It is important to note that many privacy and security techniques are inherently non-orthogonal and often complement each other in practical deployments. For example, privacy-preserving learning methods frequently incorporate multiple approaches: differential privacy for statistical privacy guarantees, secure multi-party computation for cryptographic protection, and data anonymization for additional privacy layers. Similarly, federated learning serves dual purposes as both a model adaptation technique (Sec. III-B) and a privacy protection mechanism, demonstrating the multi-faceted nature of techniques in ECCC systems. Understanding these interdependencies is crucial for designing comprehensive security frameworks that leverage synergies between different protection mechanisms while avoiding redundant overhead.
V-A Data Privacy Protection
Comprehensive protection mechanisms safeguard sensitive information throughout its lifecycle in collaborative computing paradigms, addressing critical challenges of data security within distributed environments. We begin by examining privacy-preserving learning methods, such as federated learning and secure aggregation, enabling collaborative model training without direct data sharing. Subsequently, we delve into secure multi-party computation (SMPC), which allows joint function computation on private inputs. Additionally, differential privacy techniques, adding noise to protect individual data points while preserving aggregate insights, are considered. Finally, data anonymization approaches, transforming data to prevent individual identification, are explored.
V-A1 Privacy-preserving learning methods
Privacy enhancement in federated learning leverages techniques including differential privacy (DP) and secure aggregation to protect sensitive information [166]. DP adds carefully calibrated noise to model updates before aggregation, making it difficult to infer individual data points [167]. Meanwhile, secure aggregation uses cryptographic protocols like multi-party computation (MPC) to aggregate updates without revealing individual contributions to the server [168]. Additionally, zero-concentrated differential privacy (zCDP) offers tighter privacy guarantees than traditional DP in FL [169], and CaPC learning combines MPC, homomorphic encryption, and privately aggregated teacher models to achieve both confidentiality and privacy [170]. Lightweight encryption and aggregation techniques further enhance security and efficiency in federated learning systems [171].
V-A2 Secure Multi-Party Computation
Collaborative computation frameworks enable multiple parties to jointly compute functions over private inputs without revealing sensitive data through secure multi-party computation protocols [172]. SMPC protocols ensure each party learns only the function’s output, safeguarding data privacy and security, crucial for edge-cloud collaborative AI model training. For example, multiple hospitals could train a joint model on patient data without sharing raw data due to privacy regulations [170]. Additionally, SMPC can decompose complex AI tasks into smaller sub-tasks distributed across edge devices and the cloud, improving training efficiency [173]. Robust management frameworks, such as FlexSMC [174], address the dynamic edge-cloud environment by facilitating node discovery, establishing trust, and ensuring robustness against failures. Coded computation within SMPC, as in coded-MPC (CMPC) [175], further enhances performance through resource optimization, with adaptive gap entangled (AGE) polynomial codes [176] as a promising approach. Ensuring security against malicious attacks during MPC-enabled privacy-preserving training is also addressed by approaches like the one in [30], which provides active security even with a majority of malicious parties. Systems like Helen [177] demonstrate practical applications of maliciously secure cooperative learning for linear models.
V-A3 Differential Privacy Techniques
Privacy-preserving analysis frameworks offer robust protection for individual privacy by introducing carefully calibrated noise into data or computations [178], enabling meaningful insights extraction while safeguarding sensitive information. Specifically, DP mechanisms ensure negligible impact on analysis outputs from the presence or absence of a single individual’s data, preventing identification or inference of sensitive information [179, 180]. Common approaches include adding noise directly to raw data before analysis or perturbing the results of computations, such as model parameters, with the noise calibrated based on the computation’s sensitivity [179, 180]. For instance, in federated learning, DP protects individual device data during collaborative training by adding noise to local model updates before aggregation, preventing private data reconstruction from shared updates [181]. Similarly, in collaborative inference, DP can protect intermediate outputs and mitigate membership information leakage [182]. While noise addition inevitably impacts data utility and model accuracy [183], a key challenge is balancing privacy and utility, especially given the heterogeneous nature of edge devices and distributed edge-cloud systems [184, 185].
V-A4 Data Anonymization Approaches
Identity protection techniques transform data to prevent individual identification while enabling collaborative AI development, particularly crucial in distributed processing environments with unique privacy challenges [186]. Several approaches achieve this goal, including StyleID, which anonymizes faces in image datasets through identity disentanglement [187]. Specifically, StyleID projects images onto a GAN latent space, manipulating specific features to protect identities while preserving other image characteristics. Another key technique, DP, adds noise to data or computations to protect individual privacy while preserving useful insights [97]. For instance, differentially private federated tensor completion enables secure cloud-edge collaborative AIoT data prediction [97]. Dynamic anonymization techniques, often leveraging machine learning, adapt to specific contexts and data characteristics for more flexible and effective privacy protection [186]. Data anonymization in edge-cloud collaborative AI systems, where data may be distributed across multiple devices and locations, can be combined with other privacy-preserving techniques. Complementary approaches include SMPC and federated learning to create a more comprehensive privacy framework [97]. As an example, in cloud-edge collaborative inference, anonymization complements methods like PrivaScissors, which reduces mutual information leakage between intermediate model outputs and device data [182].
V-B Model Security
Comprehensive security frameworks address critical vulnerabilities throughout AI model lifecycles via multi-layered protection mechanisms against unauthorized modifications, adversarial attacks, inference attacks, and model poisoning [188, 97]. Robust protection encompasses integrity mechanisms utilizing cryptographic hashing and trusted execution environments, attack detection through adversarial training and anomaly detection, and secure update protocols implementing authentication and encrypted transmission [189, 190, 191]. Advanced approaches including FlexibleFL demonstrate effective federated learning defense through contribution-based mitigation, while blockchain-based solutions provide enhanced protection against coordinated attacks [99, 192, 193]. Additionally, runtime monitoring and auto-updating mechanisms enable adaptive responses to emerging threats while maintaining system resilience against novel attack patterns [194, 195]. Detailed technical analysis of integrity protection, attack detection, secure updates, and trust management is provided in LABEL:sec:supp-model-security of the Supplementary Material.
V-C Communication Security
Secure data exchange across heterogeneous infrastructures requires comprehensive protection through integrated encryption, authentication, and access control mechanisms addressing distributed edge-cloud architectural challenges [196, 197]. Adaptive security strategies balance protection strength with computational efficiency through frameworks integrating symmetric encryption for large-scale transfers, asymmetric encryption for secure key exchange, and TLS/VPN protocols establishing protected channels [198, 199]. Authentication incorporates multi-factor authentication, role-based access control, and hierarchical privilege management preventing unauthorized access while maintaining operational efficiency [200, 201]. Furthermore, encryption schemes leverage attribute-based approaches for fine-grained access control, while blockchain-based solutions provide decentralized trust management and immutable audit trails [194, 202, 203]. Coverage of secure transmission protocols, authentication mechanisms, encryption schemes, and blockchain solutions is detailed in LABEL:sec:supp-communication-security of the Supplementary Material.
VI Applications
ECCC enables transformative capabilities across diverse application domains by leveraging the unique synergy between edge proximity and cloud resources. Unlike traditional centralized or purely edge-based approaches, ECCC provides distinctive advantages through: (1) Hierarchical intelligence deployment, sophisticated AI models in the cloud guide lightweight edge models for real-time decision-making; (2) Dynamic workload distribution, computational tasks are intelligently partitioned based on latency requirements, privacy constraints, and resource availability; (3) Collaborative learning, federated and continual learning approaches enable privacy-preserving model improvement across distributed deployments; (4) Adaptive resource orchestration, dynamic scaling between edge and cloud resources based on real-time demands and network conditions. The following subsections demonstrate how these ECCC-specific capabilities create unprecedented opportunities for innovation across autonomous driving, smart healthcare, industrial automation, and smart cities, highlighting applications that are either impossible or significantly limited without the collaborative edge-cloud paradigm.
VI-A Autonomous Driving
Real-time processing of complex sensor data is crucial for safe and efficient autonomous driving. Edge-cloud collaboration offers a promising solution by distributing computational workloads between resource-constrained in-vehicle edge devices and powerful cloud servers. Specifically, edge devices can handle time-sensitive tasks like object detection and perception, potentially leveraging lightweight frameworks like Edge YOLO [204], which combines pruned and compressed networks for efficient multi-scale prediction on resource-limited platforms. In addition, the collaborative nature of these systems enables integration of crowdsourced data, enhancing perception capabilities by providing a more comprehensive environmental view, as demonstrated by LiveMap [205, 206], which efficiently processes and integrates multi-vehicle data for real-time dynamic map generation. Beyond perception, LLMs are emerging as powerful tools for motion planning and decision-making [207]. For example, systems like EC-Drive utilize LLMs across both cloud and edge, adapting to data drifts and optimizing communication for reduced latency. Similarly, collaborative policy learning through DRL [205] offers a mechanism for dynamic scheduling and resource allocation, optimizing vehicle scheduling and offloading decisions for improved system robustness and adaptability. Adaptive edge-cloud collaboration, aided by large vision models like in LAECIPS [208], further enhances performance and adaptability of IoT-based perception systems through efficient hard input mining. Concurrently, the fusion of DRL and edge computing facilitates real-time monitoring and control optimization in IoT environments [21], enabling predictive system state analysis and dynamic resource allocation for enhanced control.
VI-B Smart Healthcare
Figure 6: Integration of federated learning, edge-cloud platforms, and IoMT optimization in smart healthcare.
AI-driven collaborative computing transforms smart healthcare systems into efficient, secure, and personalized ecosystems, as shown in Fig. 6. We explore how AI-driven edge-cloud collaboration revolutionizes healthcare delivery and improves patient outcomes by investigating FL for medical image analysis, edge-cloud platforms for remote patient monitoring, and resource allocation optimization for Internet of medical things (IoMT) devices. FL offers a promising technique for training AI models on distributed medical datasets without compromising patient privacy, allowing collaborative model training across multiple healthcare providers while keeping sensitive data localized [209]. Specifically, FL enables robust diagnostic model development, including COVID-19 diagnosis using multimodal data like X-rays and ultrasound images [210]. Edge-cloud platforms are crucial for enabling remote patient monitoring and personalized healthcare [211], leveraging edge device computation and cloud storage for real-time monitoring and personalized interventions. For example, wearable sensors and IoMT devices collect patient data, process it at the edge for immediate feedback, and transmit it to the cloud for analysis and storage [28]. AI algorithms deployed on edge devices analyze this data, providing personalized recommendations and improving intervention effectiveness [212]. Efficient resource allocation is also essential for optimizing IoMT device performance and ensuring timely healthcare delivery [213]. Consequently, AI-driven task offloading and resource management dynamically allocate tasks and resources between edge and cloud based on real-time demands and availability, minimizing latency and improving system efficiency. Hierarchical FL approaches further enable multi-party collaboration and efficient model training for anomaly detection in IoMT data, leveraging digital twins and edge cloudlets for real-time patient health insights and timely interventions [214]. Integrating FL within the healthcare Metaverse enhances these capabilities by improving privacy, scalability, and interoperability [215].
VI-C Industrial Automation
Intelligent automation systems benefit significantly from collaborative computing architectures that deliver real-time fault diagnosis and predictive maintenance capabilities [21]. Leveraging both cloud computational power and edge real-time processing capabilities, industrial automation systems achieve greater agility through optimized production processes and resource allocation [80], where cloud-based platforms analyze historical data to generate optimized production schedules while edge devices adapt these schedules in real-time based on local conditions. Distributed intelligence across multiple edge devices handles complex manufacturing tasks more efficiently and enhances fault tolerance since individual device failures do not necessarily compromise the entire system [216]. As edge intelligence, anomaly detection, cloud-edge collaboration, and distributed intelligence converge, industrial automation systems evolve toward more sophisticated and autonomous manufacturing processes [65].
VI-D Smart Cities
Urban infrastructure optimization leverages collaborative computing to enhance traffic management and environmental monitoring capabilities [217, 142]. Edge devices collect and analyze real-time data for adaptive traffic control and congestion reduction, while distributed sensor networks monitor environmental factors like air quality and noise levels. Subsequently, cloud platforms aggregate and analyze this data to inform urban planning and environmental protection. In addition, AI-driven resource allocation optimizes smart grid energy consumption [218]. Specifically, edge devices monitor energy usage, and cloud-based algorithms dynamically adjust energy distribution to minimize waste and improve grid stability. FL is crucial for privacy-preserving data analysis and decision-making in these applications [16, 219, 220]. Through FL, stakeholders such as government agencies and private companies can collaboratively train models on combined data without sharing sensitive information, leading to more accurate and robust models for urban planning and public safety.
VII Performance Analysis
VII-A Evaluation Metrics
Figure 7: Evaluation metrics, which are categorized into four primary groups: System performance, model efficiency, quality of service, and security and privacy. Each group contains specific evaluation dimensions (e.g., latency, inference time, availability, and data breach frequency), enabling comprehensive assessment of collaborative performance, resource usage, user experience, and trustworthiness.
Comprehensive performance assessment of collaborative computing systems requires multidimensional evaluation that addresses the unique characteristics of distributed computing paradigms. Unlike traditional centralized systems, ECCC performance evaluation must consider the complex interplay between edge proximity and cloud scalability, heterogeneous resource constraints, and dynamic network conditions. We categorize the relevant metrics into four primary groups tailored for edge-cloud environments: system performance, model efficiency, quality of service, and security and privacy. As illustrated in Fig. 7, comprehensive evaluation spans system-level, model-level, service-level, and trust-level metrics, considering real-world constraints unique to edge-cloud deployments such as network variability, compute heterogeneity, device mobility, and multi-tier resource orchestration.
Performance evaluation in ECCC environments must address several unique challenges: (1) Cross-tier latency analysis, measuring end-to-end response times that include network transmission delays, edge processing times, and cloud computation overhead; (2) Resource utilization efficiency, evaluating how effectively computational, storage, and communication resources are distributed and utilized across the edge-cloud continuum; (3) Adaptation overhead, quantifying the costs associated with dynamic model partitioning, task migration, and real-time workload redistribution; (4) Collaborative intelligence effectiveness, assessing the performance gains achieved through edge-cloud cooperation compared to standalone edge or cloud-only deployments; (5) Fault tolerance and resilience, measuring system robustness against edge device failures, network partitions, and cloud service interruptions. Specialized considerations necessitate specialized evaluation frameworks that capture the distributed, heterogeneous, and dynamic nature of edge-cloud systems while providing meaningful comparisons with traditional computing paradigms.
VII-A1 System Performance
Throughput, latency, and scalability are crucial metrics for evaluating the effectiveness and responsiveness of ECCC systems [116, 20, 221]. Throughput measures the volume of data processed or tasks completed within a given timeframe, with higher throughput indicating efficient handling of substantial workloads [222]. Conversely, latency quantifies the delay between request initiation and completion; minimizing latency is paramount for real-time and interactive applications where responsiveness is critical [221]. Scalability, in turn, refers to a system’s ability to adapt to increasing workloads or expanding infrastructure without performance degradation, accommodating growth in user demands or data volume while maintaining acceptable throughput and latency [203]. The distributed and heterogeneous nature of edge-cloud environments amplifies the significance of these metrics, as the interplay between edge and cloud resources necessitates careful optimization. For example, the study in [116] benchmarks a container-based edge computing system, demonstrating the impact of data synchronization on scalability across different deployment strategies. Similarly, the SPEC-RG reference architecture in [20] provides a framework for analyzing resource management and performance, while [221] highlights the potential for performance inversion due to edge queuing delays. Evaluating these metrics is essential for understanding system efficiency and informing design choices for optimized performance in resource-constrained and dynamic edge-cloud environments [62].
VII-A2 Model Efficiency
Efficient AI models are essential for effective deployment and operation within the resource-constrained edge-cloud continuum. Key metrics for quantifying model efficiency include computational cost, memory usage, and inference time, informing model selection and deployment strategies. Computational cost, typically measured in FLOPS [223], quantifies processing demands. Higher FLOPS values indicate greater computational requirements and potentially increased energy consumption, impacting resource-limited edge devices and overall system sustainability [224, 225]. Memory usage reflects a model’s storage footprint; large models may exceed the limited capacity of edge devices [226], necessitating compression techniques like pruning, quantization, or knowledge distillation [70]. Inference time, the duration for processing input and generating output, directly impacts application responsiveness [226]. In latency-sensitive applications like autonomous driving or real-time video analytics, minimizing inference time is crucial [20]. Efficiency optimization often involves balancing model accuracy and computational efficiency, potentially employing strategies like model partitioning across different devices [227].
VII-A3 Quality of Service
Evaluating the performance of edge-cloud collaborative systems necessitates an analysis of QoS metrics, which directly reflect the system’s capacity to fulfill user expectations and provide consistent, dependable service [228]. Key QoS metrics include availability, reliability, and user satisfaction. Availability quantifies the proportion of time a system remains operational and accessible, a crucial aspect in edge-cloud systems where interruptions significantly impact user experience and application performance [20]. High availability relies on robust system design, redundant infrastructure, and effective fault tolerance mechanisms, further enhanced by dynamic resource management and adaptive optimization techniques [229]. Reliability, in turn, assesses the system’s ability to consistently deliver the expected service level over time, encompassing fault tolerance, error recovery, and performance maintenance under varying workloads [230]. Dependency-aware offloading and collaborative resource management contribute to improved reliability by optimizing task placement and resource allocation. While often subjective, user satisfaction remains a crucial QoS indicator, reflecting the overall user experience, including performance, ease of use, and perceived value [231]. In edge-cloud systems, factors such as latency, throughput, and the quality of AI-driven services influence user satisfaction. Specifically, QoE serves as a proxy for user satisfaction, directly linked to QoS parameters like latency and loss in applications like video streaming [232]. Similarly, in industrial automation, the relationship between QoS and QoE is critical for optimizing network resource management and ensuring customer satisfaction.
VII-A4 Security and Privacy
In edge-cloud collaborative systems, robust security and privacy are crucial for maintaining trust and protecting sensitive information, especially given the sensitive nature of processed data [233]. Therefore, evaluating these systems requires metrics focusing on data breaches, encryption strength, and regulatory compliance. Data breach frequency directly indicates system vulnerability by quantifying unauthorized accesses within a given timeframe [198]. A lower frequency suggests stronger security, and analyzing breach characteristics can inform improved protocols, potentially leveraging frameworks like the transparency, accountability, and blockchain framework [233]. Encryption strength assesses cryptographic robustness, including key length, algorithms, and resistance to attacks. Strong encryption is paramount in distributed edge-cloud environments [194], and combining techniques like lightweight symmetric encryption and ciphertext-policy attribute-based encryption can enhance security [194]. Additionally, compliance with regulations like GDPR and CCPA is essential. Metrics in this area assess adherence to data anonymization, user consent, and data retention policies [234]. Evaluating privacy protectability, along with techniques like PrivaScissors, can further enhance privacy protection [235, 182].
VII-B Benchmarking Methods
Performance benchmarking of edge-cloud collaborative systems requires standardized evaluation approaches to facilitate meaningful comparisons across diverse system designs and optimization strategies.
TABLE VI: Comparison of edge-cloud benchmarking tools and platforms.
Tool Focus Area Key Features Advantages Limitations Use Case Scenarios
Edge AIBench [236] End-to-end edge AI performance Models typical edge scenarios (ICU monitoring, autonomous vehicles), comprehensive client-edge-cloud evaluation Realistic application scenarios Comprehensive performance metrics Multi-tier evaluation coverage Limited customization options Fixed scenario constraints Domain-specific focus Healthcare monitoring, autonomous driving, real-time inference applications
EdgeBench [237] Workflow-based edge computing Customizable workflow logic, configurable data storage, flexible stage distribution across computing tiers High customization flexibility Workflow-centric approach Adaptable to diverse applications Complex configuration process Requires workflow expertise Limited standardized scenarios Custom application workflows, multi-stage processing pipelines, research environments
BenchFaaS [238] Serverless edge computing Heterogeneous testbeds (VMs, Raspberry Pi, cloud platforms), FaaS performance evaluation Serverless focus Heterogeneous platform support Function-level granularity Limited to FaaS paradigm Cold start overhead Platform-dependent results Serverless applications, function-as-a-service deployment, microservices evaluation
KheOps [239] Reproducible edge-to-cloud experiments Collaborative environment, experiment repository, notebook interface, multi-platform methodology Strong reproducibility support Collaborative features Multi-platform compatibility Learning curve for new users Resource-intensive setup Limited real-time capabilities Research collaboration, experimental reproducibility, academic studies
OpenPerf [240] Open-source project benchmarking General performance evaluation framework, standardized metrics, community-driven development Open-source transparency Community contributions Standardized approach Not edge-cloud specific Generic rather than specialized Limited edge device support General performance testing, open-source projects, baseline comparisons
Table VI provides a comprehensive comparison of existing benchmarking tools and their respective strengths and limitations across different edge-cloud deployment scenarios. Standardized datasets and workloads, discussed in Sec. VII-B1, provide a common basis for evaluation, minimizing variability stemming from data discrepancies. Well-defined testing environments, explored in Sec. VII-B2, further enhance reliability and reproducibility by controlling hardware and software configurations, network setups, and other influential environmental factors. Performance comparison frameworks and evaluation methodologies, including simulation, emulation, and real-world testing approaches, offer structured methodologies and tools for identifying performance bottlenecks and areas for improvement in edge-cloud systems, with detailed analysis provided in the Supplementary Material.
The choice of appropriate benchmarking tools depends on several factors including application domain, system architecture, and evaluation objectives. Edge AIBench [236] excels in scenarios requiring realistic application modeling with comprehensive end-to-end performance analysis, making it particularly suitable for healthcare and autonomous driving applications where domain-specific workloads are crucial. EdgeBench [237] offers superior flexibility for research environments and custom applications that require tailored workflow configurations, though it demands significant expertise in workflow design and configuration. BenchFaaS [238] provides specialized evaluation capabilities for serverless edge computing paradigms, offering detailed insights into function-level performance characteristics but with limited applicability outside the FaaS domain. KheOps [239] stands out for collaborative research environments where reproducibility and experiment sharing are paramount, though it may require substantial initial setup overhead. The selection process should consider factors such as workload representativeness, customization requirements, evaluation scope, and available expertise. Additionally, the benchmarking tool should align with the specific performance metrics of interest, whether focusing on latency, throughput, energy efficiency, or comprehensive system-level analysis.
VII-B1 Standard Datasets and Workloads
Consistent and comparable performance evaluation across different studies of ECCC systems requires the use of standard datasets and workloads [241]. Unified standardization clarifies the strengths and weaknesses of various architectures, optimization techniques, and algorithms, enabling researchers to build upon existing work and accelerate the development of innovative solutions. For example, Edge AIBench [236] offers benchmarks modeling typical edge scenarios like ICU patient monitoring and autonomous vehicles, focusing on data distribution and workload collaboration across client devices, the edge, and the cloud for comprehensive end-to-end performance evaluation. Similarly, EdgeBench [237] provides a workflow-based benchmarking approach, allowing customization of workflow logic, data storage, and the distribution of workflow stages across computing tiers to tailor benchmarks to specific application scenarios. Appropriate workload selection is also crucial [20]. Specifically, workloads should reflect real-world application characteristics, including data intensity, computational complexity, and communication patterns, particularly in distributed intelligence scenarios where data and model distribution, as well as communication overhead for model training and inference, are critical factors. KheOps [239] further emphasizes reproducibility in edge-to-cloud experiments by providing a collaborative environment with an experiment repository, a notebook environment, and a multi-platform methodology, promoting transparency and rigor in performance evaluation. The availability of standard datasets and workloads also facilitates the development of new optimization techniques [242], allowing researchers to evaluate their methods on established benchmarks [227].
VII-B2 Testing Environments
Robust benchmarking of edge-cloud collaborative systems mandates controlled testing environments [243]. Specifically, accurately capturing performance characteristics requires careful configuration and control due to the complex interactions among edge devices, network infrastructure, and cloud servers. Achieving reliable benchmarking results necessitates meticulous configuration of hardware, software, and network parameters to minimize external influences and ensure reproducibility. The diverse hardware platforms in edge deployments, from resource-constrained devices to powerful edge servers, necessitate standardized testing environments for fair comparisons across different system architectures [236]. For example, BenchFaaS and KheOps employ heterogeneous testbeds, encompassing virtual machines, Raspberry Pis, and various cloud platforms, to evaluate performance under diverse hardware and network conditions [238, 239]. Additionally, accurate measurement of key performance indicators (KPIs), such as latency, throughput, and energy consumption, requires precise control over network conditions, including bandwidth, latency, and packet loss. Standardized software configurations, including operating systems, libraries, and runtime environments, are also essential for eliminating inconsistencies and ensuring comparability across experiments. Representative benchmarks that accurately reflect real-world scenarios, encompassing data volume, processing requirements, and communication patterns, are necessary for comprehensive performance evaluation.
Performance comparison frameworks offer structured methodologies for systematic evaluation of diverse ECCC architectures and optimization techniques, providing standardized approaches to performance assessment that enable identification of bottlenecks and derivation of best practices [240]. Comprehensive frameworks encompass multifaceted evaluation metrics including system performance (throughput, latency, scalability), model efficiency, quality of service, and security considerations [244], alongside standardized datasets and workloads representative of real-world applications [243]. Essential components include guidelines for establishing controlled testing environments for reproducible results [245] and tools for result analysis incorporating statistical and visualization techniques to identify performance bottlenecks and understand design trade-offs [20]. Evaluation methodologies span simulation for cost-effective scenario exploration, emulation for realistic yet controlled assessment using network emulation tools [128], and real-world testing for comprehensive validation, each offering distinct advantages in capturing system complexities while addressing specific evaluation requirements. Detailed performance comparison frameworks, systematic evaluation approaches, and comprehensive methodology analysis are provided in LABEL:sec:supp-frameworks and LABEL:sec:supp-methodologies of the Supplementary Material.
VII-C System Design Evaluation
Comprehensive performance assessment across ECCC paradigms requires multi-dimensional analysis through architectural comparisons, optimization evaluations, and systematic trade-off analysis [246, 123, 247]. Architectural configurations reveal how cloud-centric solutions prioritize scalability versus edge-centric approaches like Armada emphasizing minimal latency, while hybrid systems including AppealNet and EdgeMatrix leverage predictive offloading and multi-agent algorithms for optimal balance. Additionally, optimization technique analysis examines load balancing effectiveness through two-timescale frameworks, collaborative edge-edge cooperation, and reinforcement learning-based adaptive management demonstrating superior performance in distributed scenarios [248, 249, 126]. Furthermore, trade-off analysis addresses design considerations balancing performance versus energy consumption, latency-accuracy optimization, and security-usability considerations [142, 127, 62]. Detailed architectural comparisons, optimization analysis, and trade-off methodologies are provided in LABEL:sec:supp-system-design-eval of the Supplementary Material.
VIII Challenges and Future Directions
VIII-A Technical Challenges
Figure 8: Technical challenges, which span across model optimization, system integration, and performance bottlenecks, with domain-specific requirements shaping solution design.
As illustrated in Fig. 8, AI-driven ECCC systems face multifaceted technical challenges spanning model optimization under resource constraints, integration of heterogeneous hardware/software, performance bottlenecks, and addressing domain-specific concerns such as real-time processing and scalability. Addressing the interconnected nature of challenges requires comprehensive solutions that balance accuracy, efficiency, and scalability while ensuring seamless interoperability and overcoming practical deployment barriers.
VIII-A1 Model Optimization Issues
Model optimization for edge-cloud deployment fundamentally challenges the accuracy-efficiency trade-off, where complex models demand substantial resources unsuitable for edge deployment while smaller models sacrifice accuracy [250]. Resource limitations including processing power, memory capacity, and energy availability, particularly for IoT devices [142, 251], compound deployment complexities alongside scaling challenges from communication overhead, resource contention, and device heterogeneity [252, 253, 15]. Effective optimization necessitates integrated approaches combining pruning, quantization, knowledge distillation [48] with resource-aware strategies like dynamic model selection, offloading [62, 39], energy-aware scheduling [254], and adaptive model partitioning [195]. LLMs further intensify optimization challenges, requiring specialized inference techniques for resource-constrained environments, with solutions like Shoggoth [255] demonstrating online knowledge distillation to enhance accuracy while managing data drift and computational offloading [256].
VIII-A2 System Integration Challenges
System integration in ECCC demands sophisticated heterogeneity management across diverse hardware capabilities, operating systems, network bandwidths, and communication protocols between resource-constrained edge nodes and powerful cloud servers [257, 225]. Interoperability challenges stem from differing data formats, communication protocols, and software interfaces, necessitating complex data transformations and sophisticated scheduling algorithms that consider device-specific capabilities [258, 259, 128]. Standardized protocols and well-defined interfaces become essential for seamless data exchange, simplified development, and enhanced scalability, with containerization (Docker [257]) and WebAssembly [258] providing portability solutions alongside AI-driven task placement optimization [260, 259]. Standardization particularly benefits critical applications like autonomous driving [53] and smart healthcare [142] requiring reliable real-time data exchange, while promoting vendor interoperability and fostering open ecosystems [247, 261, 229].
VIII-A3 Performance Bottlenecks
ECCC systems encounter significant performance bottlenecks from communication overhead introducing latency and bandwidth consumption that impacts real-time applications [116, 221, 262], while edge devices struggle with limited processing power, memory, and energy constraints when executing complex AI models and handling growing IoT data volumes [66, 263, 264, 29]. Resource limitations become particularly acute during test-time adaptation scenarios where models must adjust to new data distributions while maintaining strict latency and energy budgets [115]. Storage limitations exacerbated by device heterogeneity necessitate intelligent caching, data compression, distributed storage solutions like EdgeKV [63], and comprehensive data lifecycle management [48, 265, 266, 267]. Data distribution shifts across geographically distributed edge nodes create additional performance challenges, requiring adaptive load balancing and intelligent workload redistribution to maintain system efficiency. Real-time applications including autonomous driving, industrial automation, and smart healthcare demanding ultra-low latency [208, 268, 21] require integrated architectural innovations combining hierarchical edge computing with intelligent resource management [269], utilizing frameworks like EdgeMatrix [247] for multi-agent scaling, KubeEdge [195] for automated provisioning, predictive offloading [123], intelligent caching [270], and hardware acceleration technologies for optimal performance.
VIII-A4 Practical Implementation Barriers
Real-world ECCC deployment reveals substantial barriers extending beyond theoretical optimization to encompass operational complexities, economic constraints, and integration challenges, where device heterogeneity necessitates standardized middleware layers and adaptive frameworks for automatic capability detection [20, 28, 25, 60, 49]. Network connectivity issues require robust offline operation modes and intelligent synchronization mechanisms [35], while deployment complexity often results in 40-60% timeline overruns due to integration challenges and vendor compatibility issues [226]. Economic viability concerns balance substantial upfront investments against long-term benefits, with edge computing reducing bandwidth costs by 30-50% but requiring 3-5 year payback periods [142, 45], while technical pitfalls include underestimated communication overhead [21, 253], data consistency issues requiring eventual consistency models [10], and resource contention demanding sophisticated isolation algorithms [13, 29]. Successful implementations emphasize phased rollout strategies, hybrid architectures, comprehensive security frameworks, and dedicated cross-functional teams [18, 19, 14].
VIII-B Future Research Directions
Future research in ECCC must address critical scalability challenges while advancing revolutionary computing paradigms [57]. Short-term objectives include developing adaptive model partitioning algorithms [58] and standardized APIs for heterogeneous device integration [200], alongside lightweight containerization technologies optimized for resource-constrained environments. Furthermore, medium-term objectives should focus on neuromorphic computing integration [271, 272] for ultra-low-power edge inference and quantum-resistant security protocols [273, 274] for long-term protection. Additionally, long-term research directions encompass quantum edge computing [275, 276] for solving intractable optimization problems and fully autonomous edge-cloud ecosystems capable of self-optimization [259]. Consequently, implementation feasibility analysis indicates that standardization efforts represent the most immediately achievable objectives, while quantum computing integration and autonomous systems present longer-term challenges requiring significant advances in underlying technologies [17].
VIII-B1 Advanced AI Techniques
Figure 9: Overview of advanced AI techniques.
Advanced AI technologies represent the next frontier in ECCC, offering transformative capabilities through intelligent automation [277] and revolutionary computational paradigms. As illustrated in Fig. 9, algorithm optimization techniques including NAS [15] and AutoML [278, 279] enable autonomous system optimization by automatically discovering optimal architectures tailored to edge-cloud constraints, thereby reducing manual configuration overhead across heterogeneous infrastructure. Furthermore, intelligent interaction capabilities encompass AI agents [260] that facilitate natural human-system interaction and autonomous decision-making for adaptive responses to dynamic conditions. Additionally, hardware & infrastructure innovations particularly neuromorphic computing [271], quantum edge computing [275], and 6G network integration [26, 280] promise revolutionary advances in computational efficiency and connectivity. Consequently, the synergistic integration of these technologies enables unprecedented system autonomy, where NAS optimizes architectures for edge constraints while AI agents coordinate workflows and quantum computing tackles intractable optimization problems, collectively transforming edge-cloud system capabilities.
$\bullet$ Neural Architecture Search  NAS automates the design and optimization of neural network architectures, aiming to discover models that outperform manually designed ones in accuracy and efficiency for edge-cloud collaborative environments [281]. NAS algorithms explore a vast search space of potential architectures, evaluating their performance to identify architectures tailored to the specific constraints of edge-cloud systems, such as limited resources and bandwidth [144, 216].  Approaches like reinforcement learning, evolutionary algorithms, and gradient-based NAS show promise in finding optimal architectures for specific tasks and resource constraints [58], and combining NAS with model optimization techniques like compression and adaptation further improves performance and efficiency [282, 283].  Ultimately, integrated optimization addresses the challenges of deploying complex AI models in heterogeneous edge-cloud environments [284].
$\bullet$ AutoML Automated machine learning (AutoML) streamlines AI model development and deployment in edge-cloud systems by automating processes like model selection, hyperparameter tuning, and deployment [285, 277]. Automated optimization becomes crucial in edge-cloud environments where diverse hardware and software coupled with resource constraints make manual optimization complex [286].  For example, techniques like FAST-DAD [287] distill complex models into smaller, deployable ones for edge devices, while frameworks like H2O AutoML simplify model selection and hyperparameter optimization [277].  Language-driven systems like AutoMMLab [278] further democratize AI development.  However, challenges remain in transparency, flexibility for complex scenarios, and consistent coverage of the entire ML workflow [279]. In large-scale environments, sampling-based AutoML and human-centered systems like OmniForce [288, 286] show promise, while automatically generating data-dependent configurations improves model performance [289].
$\bullet$ Large Language Models Optimizing LLM inference for edge devices involves addressing limited resources and real-time performance, necessitating research on optimized inference techniques and model compression methods like Agile-Quant [290, 291].  Approaches like expert-wise bitwidth adaptation in EdgeMoE [290] can further enhance on-device efficiency.  Integrating LLMs for edge-cloud orchestration, including dynamic resource allocation and task offloading, promises more efficient systems, as demonstrated by VELO and EC-Drive [292, 207].  Developing specialized LLM architectures for edge-cloud environments, including low-rank adaptation (NetGPT [293]), model partitioning, and split learning/inference [256], is a critical research direction.
$\bullet$ AI Agents Intelligent agents capable of autonomous decision-making and learning within edge-cloud systems enhance adaptability and efficiency [15, 294].  AI agents facilitate continuous learning on edge devices, addressing bandwidth limitations and connectivity issues through approaches like NeuroEvolutionary learning [282]. Collaborative learning improves model accuracy, particularly in heterogeneous networks [11, 10], and these agents optimize system performance through dynamic task offloading and resource allocation [253, 284].  Integration with lifelong learning frameworks like KubeEdge-Sedna further enhances automation [195].
$\bullet$ Embodied AI Embodied AI, focusing on agents that interact with their environment, requires new algorithms and frameworks tailored for resource-constrained edge devices [295].  Algorithms for robot control and coordination must consider latency and bandwidth limitations [296], while efficient task offloading and resource allocation optimize performance [247]. Furthermore, exploring embodied AI applications for multi-robot collaboration and dynamic visual servoing presents significant research opportunities [297, 296]. Ultimately, developing robust and adaptive algorithms that effectively leverage distributed computational resources represents a critical research direction for advancing embodied AI in edge-cloud environments [298].
$\bullet$ Neuromorphic Computing Inspired by the biological brain, neuromorphic computing promises energy-efficient AI models for edge-cloud environments [271, 299]. Neuromorphic hardware, like SNNs, offers potential energy savings compared to traditional hardware [272], while inherent fault tolerance improves robustness [300].  Integrating neuromorphic hardware into data centers presents challenges in algorithms, software, and benchmarks [301].  However, its potential benefits for edge robotics drive research towards ultra-low-power designs and 3D integration technologies [298, 302].
$\bullet$ Quantum Edge Computing Quantum computing offers revolutionary potential for edge-cloud applications by tackling classically intractable problems [275]. Consequently, developing edge-tailored quantum algorithms [276] and robust quantum communication protocols [303], potentially in hybrid approaches, are key research avenues. However, significant security [273] and privacy concerns [274] necessitate quantum-resistant techniques. Investigating the associated societal impacts [17] is also crucial for responsible deployment.
$\bullet$ 6G Network Integration 6G networks, with ultra-low latency and high bandwidth, promise to enhance edge-cloud systems [26].  Low latency enables real-time processing, while high bandwidth facilitates seamless data transfer, supporting resource-intensive AI tasks [304, 280].  Advanced features like network slicing and cell-free MIMO optimize performance [280], and integration with digital twins and pervasive AI unlocks new possibilities [305, 306].  Addressing challenges in network architecture, resource allocation, and security through AI-driven solutions is crucial [307, 27].
VIII-B2 System Improvements
Efficient and dependable AI-driven ECCC necessitates crucial system improvements, primarily focusing on dynamic resource management, adaptive optimization, and fault tolerance.
$\bullet$ Dynamic Resource Management Dynamic resource management optimizes performance and resource utilization in the fluctuating demands of edge-cloud environments through efficient, real-time resource allocation and reallocation.  Key aspects include algorithms that dynamically adjust resource allocation based on real-time workload characteristics and available resources [126], such as the energy-aware and trust-collaboration cross-domain resource allocation (ETCRA) algorithm, which minimizes system functions while adhering to latency and trust constraints [130].  Collaborative resource management coordinates resource allocation across multiple edge nodes and the cloud for enhanced system efficiency [15].  Additionally, integrating RL allows systems to learn optimal resource allocation strategies based on environmental feedback [21], while dynamic algorithms with initial allocation and real-time updating based on trust assessment improve performance [51]. Multi-timescale algorithms address the varying optimization periods of different sub-problems like service placement and task scheduling within edge-cloud cooperation [126].
$\bullet$ Adaptive Optimization  Adaptive optimization dynamically adjusts system parameters and optimization strategies in response to real-time data and evolving conditions to maximize efficiency. For instance, the dynamic resource allocation mechanism in [21] ensures efficient scheduling of edge resources for global optimization.  Adaptive training with small batches, as proposed in Shoggoth [255], allows edge model adaptation under limited computing power.  Similarly, incremental learning approaches periodically retrain edge AI models using real-time data [267], optimizing prediction accuracy and reducing cloud dependency.  Dynamic model selection based on resources and application requirements [15] and decentralized adaptive scheduling using A3C learning and R2N2 [308] are other crucial aspects.  Additionally, the two-stage iterated greedy optimization (TIGO) algorithm addresses changing user requirements for dynamic microservice placement [120], while adaptive switching between fog and cloud instances based on response time enhances performance [309].
$\bullet$ Fault Tolerance  Fault tolerance ensures reliable operation by implementing fault detection, isolation, and recovery strategies.  DeepFT [310], for example, uses a deep surrogate model for proactive fault prediction and optimized task scheduling.  Ensuring service resilience for critical nodes is addressed by confidence-aware resilience models like CAROL [311], which leverages generative neural networks for QoS prediction and system recovery.  Programmable asset orchestration at the edge contributes to operational resilience, with systems like RECS [312] dynamically managing services and resources during incidents.  Finally, distributed data auditing schemes, often incorporating techniques like FSS [188], ensure data integrity for reliable AI applications.
IX Conclusion
In this survey, we comprehensively examine AI-driven edgecloud collaborative computing, focusing on distributed intelligence and model optimization. We first outline foundational architectures, then categorize key techniques into model optimization, resource management, and privacy-security enhancements. Additionally, we highlight trends in AI-driven resource allocation and model adaptation. Despite notable progress, challenges in system heterogeneity, security, and application diversity persist, urging further research into advanced AI techniques and emerging technologies like 6G, LLM integration, AI agents, embodied AI, edge quantum computing, etc.